### v1

# Fully Dynamic Matching and Ordered Ruzsa-Szemerédi Graphs

Abstract
We study the fully dynamic maximum matching problem. In this problem, the goal is to efficiently maintain an approximate maximum matching of a graph that is subject to edge insertions and deletions. Our focus is particularly on algorithms that maintain the edges of a $(1-\varepsilon)$-approximate maximum matching for an arbitrarily small constant $\varepsilon>0$. Until recently, the fastest known algorithm for this problem required $\Theta(n)$ time per update where $n$ is the number of vertices. This bound was slightly improved to $n/(\log^{*}n)^{\Omega(1)}$ by Assadi, Behnezhad, Khanna, and Li [STOC’23] and very recently to $n/2^{\Omega(\sqrt{\log n})}$ by Liu [ArXiv’24]. Whether this can be improved to $n^{1-\Omega(1)}$ remains a major open problem.
In this paper, we present a new algorithm that maintains a $(1-\varepsilon)$-approximate maximum matching. The update-time of our algorithm is parametrized based on the density of a certain class of graphs that we call Ordered Ruzsa-Szemerédi (ORS) graphs, a generalization of the well-known Ruzsa-Szemerédi graphs. While determining the density of ORS (or RS) remains a hard problem in combinatorics, we prove that if the existing constructions of ORS graphs are optimal, then our algorithm runs in $n^{1/2+O(\varepsilon)}$ time for any fixed $\varepsilon>0$ which would be significantly faster than existing near-linear in $n$ time algorithms.
Our second main contribution is a better upper bound on density of both ORS and RS graphs with linear size matchings. The previous best upper bound was due to a proof of the triangle-removal lemma from more than a decade ago due to Fox [Annals of Mathematics ’11].
1 Introduction
We study dynamic algorithms for the maximum matching problem, a cornerstone of combinatorial optimization. Given a graph $G=(V,E)$ a matching $M\subseteq E$ is a collection of vertex-disjoint edges. A maximum matching is a matching of largest possible size in $G$. We study the maximum matching problem in fully dynamic graphs. In this problem, the input graph $G$ changes over time via a sequence of edge insertions and deletions. The goal is to maintain an (approximate) maximum matching of $G$ at all times, without spending too much time after each update.
Background on Dynamic Matching:
The dynamic matching problem has received a lot of attention over the last two decades [33, 4, 5, 32, 27, 18, 11, 12, 17, 16, 35, 21, 1, 13, 9, 10, 37, 14, 15, 34, 29, 26, 8, 20, 7, 19]. There is a relatively simple algorithm that maintains a $(1-\varepsilon)$-approximate maximum matching, for any fixed $\varepsilon>0$, in just $O(n)$ time per update (see e.g. [27]). The update-time can be significantly improved if we worsen the approximation. For instance, a 1/2-approximation can be maintained in $\operatorname{poly}\log n$ time per update [36, 4, 9], or an (almost) 2/3-approximation can be maintained in $O(\sqrt{n})$ time [12]. Nonetheless, when it comes to algorithms with approximation ratio better than $2/3$, the update-time stays close to $n$. A slightly sublinear algorithm was proposed by Assadi, Behnezhad, Khanna, and Li [3] which runs in $n/(\log^{*}n)^{\Omega(1)}$ time per update and maintains a $(1-o(1))$-approximation. In a very recent paper, Liu [30] improved this to $n/2^{\Omega(\sqrt{\log n})}$ via a nice connection to algorithms for the online matrix-vector multiplication (OMv) problem. Despite this progress, the following remains a major open problem:
Open Problem 1.
Is it possible to maintain a $(1-\varepsilon)$-approximte maximum matching in a fully dynamic graph, for any fixed $\varepsilon>0$, in $n^{1-\Omega(1)}$ update-time?
We note that there is an orthogonal line of work on fully dynamic algorithms that instead of maintaining the edges of the matching, maintain only its size [7, 20, 19]. For this easier version of the problem, Bhattacharya, Kiss, and Saranurak [19] positively resolved the open problem above. However, their algorithm crucially relies on only estimating the size and does not work for the problem of maintaining the edges of the matching. We refer interested readers to [7] where the difference between the two versions of the problem is mentioned.
1.1 Our Contributions
Contribution 1: Dynamic Matching.
In this paper, we make progress towards 1 by presenting a new algorithm whose update time depends on the density of a certain class of graphs that we call Ordered Ruzsa-Szemerédi (ORS) graphs, a generalization of the well-known Ruzsa-Szemerédi (RS) graphs.
Let us start by defining RS graphs.
Definition 1 (Ruzsa-Szemerédi Graphs).
An $n$-vertex graph $G=(V,E)$ is an $\mathsf{RS}_{n}(r,t)$ graph if its edge-set $E$ can be decomposed into $t$ edge-disjoint induced matchings each of size $r$. We use $\mathsf{RS}_{n}(r)$ to denote the maximum $t$ for which $\mathsf{RS}_{n}(r,t)$ graphs exists.
Instead of each matching being an induced matching in the whole graph, the edges of an ORS graph should be decomposed into an ordered list of matchings such that each matching is induced only with respect to the previous matchings in the ordering. The following formalizes this.
Definition 2 (Ordered Ruzsa-Szemerédi Graphs).
An $n$-vertex graph $G=(V,E)$ is an $\mathsf{ORS}_{n}(r,t)$ graph if its edge-set $E$ can be decomposed into an ordered list of $t$ edge-disjoint matchings $M_{1},\ldots,M_{t}$ each of size $r$ such that for every $i\in[t]$, matching $M_{i}$ is an induced matching in $M_{1}\cup\ldots\cup M_{i}$. We use $\mathsf{ORS}_{n}(r)$ to denote the maximum $t$ for which $\mathsf{ORS}_{n}(r,t)$ graphs exists.
Note that every $\mathsf{RS}_{n}(r,t)$ graph is an $\mathsf{ORS}_{n}(r,t)$ graph but the reverse is not necessarily true.
Our main result can now be stated as follows:
{thm} []
Theorem 1.
Let $\varepsilon>0$ be fixed. There is a fully dynamic algorithm that maintains the edges of a $(1-\varepsilon)$-approximate maximum matching in $O\left(\sqrt{n^{1+\varepsilon}\cdot\mathsf{ORS}_{n}(\Theta(\varepsilon^{2}n))}\operatorname{poly}(\log n)\right)$ amortized update-time. The algorithm is randomized but works against adaptive adversaries.
To understand the update-time in Theorem 1, we need to understand the density of ORS graphs for linear size matchings. Let $50<c<1/$ 5 be a constant. Since $\mathsf{ORS}_{n}(cn)\geq\mathsf{RS}_{n}(cn)$ as every RS graph is also an ORS graph with the same parameters, it is natural to first look into the more well-studied case of RS graphs. In other words, how dense can RS graphs with linear size matchings be? We note that this question has been of interest to various communities from property testing [22] to streaming algorithms [25, 3, 2] to additive combinatorics [24]. Despite this, the value of $\mathsf{RS}_{n}(cn)$ remains widely unknown.
The best lower bound on $\mathsf{RS}_{n}(cn)$— i.e., the densest known construction—is that of [22] from more than two decades ago which shows $\mathsf{RS}_{n}(cn)\geq n^{\Omega_{c}(1/\log\log n)}=n^{o(1)}$. This is indeed the densest known construction of ORS graphs we are aware of too. If this turns out to be the right bound, Theorem 1 implies a $(1-\varepsilon)$-approximation in $\widetilde{O}(n^{1/2+\varepsilon})$ time, an almost quadratic improvement over prior near-linear in $n$ algorithms of [3, 30]. In fact, we note that so long as $\mathsf{ORS}_{n}(cn)$ is moderately smaller than $n$ (say $\mathsf{ORS}_{n}(\Theta(\varepsilon^{2}n))\ll n^{1-2\varepsilon}$) Theorem 1 implies a truly sublinear in $n$ update-time algorithm, positively resolving 1.
Finally, we note that there is a long body of work on proving conditional lower bounds for dynamic problems. For instance, the OMv conjecture can be used to prove that maintaining an exact maximum matching requires near-linear in $n$ update-time [28]. Adapting these lower bounds to the $(1-\varepsilon)$-approximate maximum matching problem has remained open since then. Our Theorem 1 implies that proving such lower bounds either requires a strong lower bound of near-linear on $\mathsf{ORS}_{n}(\varepsilon n)$, or requires a conjecture that implies this.
Contribution 2: Better Upper Bounds for ORS and RS:
Unfortunately there is a huge gap between existing lower and upper bounds for $\mathsf{RS}_{n}(cn)$ (and as a result also for $\mathsf{ORS}_{n}(cn)$). To the best of our knowledge, the best known upper bound on $\mathsf{RS}_{n}(cn)$ follows from the improved triangle-removal lemma due to Fox [23] from more than a decade ago which shows $\mathsf{RS}_{n}(cn)\leq n/2^{\Omega(\log^{*}n)}=n^{1-o(1)}$ (see also [24, 2]).
Our second main result in this paper is the following upper bound.
Theorem 2.
For any $c>0$, it holds that $\mathsf{ORS}_{n}(cn)=O(n/\log^{(\ell)}n)$ for some $\ell=\operatorname{poly}(1/c)$, where $\log^{(x)}$ is the iterated log function.
Since every RS graph is also an ORS graph with the same parameters, Theorem 2 immediately implies the following bound on the density of RS graphs:
Corollary 3.
For any $c>0$, it holds that $\mathsf{RS}_{n}(cn)=O(n/\log^{(\ell)}n)$ for some $\ell=\operatorname{poly}(1/c)$.
Note that $\log^{(\ell)}n\gg 2^{\Omega(\log^{*}n)}$ for any fixed $c$. As such, Corollary 3 also improves the upper bound for RS graphs after more than a decade since the work of [23].
1.2 Perspective: ORS vs RS
Summarizing the above-mentioned bounds, the state-of-the-art bounds for the values of $\mathsf{RS}_{n}(cn)$ and $\mathsf{ORS}_{n}(cn)$ can be summarized as
[…]
While the value of $\mathsf{RS}_{n}(cn)$ remains widely unknown, one might argue that $\mathsf{RS}_{n}(cn)=n^{o(1)}$ is a plausible outcome, given that the construction of [22] has resisted any improvements for over two decades despite significant interest. But should we believe that $\mathsf{ORS}_{n}(cn)$ is also small in this case? Unfortunately the authors could not prove any formal relation between the densities of RS and ORS graphs beyond the upper bound of Theorem 2. In particular, we believe the following is an extremely interesting question for future work:
Open Problem 2.
Does it hold that $\mathsf{ORS}_{n}(cn)=\operatorname{poly}(\mathsf{RS}_{n}(O(cn)))$?
In the event that the answer to 2 is positive and $\mathsf{RS}_{n}(cn)=n^{o(1)}$ for any fixed $c>0$, we also get that $\mathsf{ORS}_{n}(cn)=n^{o(1)}$. Therefore Theorem 1 would imply an $n^{1/2+O(\varepsilon)}$ time algorithm in this case.
In the event that the answer to 2 is negative, one might wonder whether we can improve Theorem 1 by parametrizing it based on $\mathsf{RS}$ instead of $\mathsf{ORS}$. Put differently, suppose that $\mathsf{ORS}_{n}(cn)=n^{1-o(1)}$ and $\mathsf{RS}_{n}(cn)=n^{o(1)}$. Can we somehow utilize the sparsity of RS graphs (instead of ORS graphs) in this case to improve existing dynamic matching algorithms? We start Section 2 by providing an input construction which informally shows ORS is the right parameter for Theorem 1 even if RS is much sparser.
2 Our Techniques
Before describing the intuition behind our algorithm of Theorem 1, let us start with a sequence of updates that, in a sense, explains why existence of dense ORS graphs would make it challenging to maintain a $(1-\varepsilon)$-approximate maximum matching in a fully dynamic setting. We then proceed to show that this input construction is essentially the only instance that we do not know how to solve fast.
An input construction based on ORS graphs:
Consider a fully dynamic input graph $G=(V,E)$ that is composed of two types of vertices: the ORS vertices $V_{ORS}\subseteq V$ which is a subset of $n$ vertices, and the singleton vertices $V_{S}$ which is a subset of $(1-2\varepsilon)n$ vertices. We start by inserting an ORS graph in the induced subgraph $G[V_{ORS}]$. Namely, take an $\mathsf{ORS}_{n}(\varepsilon n,t)$ graph on $n$ vertices. We make the induced subgraph $G[V_{ORS}]$ isomorphic to this ORS graph by inserting its edges one by one to $G[V_{ORS}]$. Let $M_{1},\ldots,M_{t}$ be the ordered induced matchings of $G[V_{ORS}]$ as defined in Definition 2. Then the sequence of updates is as follows:
  • For $i=t$ to $1$:
    • Delete all existing edges of $V_{S}$.
    • If $i\not=t$, delete the edges of $M_{i+1}$.
    • Let $M_{i}$ be the current induced matching in $G[V_{ORS}]$.
    • Insert a perfect matching between the $(1-2\varepsilon)n$ vertices of $V_{ORS}$ left unmatched by $M_{i}$ and the $(1-2\varepsilon)n$ vertices in $V_{S}$.
Take the graph after the iteration $i$ of the for loop. Note that there is a perfect matching in $G$: match all singleton vertices to the $V_{ORS}$ vertices not matched by $M_{i}$, and match the rest of the vertices in $V_{ORS}$ through $M_{i}$. Importantly, since all matchings $M_{i+1},\ldots,M_{t}$ have already been deleted from the graph, $M_{i}$ must be an induced matching of the remaining graph $G$ (the other matchings $M_{1},\ldots,M_{i-1}$ cannot have any edge with both endpoints matched by $M_{i}$ due to Definition 2). Because of this, it can be confirmed that any $(1-\varepsilon/2)$-approximate maximum matching of $G$ must include at least half of the edges of $M_{i}$. The naive algorithm for finding an edge of $M_{i}$ for some vertex $v$ would scan the neighbors of $v$, which could take $\Omega(t)$ time per vertex (as this is the degrees in the ORS graph) and thus $nt$ time in total after every iteration of the loop consisting of $n$ updates. Hence, the amortized update-time of this algorithm must be at least $\Omega(nt/n)=\Omega(\mathsf{ORS}_{n}(\varepsilon n))$.
The input construction above implies that to maintain a $(1-\varepsilon)$-approximation of maximum matching, either we have to find a way to identify induced matchings of an ORS graph fast (without scanning the neighbors of each vertex) or have to parameterize our algorithm’s update-time by $t$, the density of ORS graphs.
Overview of our algorithm for Theorem 1:
Let us for this informal overview of our algorithm assume that the maximum matching size is at least $\Omega(n)$. Having this assumption allows us to find the matching once, do nothing for the next $\varepsilon n$ updates, and then repeat without hurting the size of the approximate matching that we find by more than a $1+O(\varepsilon)$ factor.
As it is standard by now, to find a $(1-\varepsilon)$-approximate matching it suffices to design an algorithm that given a subset $U\subseteq V$, finds a constant approximate maximum matching in $G[U]$. If this algorithm runs in $T$ time, we can find a $(1-\varepsilon)$-approximate maximum matching of the whole graph also in $O_{\varepsilon}(T)$ time. Amortized over $\varepsilon n$ updates, this runs in $O(T/n)$ total time for constant $\varepsilon>0$. However, just like the challenging example discussed above, in case the maximum matching in the induced subgraph $G[U]$ is an induced matching, we do not know how to find a constant fraction of its edges without spending $\Omega(n^{2})$ time. However, if we manage to bound the total number of such hard subsets $U$ for which we spend a lot of time, then we can bound the update-time of our algorithm. Intuitively, we would like to guarantee that if our algorithm takes $\Omega(n^{2})$ time to solve an instance $G[U]$, then the maximum matching in $G[U]$ must be an induced matching of the graph $G$, and charge these heavy computations to ORS which provides an upper bound on the number of edge-disjoint such induced matchings. However, there are two main problems: (1) it may be that the maximum matching in $G[U]$ is not an induced matching of the graph, yet it is sparse enough that it is hard to find; (2) even if $G[U]$ forms an induced matching, we have to ensure that its edges do not belong to previous induced matchings that we have charged, as ORS only bounds the number of edge-disjoint ordered induced matchings.
For the first problem discussed above, we present an algorithm that runs in (essentially) $O(n^{2}/d)$ time to find the maximum matching in $G[U]$. Here $d$ is a parameter that depends on the structure of $G[U]$ that measures how easy it is to find an approximate maximum matching of $G[U]$. Intuitively, if the average degree within $G[U]$ is $d$, we can random sample pairs of vertices to add to the matching. If each vertex is adjacent to $d$ others, we only need $O(n/d)$ samples to match it and the algorithm runs in $O(n^{2}/d)$ time. Of course, this can take up to $\Omega(n^{2})$ time if $G[U]$ is sparse – e.g. when it is an induced matching. Then instead of charging a matching in $G[U]$ that is an induced matching, we charge this matching of average degree at most $d$ inside. Let $M_{1},\ldots,M_{t}$ be the matchings that we charge and let $d_{i}$ be the average degree of the $i$-th matching and suppose that these matchings are edge-disjoint. We show in our update-time analysis that $\sum_{i=1}^{t}1/d_{i}$ can be at most $\mathsf{ORS}(\varepsilon^{2}n)$, and therefore the total time spent by our algorithm can be upper bounded by $n^{2}\mathsf{ORS}(\varepsilon^{2}n)$.
For the second problem, or in other words, to ensure that the matchings that we charge are edge-disjoint, we maintain a set $S$ and add all edges of any matching that we charge to this set. Thereafter, before solving $G[U]$, we first go over the edges stored in this set and see whether they can be used to find a large matching in $G[U]$. If they do, we do not run the random sampling algorithm discussed above. If not, the matching that we find must be edge-disjoint. We have to be careful that we do not make $S$ too dense though as we spend linear time in the size of $S$. Our final algorithm resets $S$ after a certain number of updates.
Overview of our upper bound in Theorem 2:
To obtain our upper bound of Theorem 2, we partition the matchings into two subsets, $\mathcal{M}$ and $\mathcal{M}^{\prime}$, based on their order in the sequence. A key insight is the following: take a vertex $v$ and suppose that it is matched by some matching $M$ in $\mathcal{M}^{\prime}$. If we remove all neighbors of vertex $v$ in $\mathcal{M}$, then it can be proved that no vertex of matching $M$ is removed because otherwise there must be an edge from $\mathcal{M}$ that matches two vertices of $M$, violating the inducedness property. Intuitively, this shows that if vertices have large degrees in $\mathcal{M}$, we can remove a relatively large number of vertices without hurting the matchings that include this vertex. To derive the upper bound, we carefully select a set of pivots based on the degrees in $\mathcal{M}$, and remove the neighbors of these pivots. We show that this reduces the number of vertices significantly enough, and keeps the size of a small (but sufficiently many) of the matchings unchanged. If the initial number of matchings is so large, we show that we can iteratively applying this procedure. Because the size of matchings do not change but the number of vertices drops, we get that the process should eventually stop. This implies the upper bound on the number of matchings in the starting graph.
3 Preliminaries
A fully dynamic graph $G=(V,E)$ is a graph defined on a fixed vertex set $V$ that is subject to edge insertions and deletions. We assume that each edge update is issued by an adaptive adversary that can see our algorithm’s previous outputs and can accordingly decide on the next update. We say an algorithm has amortized update-time $U$ if the total time spent after $T$ updates is $U\cdot T$ for some sufficiently large $T=\operatorname{poly}(n)$.
Tools from prior work:
Here we list some of the tools we use from prior work in our result.
The following proposition, implied by the streaming algorithm of McGregor [31] (see also [19] for its dynamic adaptation), shows that to find a $(1-\varepsilon)$-approximate matching, it suffices to solve a certain induced matching problem a constant number of times.
Proposition 4 ([31]).
Let $G=(V,E)$ be any (possibly non-bipartite) $n$-vertex graph. Suppose that for some parameters $\varepsilon\in(0,1)$ we have an algorithm $\mathcal{A}$ that provided any vertex subset $U\subseteq V$ with $\mu(G[U])\geq\varepsilon n$, finds a matching of size at least $\varepsilon\mu(G[U])$ in $G[U]$. Then there is an algorithm that finds a matching of size at least $\mu(G)-O(\varepsilon n)$ in $G$ by making $\widetilde{O}_{\varepsilon}(1)$ adaptive calls to $\mathcal{A}$ where preparing the subset $U$ for each call can be done in $\widetilde{O}_{\varepsilon}(n)$ time.
We also use the following sublinear-time algorithm of Behnezhad [6] for estimating the size of maximum matching. We note that even though we use this algorithm, we maintain the edges of the maximum matching explicitly.
Proposition 5 ([6]).
Let $G=(V,E)$ be any (possibly non-bipartite) $n$-vertex graph. For any $\varepsilon>0$, there is an algorithm that makes $\widetilde{O}(n\operatorname{poly}(1/\varepsilon))$ adjacency matrix queries to $G$ and provides an estimate $\widetilde{\mu}$ of the size of maximum matching $\mu(G)$ in $G$ such that with probability $1-1/\operatorname{poly}(n)$, it holds that
$0.5\mu(G)-\varepsilon n\leq\widetilde{\mu}\leq\mu(G).$
Finally, we use the following vertex sparsification idea which was adapted to the dynmaic setting in the work of Kiss [29].
Definition 6.
Let $V$ be a vertex set and denote $k\leq|V|/2$ be a parameter. A set $\mathcal{P}=\{P_{1},\ldots,P_{\ell}\}$ of $\ell$ partitionings of $V$ is a $(\ell,k,\varepsilon)$-vertex sparsifier if the following hold:
  1. 1.
    Each partitioning $P_{i}$ partitions $V$ into $O(k/\varepsilon^{2})$ subsets.
  2. 2.
    For each subset $S\subseteq V$ of size $2k$, there is at least one partitioning $P_{i}$ such that there are at least $(1-\varepsilon)|S|$ subsets in $P_{i}$ that have exactly one vertex of $S$.
Proposition 7 ([29]).
Let $\mathcal{P}$ include $\ell=\lceil 512\log n/\varepsilon^{2}\rceil$ random partitionings $P_{1},\ldots,P_{\ell}$ of $V$, where each $P_{i}$ partitions $V$ into at least $8k/\varepsilon$ subsets randomly. Then, with probability $1-1/\operatorname{poly}(n)$, $\mathcal{P}$ is a $(\ell,k,\varepsilon)$-vertex sparsifier of $V$.
The notion of graph contraction here refers to the process of simplifying the graph based on the partitionings provided by the vertex sparsifiers. Each set of partitions effectively represents a contracted version of the graph, where the vertices within the same subset can be considered as a single entity for the purpose of matching. This contraction not only reduces the overall size of the graph but also retains the essential structural properties that influence the size and composition of maximum matchings.
Definition 8.
Denote $M$ be a given matching with $2r$ vertices in graph $G$. Let us assume the $d_{v}$ for a vertex $v$ in the matching is the number of neighbors of $v$ in $V(M)$. we define average degree of $M$ as $\sum_{v\in V(M)}d_{v}/2r$.
4 Dynamic Approximate Matching
We present our dynamic matching algorithm in this section and prove Theorem 1.
Let us start with an overview of how our proposed algorithm is built. We randomly sample edges from graph $G$ to find a matching with $\varepsilon$-approximation, which is described in Section 4.1. We propose that if we need many samples to find this matching then the average degree of this matching is small in the remaining edges (what we mean by the remaining edges is that, when we output a matching, the adversary can remove its edges from the graph, while searching for a new matching we should discard the edges found before), resulting in the matching to be induced in the remaining graph. In this case, we charge the cost of the algorithm to ORS, with characteristics similar to the matching we found. Otherwise, we will succeed in finding a matching, relatively fast.
Note that, in this model, we can argue that if the matching size in the graph is $\mu$ then the matching size will not change after $\varepsilon^{2}n$ updates. We will find a matching once in $\varepsilon^{2}n$ edge updates to optimize amortized running time. This motivates us to run Section 4.1 to find a matching after $\varepsilon^{2}n$ updates. To keep track of the large matchings we already found we introduce Section 4.2. This algorithm simplifies our problem where the matching size does not change. If we maintain this algorithm for different matching sizes, knowing the approximate matching we will know which maintained algorithm to look at to find the $\varepsilon$-approximate matching.
To boost our approximation to $(1-\varepsilon)$ we use Proposition 4 to output the subsets of $U$ that we need to search for the matchings of size $\varepsilon\mu$. We use the subsets in Section 4.1 which is a $\varepsilon$-approximation algorithm.
Note that we provide a multiplicative approximation maximum matching algorithm for the dynamic model, which leads to the fact that we need to handle the cases where the matching size is not of order $n$. This adds some layers to the intuition we stated above.
  1. 1.
    If the matching size is $n^{\prime}$ we output a matching after $\varepsilon^{2}n^{\prime}$ updates. This means we need to handle matching with similar sizes together, we will have $\log n$ partition to keep track of the matchings outputted by Section 4.1 of almost the same size simultaneously.
  2. 2.
    To handle the graph with small matchings of $O(n^{\prime})$ we need to contract the graph to one with $n^{\prime}$ vertices. We build a contracted graph by creating $n^{\prime}$ buckets of $O(n/n^{\prime})$ vertices, the vertices are distributed randomly in buckets. Since we want to preserve the edges of the maximal matching we need to randomly maintain $O(\log n)$ such contractions for $\log n$ values of $n^{\prime}$.
  3. 3.
    The fact that the edges of the matching found in the earlier steps of Section 4.2, could be removed by the adversary leads us to a definition of the degree of matching that only takes into account the matchings that we will be finding in the next rounds (we do not care about the edges removed from the graph in the dynamic steps). This concept leads us to define the set of matchings named ORS, where every matching $M$ is induced in the set of matchings we observe after $M$.
4.1 Random Sampling Algorithm
Here we introduce a sampling algorithm that given a graph $G$, with a matching of size $\mu$ can output a matching of size $\varepsilon\mu$. We bound the algorithm running time with the average degree of the outputted matching. The way the algorithm works is that for each vertex we sample $b$ vertices and check if the edge is in the graph using the adjacency matrix of the graph. We add the edges to the matching if none of the endpoints have been used previously. Note that by starting with a small $b$ for the budget, we ensure to find the edges from dense parts of the graph which is easier to find. As we go further in the algorithm we make the budget larger to find edges in more sparse parts of the graph. Now, we state the random sampling algorithm and prove its efficiency.
Lemma 9.
Given a graph $G=(V,E)$, vertex subset $U\subseteq V$, and parameters $\varepsilon\in(0,1)$, such that maximal matching of $G[U]$ is $\mu\geq\varepsilon n$, there is an algorithm that finds a matching $M\subseteq G[U]$ of size at least $\varepsilon\mu$. The algorithm runs in time $O(n^{2+\varepsilon}/d)$ where $d=|E\cap V^{2}(M)|/O(n)$.
Proof.
Algorithm 1: RandomSamplingAlgorithm$(G[U],\varepsilon,\mu)$   Parameters: $\varepsilon\in(0,1]$. $M\leftarrow\emptyset$. $M_{1},...,M_{1/(2\varepsilon)}\leftarrow\emptyset$. $S\leftarrow\emptyset$ $b\leftarrow 1$. $c=0$. while $|M|<\varepsilon\mu$ do 1       if $b>\varepsilon c\cdot\log n$ then 2             $M_{c+1}\leftarrow M$. $M\leftarrow\emptyset$. $c\leftarrow c+1$. 3      for vertex $v\in U\setminus V(M)$ in random order do 4             $S\leftarrow$ Sample $b$ vertices in $U$. if $\exists u\in S$ such that $u\in U$ and $u\notin V(M)$ and $u$ is a neighbor of $v$ then 5                   $M\leftarrow(v,u)\cup M$. 6             7      $b\leftarrow 2b$ return $M$
Let us consider Section 4.1, First we prove that this algorithm terminates. We need to prove that there exists $c$ such that sampling vertices with a budget $b$ in the interval of $[\varepsilon c\cdot\log n,\varepsilon(c+1)\cdot\log n]$ outputs a matching of size at least $\varepsilon\mu$. Let us assume there are no such intervals. This means that for any integer $c\in[0,1/(2\varepsilon)]$, running the algorithm for $b$ in the interval $[\varepsilon c\cdot\log n,\varepsilon(c+1)\cdot\log n]$ outputs a matching with less than $\varepsilon\mu$ edges. Note that for each edge $e$ in $M_{1}\cup\dots\cup M_{1/(2\varepsilon)}$, at most 2 edges in the certificate matching $\mu$ have common vertices with $e$. Summing over all the $1/(2\varepsilon)$ intervals we have:
$2\sum_{c=1}^{1/(2\varepsilon)}{|M_{i}|}<2\cdot\frac{1}{2\varepsilon}\cdot\varepsilon\mu=\mu.$
The fact that we could not find a matching of size $\mu$ is a contradiction with the lemma statement ensuring the existence of such matching.Therefore, we have at least the size of one of the matchings $M_{1},...,M_{1/(2\varepsilon)}$ is $\varepsilon\mu$.
Now, we prove the statement of the algorithm. We assume that this algorithm terminates at $\varepsilon(c-1)\cdot\log n\leq b<\varepsilon c\cdot\log n$, for a constant $\delta$, we spend $n[b+b/2+..+1]<\delta nb\leq\delta\varepsilon c\cdot n\log n<\delta\varepsilon cn^{1+\varepsilon}$ time to find the matching $M_{c}$ such that $|M_{c}|\geq\varepsilon\mu$. Note that if it takes at most $n^{2+\varepsilon}/d$ time to find the matching of size $\varepsilon n$, then we have $n^{2+\varepsilon}/d<\delta\varepsilon cn^{1+\varepsilon}$ therefore, $d>\frac{n}{\delta\varepsilon c}$.
For budget $b$, we sample vertices for all the vertices, we argue that for the vertices that are not in $M$, and after sampling $b$ vertices we could not find their neighbor then the degree of vertices out of the matching is less than $n/b$. Assuming that the degree of vertex $v$ is $d_{v}$, with probability $(1-\frac{d_{v}}{n})^{b}$ we sampled a neighbor, and this probability converges to $1/e$ for $d_{v}>O(n/b)$. Therefore in this case with a constant probability we already matched $v$. This implies that if we find the first edge of this matching at a budget of $2b$, we get that the degree of any vertex out of the matching we found until this point is less than $O(n/b)$. Therefore, the total sum of the degree of the matching is less than $O(\varepsilon\mu\cdot n/b)<O(\varepsilon\mu n/((c-1)\log n))<O(\delta\varepsilon^{2}cdn/((c-1)\log n))<O(nd)$. ∎
4.2 The Overall Algorithm via Certificates and Caching Them
In the next step, we aim to use Section 4.1 efficiently to output an approximate matching in the dynamic model. First, we prove that given the matching size is equal to $O(n^{\prime})$, after $\varepsilon^{2}n^{\prime}$ updates the matching size does not change substantially. This means we can output a matching after each $\varepsilon n^{\prime}$ update and get an $\varepsilon$-approximate matching.
Observation 10.
Given a graph $G=(V,E)$ with $n$ vertices, if the maximum matching size is $O(\mu)$ and we provide an $\varepsilon$-approximate matching $M^{\prime}$, after $\varepsilon^{2}\mu$ updates in the dynamic model, $M^{\prime}$ or one of its subsets remains an $\varepsilon$-approximate matching for the updated graph.
Proof.
We assume that in a graph $G$ the maximal matching size is $c\mu$ where $c$ is a constant, after $\varepsilon^{2}\mu$ updates in the graph the size of the maximal matching is between $(c-\varepsilon^{2})\mu,(c+\varepsilon^{2})\mu$. This implies that if we have an algorithm that outputs an $\varepsilon c\mu$ matching, after $\varepsilon^{2}\mu$ updates, the $\varepsilon$ approximate matching is preserved. This is because we lose at most $\varepsilon^{2}$ fraction of the outputted matching. ∎
Note that to use Section 4.1 we need to know about the existence of a matching $\mu$, the algorithm described in Proposition 5, given a graph with $n$ vertices, outputs if there is matching of size $O(n)$. Using this algorithm alone will not help us with the cases where the matching size is not of order $n$. This motivates us to find an algorithm that reduces the number of vertices on the graph while maintaining the maximum matching size. Proposition 7 will maintain $O(\log n)$ contraction on the graph for an integer $k\in[1,\log n]$, such that for any matching $M$ with high probability, there is a contraction preserving a $1-\varepsilon$ approximation of $M$. After running the reduction on the graph, we find the contraction that preserves the maximum matching. Let us assume that $G^{\prime}$ is the contracted graph, we run Section 4.1 on $G^{\prime}$ to output the $\varepsilon$ approximation matching.
Note that throughout the changes in the dynamic model, both the matching size and the contracted graph that preserves the maximum matching, change. An edge insertion in $G$ is a weight addition to an edge in graph $G^{\prime}$, and edge deletion is a weight subtraction in $G^{\prime}$. In the following algorithms, when we find a matching in $G^{\prime}$, we do not care about the weight of edges, an edge in $G^{\prime}$ exists if its weight is more than 0, otherwise, This edge does not exist in $G^{\prime}$.
First, we solve a simplified version of the problem where the matching size does not change throughout the updates, and a single contracted graph $G^{\prime}$ preserves matching edges in the dynamic model. We prove that there exists an efficient algorithm maintaining a $(1-\varepsilon)$-approximation matching in the dynamic model.
We introduce Section 4.2. We assume that the updates in this algorithm are in the case that the matching size in $O(n^{\prime})$, and $G^{\prime}$ preserves the maximum matching of the graph. This algorithm works in phases, each phase represents the number of updates we want to use the matchings we found in the last updates.
Algorithm 2: Algorithm for Dynamic Model in Contracted Graphs   Parameters: $\varepsilon\in(0,1]$, $t$ the threshold of the algorithm. $G^{\prime}\leftarrow$ the contracted graph. $S_{1}\leftarrow\emptyset$. $S_{2}\leftarrow\emptyset$. while  edge updates $<t$  do 1       if There are $\varepsilon^{2}n^{\prime}$new updates then 2             $S_{1}\leftarrow S_{1}\cup$ Insertions. $M\leftarrow\emptyset$. while Proposition 4 outputs a subset $U$ that has a matching of size at least $\mu=\varepsilon n^{\prime}$  do 3                   $M^{\prime}\leftarrow$ a maximal matching found by a greedy algorithm on edges of $S_{1}\cup S_{2}$ in $U$. For each edge in this set, we check if it exists in the dynamic graph. if  $|M^{\prime}|>\varepsilon\mu$  then 4                        $M\leftarrow M^{\prime}$. 5                  else 6                         $M_{U}\leftarrow$ $RandomSamplingAlgorithm(G[U\setminus V(M^{\prime})],\frac{\varepsilon}{1-2\varepsilon},(1-2\varepsilon)\mu).$ $M\leftarrow M\cup M_{U}$. 7                  $S_{2}\leftarrow S_{2}\cup M_{U}.$ Output $M$. 8             9
First, we prove that this algorithm obtains a $\varepsilon$-approximation matching.
Lemma 11.
Given a contracted graph $G^{\prime}$ with $n^{\prime}$ vertices that the maximum matching size in $G^{\prime}$ is $O(n^{\prime})$ in $t$ updates, Section 4.2 maintains an $(1-\varepsilon$)-approximate maximum matching for $G^{\prime}$ in the dynamic model.
Proof.
Note that we proved in Observation 10 that if the maximal matching size is of order $n^{\prime}$, then we do not need to output a matching in $\varepsilon^{2}n^{\prime}$ updates.
By Section 4.2, set $S_{1}$ contains edge insertions, and set $S_{2}$ contains all the matchings we found earlier in the algorithm.
Using Proposition 4, we are given constant subsets of vertices that contain matchings of size at least $\varepsilon n$, we can run Section 4.1 on these subsets to output a $\varepsilon$-approximate matching in given subset $U$, if there are no large enough activated matchings in set $S_{2}\cup S_{1}$. Given that we aim to find edge-disjoint matchings, we need to find a maximal matching in $S_{2}\cup S_{1}$ and remove its endpoints from $U$.
After finding a matching $M^{\prime}$ in $S_{2}\cup S_{1}$ of size $\varepsilon\mu$, we output this matching; otherwise, we deactivate all the endpoints in $M^{\prime}$. Note that we are certificated by Proposition 4 to have a matching of size $\mu$ in $U$. Since for maximal matching $M^{\prime}$ in $S_{1}\cup S_{2}$ we have $|M^{\prime}|<\varepsilon\mu$, then the maximum matching size in $S_{1}\cup S_{2}$ is at most of size $2\varepsilon\mu$. This certifies the existence of a matching of size $(1-2\varepsilon)\mu$ in the rest of $U$. Now, we run Section 4.1 on graph $U$ with endpoints of $M^{\prime}$ removed with parameter $\varepsilon/(1-2\varepsilon)$. This implies that we get a matching of size $\varepsilon\mu$. Since we deactivate vertices of $S_{2}$ we only add edge-disjoint matchings to set $S_{2}$.
Note that any vertex in $G^{\prime}$ represents a bucket of vertices of $G$. By maintaining the number of vertices between two buckets in $G$, we can observe whether there is an edge between two vertices in $G^{\prime}$. This is how we can certify if a matching $M^{\prime}$ in $S_{2}$ is in $G^{\prime}$.
We proved that Section 4.2 outputs a $\varepsilon$-approximate matching $M_{U}$ in each subset of $U$ presented by Proposition 4. Note that given $M_{U}$ for each subset $U$, Proposition 4 finds a $(1-\varepsilon)$-approximation matching in $G^{\prime}$. This is true by the assumption that the maximum matching size of $G^{\prime}$ is $O(n^{\prime})$. ∎
Now that we introduced an algorithm for each contraction, we explore how to build and maintain each contraction. To partition updates to the contraction graphs that we maintain we need to run Proposition 5 to approximate the matching size, and then use Proposition 7 to figure out the contracted graph preserving the maximum matching, which we will do after each $\varepsilon^{2}n^{\prime}$ updates, given that the last contracted graph had $n^{\prime}$ vertices. We introduce Section 4.2 that given updates calculates the matching size and finds a suitable contraction for the graph at each state. For each update, Section 4.2 calls Section 4.2 on all the contractions, however, it only outputs the matching build by the matching preserving contraction.
We will run $O(\log^{2}n)$ parallel versions of Section 4.2 each for a maintained contracted graphs that we maintain. More formally we introduce Section 4.2:
Algorithm 3: Main   Parameters: $\varepsilon\in(0,1]$. $k\leftarrow\log n$. $n_{1},n_{2},\dots,n_{k}\leftarrow 2,4,\dots,2^{k}$. $t\leftarrow$ threshold of algorithm Section 4.2. $\mathcal{(}C)\leftarrow$ set of maintained $O(\log n)$ contracted graphs for each $n_{c}\in\{n_{1},\dots,n_{k}\}$ using Proposition 7. $\mathcal{P}\leftarrow$ set of maintained parallel edges between two contracted sets of vertices after each update. Run a new phase Section 4.2 for each contraction after $t$ updates. For each contraction $c$ with $c_{n}$ vertices, run Proposition 5 on $c$ after $\varepsilon^{2}n_{c}$ updates to see whether $c$ preserves the maximum matching in $G$. Find the contraction $c$ preserving the largest matching of $G$ and use the output of Section 4.2 for the corresponding contraction, to output an $(1-\varepsilon)$-approximate matching.
Now, we prove that this algorithm outputs a $(1-\varepsilon)$-approximation matching.
Lemma 12.
Given graph $G$ with $n$ vertices, Section 4.2 maintains an $(1-\varepsilon$)-approximate maximum matching for $G$ in the dynamic model.
Proof.
In this algorithm, we maintain $O(\log^{2}n)$ contraction of graph $G$, where $G$ is the updated graph in the current batch of $t$ updates. If the maximum matching size of $G$ is $\mu$, by Proposition 7, there exists a contraction $c$ with $n_{c}$ vertices that preserves the maximum matching with $1-\delta$ factors and $\mu=O(n_{c})$. Denote $\delta$ a parameter sufficiently smaller than $\varepsilon$. We call $c$ a preserving contraction of $G$. This means that by Lemma 11, Section 4.2 outputs a $(1-\delta)$-approximation matching for $G^{\prime}$, that is $(1-\varepsilon)$-approximation matching for $G$. Note that, by running Proposition 5 on the maintained reduction, we find the preserving contraction of $G$.This concludes the $(1-\varepsilon)$-approximate. ∎
4.3 Running Time Analysis
Now, we explore the time per update for Section 4.2.
Lemma 13.
If we set the threshold of Section 4.2 equal to $t=n^{1+\varepsilon/2}\sqrt{n\log n\cdot\mathsf{ORS}_{n}(\varepsilon\mu)}$, then the time spent on single contraction $G^{\prime}$ in both Section 4.2 is $O(\sqrt{n^{1+\varepsilon}\log n\cdot\mathsf{ORS}_{n}(\varepsilon\mu)})$ per each update, where $\mu$ is the maximum matching size in $G^{\prime}$.
Proof.
Fix a contracted graph $G^{\prime}=(V^{\prime},E^{\prime})$ with $n$ vertices. We explore the algorithm running time for this graph.
First, let us determine the time the algorithm spends finding the size of the certificate matching, $\mu$, after $\varepsilon^{2}n$ updates. Note that we run Proposition 5 after each $\varepsilon^{2}n$ vertex update, this algorithm outputs $\mu$, which is a constant approximation of the maximum matching size of the graph. Moreover, this algorithm certifies the existence of a matching of size $\mu$ in the graph, the running time for which is $\widetilde{O}(n\operatorname{poly}(1/\varepsilon))$.
By Proposition 4, after each $\varepsilon^{2}n$ update we spend $\widetilde{O}_{\varepsilon}(n)$ on finding the appropriate subsets of vertices such as $U$, including matching of size at least $\varepsilon n$.
Now, we calculate the time Section 4.2 takes going through the set $S_{1}$ and $S_{2}$, after $\varepsilon^{2}n$ updates to find a matching $M^{\prime}$. Note that, after $\varepsilon^{2}n$ updates we go over the set $S_{1}\cup S_{2}$ to find an active matching. We add an active edge to the matching if none of the endpoints was added before it. The algorithm outputs a maximal matching this way. The total running time per $\varepsilon^{2}n$ updates is $|S_{1}|+|S_{2}|$.
Finally, we explore the running time of Section 4.1. Let us assume that at the end of a phase of Section 4.2 we have $S_{2}=\{M_{1}\cup\dots\cup M_{l}\}$, We prove in Lemma 11 that these matchings are edge-disjoint. Note that the total edge-disjoint matchings in $G^{\prime}$ can be more than $l$ matching, $M_{1},\dots,M_{l}$ are all the matchings we find during the $t$ edge updates. We assume that the total edge-disjoint matchings in $G^{\prime}$ is $l_{t}$. We label these matchings as $M_{1},\dots,M_{l},M_{l+1},\dots,M_{l_{t}}$. Let us assume that for $i\in\{1,\dots,l\}$ we denote $d_{i}$ the average degree of matching $M_{i}$ in the edge set of $\{M_{i}\cup\dots\cup M_{l_{t}}\}$. Denoting by $M^{\prime}$ a maximal matching in $S_{1}\cup S_{2}$ found in Section 4.2, we prove that if Section 4.1 in graph $G^{\prime}\setminus V(M^{\prime})$, takes $O(n^{2+\varepsilon}/d)$ to output matching $M_{i}$ then we have $O(d)=d_{i}$. Note that by Lemma 9 we know that the average degree of matching $M_{i}$ in $G^{\prime}\setminus(S_{1}\cup S_{2})$ is $O(d)$. Note that, for matching $M_{1},\dots,M_{i-1}$, either their edges are removed from the graph or we deactivate the vertices that made a large matching. This implies that no edges from $M_{1},\dots,M_{i-1}$ remain in the graph. If we assumed that an edge $e$ was still in the graph with both endpoints not deactivated, we could add this edge to $M^{\prime}$, this is in contradiction with the fact that $M^{\prime}$ is a maximal matching. This implies that $d_{i}=O(d)$. This results in the conclusion that the running time of Section 4.1 is less than $n^{2+\varepsilon}/d_{i}$.
This definition of matching average degree motivates us to define ORS graphs. Recall that the ORS graph is a set of matching $M_{1},\dots M_{l}$ that each matching is induced in the set of matching with higher indices. We then upper bound the running time of our algorithm by the size of ORS graphs according to Lemma 14.
Now, let us compute the total running time per update. Note that since we add a matching of size $\varepsilon\mu$ to $S_{2}$ after $\varepsilon^{2}n$ updates, $|S_{2}|<|S_{1}|$. Let us assume that we add $M_{1},\dots M_{l}$ to $S_{2}$ by running Section 4.1. The total running time per update is:
$\frac{\frac{|S_{1}|}{\varepsilon^{2}n}\cdot(|S_{1}|+|S_{2}|+\widetilde{O}(n))+n^{2+\varepsilon}\cdot\sum_{i=1}^{l}1/d_{i}}{t}.$
Using Lemma 14 we have the running time as:
$\frac{\frac{|S_{1}|}{\varepsilon^{2}n}\cdot(|S_{1}|+|S_{2}|+\widetilde{O}(n))+n^{2+\varepsilon}\cdot O(\mathsf{ORS}_{n}(\varepsilon\mu))\log n}{t}.$
Now, if we set the threshold as $t=n^{1+\varepsilon/2}\sqrt{n\log n\cdot\mathsf{ORS}_{n}(\varepsilon\mu)}$ we have:
$|S_{2}|,|S_{1}|<t=n^{1+\varepsilon/2}\sqrt{n\log n\cdot\mathsf{ORS}_{n}(\varepsilon\mu)},$
This is because we could add a matching of size $\varepsilon\mu$ to $S_{2}$ after $\varepsilon^{2}n$ updates, and the total insertions are less than the total updates. The total running time per update is:
$O(\sqrt{n^{1+\varepsilon}\log n\cdot\mathsf{ORS}_{n}(\varepsilon\mu)}).\qed$
Lemma 14.
Given a set of matching $M_{1},\dots,M_{t}$ with distinct edges, of size at least $r$ in graph $G$, for any index $i\in[1,t]$ let us define $d_{i}$ as the average degree of $V(M_{i})$ in $M_{1}\cup\dots\cup M_{i}$. For any $\alpha\in(0,1)$, we have $\sum_{i=1}^{t}1/d_{i}=O(\mathsf{ORS}_{n}(r(1-\alpha))\cdot\log n)$.
Proof.
Note that the maximum average degree of a matching $M$ is $n-1$, and the minimum is 1. Therefore, if we partition the input matchings into $\log n=k$ groups, the average degree of any matching in each partition is at most within 2 factors of the minimum average degree.
We first prove that if we have $t^{\prime}$ matchings $M^{\prime}_{1},\dots,M^{\prime}_{t^{\prime}}$ a subset of $\{M_{1},\dots,M_{t}\}$, and their average degree as $d^{\prime}_{1}\star,\dots,d^{\prime}_{t^{\prime}}$ that for any $i\in[1,t^{\prime}]$, $1/d^{\prime}_{i}$ is in a $2$ factor of $\min_{i=1}^{t^{\prime}}1/d^{\prime}_{i}$. For any $\alpha\in(0,1)$, we have $\sum_{i=1}^{t^{\prime}}1/d^{\prime}_{i}=O(\mathsf{ORS}_{n}(r(1-\alpha)))$.
To prove this, we first sample some of the matchings and the edges that are not induced in proper subsets of matchings. We prove there are plenty number of matchings remaining that are induced within their corresponding set of matchings.
Let us fix a small variable $\delta<\varepsilon,\delta,\varepsilon\in(0,1)$. We set $H=M^{\prime}_{1}\cup...\cup M^{\prime}_{t^{\prime}}$.
  • Step 1. Let us define $x_{e}$ for any edge $e=(u,v)$ appearing in matching $M^{\prime}_{\alpha_{j}}$ as follows
    $\sum_{i\in\{j,j+1,\dots,t^{\prime}\};u,v\in V^{2}(M_{\alpha_{i}})}\frac{1}{d^{\prime}_{\alpha_{i}}}.$
    We define $E_{1}=\{e|x_{e}>1/\varepsilon\}$ We remove any matching from $H$ with more than $2\varepsilon r$ edges in $E_{1}$ Let us call the remaining set of matchings $H_{1}$.
  • Step 2. Let us pick a matching $M^{\prime}_{i}\in H$ with probability $\frac{\delta}{2d^{\prime}_{i}}$. Let us assume that at this step, we pick the following $k$ matchings $M^{\prime}_{\alpha_{1}},\dots,M^{\prime}_{\alpha_{k}}$. we define this subset of the graph as $H_{2}$.
  • Step 3. Now, we prone any matching $M^{\prime}_{i}$ chosen in $H_{1}$ and has less than $(1-\delta)cn$ edges induced in $(M^{\prime}_{1}\cup\dots\cup M^{\prime}_{i})\cap(H_{2})$. Let us call the remaining graph $H_{3}$.
To see that the claim is true, first, we need to prove that in Step 1 we remove at most $\varepsilon$ fraction of the edges in $H$. Second, we prove that $H_{2}$ is large enough.
Let us assume that after Step 1 we remove more than $\varepsilon rt$ of the edges in $H$ from the total $rt^{\prime}$ edges. Therefore, we have
$\sum_{e\in E_{1}}x_{e}>rt^{\prime}.$
To compute $\sum_{e\in H}x_{e}$, we notice how much a matching $M^{\prime}_{\alpha_{i}}$ contributes to the total sum. Note that since the average degree of $M^{\prime}_{\alpha_{i}}$ in $\{M^{\prime}_{\alpha_{1}},\dots,M^{\prime}_{\alpha_{i}}\}$ is at most $d^{\prime}_{\alpha_{i}}$ therefore, the total contribution of matching $M^{\prime}_{\alpha_{i}}$ is $r\cdot d^{\prime}_{\alpha_{i}}\cdot\frac{1}{d^{\prime}_{\alpha_{i}}}$. In conclusion, the total sum of $x_{e}$ for any edge $e\in H$ is at most
$\sum_{i\in\{1,\dots,t^{\prime}\}}d^{\prime}_{\alpha_{i}}\cdot r/d^{\prime}_{\alpha_{i}}=rt^{\prime},$
which is a contradiction suggesting that we remove at most $\varepsilon rt^{\prime}$ edges.
Now that we only removed at most $\varepsilon$ fraction of the edges, note that $H_{1}$ contains at least half of the matchings in $H$. This is because if more than $|H|/2$ matchings had more than $2\varepsilon r$ edges removed, then the total number of edges in $E_{1}$ would be more than $\varepsilon rt^{\prime}$ which is a contradiction.
We aim to prove that $H_{3}$ contains $\delta|H_{1}|/2$ matchings. Let us fix a matching $M\in H_{1}$. Note that for any remaining edge $e$ in $M$, we have $x_{e}<1/\varepsilon$. Therefore, the probability that $e$ belongs to another matching is less than $\delta/2\varepsilon$ since we picked each matching $M^{\prime}_{i}$ with probability $\delta/2d^{\prime}_{i}$. This implies that the expected number of edges in $E(^{\prime})\cap E(H_{3})$ is more than $(1-2\varepsilon)(1-\delta/2\varepsilon)r$. Now, by applying Markov’s inequality to the random variable $r(1-2\varepsilon)-|E(M)\cap E(H_{3})|$, here we assume that the maximum number of edges is $r(1-2\varepsilon)$, and we get the following.
$\Pr[r(1-2\varepsilon)-|E(M)\cap E(H_{3})|\geq(1-2\varepsilon)(\delta/\varepsilon)r|M\in H_{2}]\leq\frac{\operatorname{\textbf{E}}[r(1-2\varepsilon)-|E(M)\cap E(H_{3})|M\in H_{2}]}{(1-2\varepsilon)(\delta/\varepsilon)r}.$
Since $\operatorname{\textbf{E}}[r(1-2\varepsilon)-|E(M)\cap E(H_{3})||M\in H_{2}]\leq((1-2\varepsilon)\delta/2\varepsilon)r$ we have
$\Pr[r(1-2\varepsilon)-|E(M)\cap E(H_{3})|\geq(1-2\varepsilon)(\delta/\varepsilon)r|M\in H_{2}]\leq\frac{(\delta/2\varepsilon)r}{(\delta/\varepsilon)r}=1/2.$
Therefore, we have
$\Pr[r(1-2\varepsilon)-|E(M)\cap E(H_{3})|\leq(1-2\varepsilon)(\delta/\varepsilon)r|M\in H_{2}]>1/2,$
which means
$\Pr[|E(M)\cap E(H_{3})|\geq(1-2\varepsilon)(1-\delta/\varepsilon)r|M\in H_{2}]>1/2.$
Multiplying with the probability that $M\in H_{2}$, we get that the probability of matching to have more than $(1-2\varepsilon)(1-\delta/\varepsilon)r$ edges and be selected is more than $\delta/(4\varepsilon d^{\prime}_{i})$.
Now, summing over all matchings in $H_{1}$ we get
$\operatorname{\textbf{E}}[|H_{3}|]\geq(\delta/(4\varepsilon))\sum_{i,M^{\prime}_{i}\in H_{1}}1/d^{\prime}_{i}.$
Since $H_{3}$ is an ORS graph we get
$(4\varepsilon)\mathsf{ORS}_{n}((1-2\varepsilon)(1-\delta/\varepsilon)r)/\delta\geq\sum_{i,M^{\prime}_{i}\in H_{1}}1/d^{\prime}_{i}.$
Now, using that for any $j\in\{1,..,t^{\prime}\}$ we have $1/d^{\prime}_{j}\leq 2\min_{i=1}^{t^{\prime}}1/d^{\prime}_{i}$, for any $i,j\in\{1,...,t^{\prime}\}$ we get $1/d^{\prime}_{i}\leq 2/d^{\prime}_{j}$.Therefore,
$\sum_{i,M^{\prime}_{i}\in H_{1}}1/d^{\prime}_{i}\geq\left(\sum_{j,M^{\prime}_{j}\in H\setminus H_{1}}1/d^{\prime}_{j}\right)/2.$
Now, we give an upper bound for $\sum_{i,M_{i}\in H}1/d^{\prime}_{i}=\sum_{i,M^{\prime}_{i}\in H_{1}}1/d^{\prime}_{i}+\sum_{j,M^{\prime}_{j}\in H\setminus H_{1}}1/d^{\prime}_{j}$. We have the following.
$(12\varepsilon)\mathsf{ORS}_{n}((1-2\varepsilon)(1-\delta/\varepsilon)r)/\delta\geq 3\sum_{i,M^{\prime}_{i}\in H_{1}}1/d^{\prime}_{i}\geq\sum_{i,M^{\prime}_{i}\in H_{1}}1/d^{\prime}_{i}+\sum_{j,M^{\prime}_{j}\in H\setminus H_{1}}1/d^{\prime}_{j}.$
Setting $\delta=\varepsilon/2,\varepsilon=4\alpha/13$ we have:
$\sum_{i,M^{\prime}_{i}\in H}1/d^{\prime}_{i}=O(\mathsf{ORS}_{n}((1-\alpha)r)).$
Note that we showed an upper bound for the sum of the inverse average degree of each matching in one partition. Summing up the inverse average degree for all the partitions we get:
$\sum_{i=1}^{t}1/d_{i}=O(\mathsf{ORS}_{n}((1-\alpha)r)\cdot\log n.\qed$
Lemma 15.
If we set the threshold of Section 4.2 equal to $t=n^{1+\varepsilon/2}\sqrt{n\log n\cdot\mathsf{ORS}_{n}(\Theta(\varepsilon^{2}n)}$, then the total time per update is $O(\operatorname{poly}\log n)\cdot\sqrt{n^{1+\varepsilon}\log n\cdot\mathsf{ORS}_{n}(\Theta(\varepsilon^{2}n))}$.
Proof.
We run the $O(\log^{2}n)$ parallel version of Section 4.2, and maintaining the contracted graphs takes $\operatorname{poly}\log n$ time. We will maintain the number of parallel edges in $G^{\prime}$ between any two contracted vertex in $\operatorname{poly}\log n$ time. Denote the set of contractions as $\mathcal{C}$ in one phase of the algorithm.
Note that if the size of a contraction $c$ is $n_{c}$, and the maximum matching in $G$ is of order $n_{c}$ we run the algorithm for this contraction. Using Proposition 7 there is a reduction that preserves that matching in $G$ with a factor of $1-\varepsilon$. Now, integrating over the contractions in $\mathcal{C}$ we find the corresponding contraction by running Proposition 5. If the maximum matching in contraction $c$ is $M_{c}$, this algorithm certifies the existence of a matching of size $|M_{c}|/2<\mu\leq|M_{c}|$ in at least one contraction with $n_{c}$ vertices. On the other hand, Proposition 4 outputs subsets with a matching of size at least $\varepsilon n_{c}$. This means that the $\varepsilon$-approximation matching found on each subset determined by Proposition 4 is at least of size $\varepsilon^{2}n$.
Now, since we run Section 4.2 on all contractions, nevertheless they contain the maximum matching or not, assuming $n_{c}$ is the number of vertices in $c$, we have the total running time as:
$\sum_{c\in\mathcal{C}}\sqrt{n_{c}^{1+\varepsilon}\log n_{c}\cdot\mathsf{ORS}_{n_{c}}(\Theta(\varepsilon^{2}n_{c}))}\leq O(\operatorname{poly}\log n)\cdot\sqrt{n^{1+\varepsilon}\log n\cdot\mathsf{ORS}_{n}(\Theta(\varepsilon^{2}n))}.\qed$
Now, using Lemma 12 and Lemma 15 the proof of Theorem 1 is complete.
5 Bounding Density of (Ordered) Ruzsa-Szemerédi Graphs
In this section we prove upper bounds on the density of ORS graphs. Our main result is a proof of Theorem 2 which we present in Section 5.1. We then show that a much stronger upper bound can be proved if matchings cover more than half of vertices in Section 5.2.
5.1 Linear Matchings
The following lemma, which is one of our main tools in proving Theorem 2, shows that so long as vertex degrees are larger than a threshold, we can reduce the number of vertices rather significantly without hurting the size of a relatively large number of induced matchings.
Lemma 16.
Let $G$ be an $\mathsf{ORS}_{n}(r,t)$ graph with $r=cn$ with even $t$. Let $M_{1},\ldots,M_{t}$ be the $t$ matchings in $G$ as defined in Definition 2. Let us partition these matchings into two subsets
[…]
Suppose that for every vertex $v$, it holds that $\deg_{\mathcal{M}^{\prime}}(v)\geq\delta ct$ for some $\delta>0$. Then there exists an $\mathsf{ORS}_{n^{\prime}}(cn,t^{\prime})$ graph $H$ on $n^{\prime}<(1-\delta c^{2}/2)n$ vertices where $t^{\prime}\geq\delta c^{2}t/(8\cdot 2^{x})$ and $x=4n/ct$.
Let us start with an observation that is extremely helpful in proving Lemma 16.
Observation 17.
Take a matching $M^{\prime}\in\mathcal{M^{\prime}}$ and take any vertex $v$ matched by $M^{\prime}$. Let $N_{\mathcal{M}}(v)$ be the set of neighbors of $v$ in $\mathcal{M}$. That is, $u\in N_{\mathcal{M}}(v)$ iff there is $M\in\mathcal{M}$ such that $(v,u)\in M$. Then no vertex in $N_{\mathcal{M}}(v)$ can be matched in $M^{\prime}$.
Proof.
Suppose there is $u\in N_{\mathcal{M}}(v)$ that is matched by $M^{\prime}$. Let $(u,w)\in M^{\prime}$ be the matching edge involving $u$. Also take the edge $(v,y)\in M^{\prime}$ that involves $v$ (which exists by definition of $v$). Note that since the matchings $M_{1},\ldots,M_{t}$ are edge-disjoint by Definition 2, $v$ and $u$ cannot be matched together in $M^{\prime}$ (as $v$ and $u$ must be matched in some other matching in $\mathcal{M}$ by definition of $N_{\mathcal{M}}(v)$). Now since $(u,w),(v,y)\in M^{\prime}$ but $(v,u)$ belongs to some matching in $\mathcal{M}$ that comes before $M^{\prime}$ in the ordering, matching $M^{\prime}$ cannot be an induced matching among its previous matchings, contradicting Definition 2 for ORS graphs. ∎
We are now ready to prove Lemma 16.
Proof of Lemma 16.
We explain the proof in a few steps.
Step 1: The pivots.
We iteratively take a vertex of highest remaining degree in $\mathcal{M}$, and remove its neighbors in $\mathcal{M}$ from the graph. More formally, take the graph $G_{0}=(V_{0},E_{0})$ where $V_{0}=V(G)$ and $E_{0}$ is the union of the matchings in $\mathcal{M}$. At step $i$, we choose $v_{i}$ in $V_{i-1}$ with the maximum degree in $G_{i-1}$. Let us define $N_{i}$ as the neighbors of $v_{i}$ in $G_{i-1}$. We then remove $v_{i}$ and its neighbors in $G_{i-1}$ to define $G_{i}=(V_{i},E_{i})$. Namely, $V_{i}=V_{i-1}\setminus(\{v_{i}\}\cup N_{i})$ and $E_{i}$ includes all edges in $E_{i-1}$ except those that have at least one endpoint that does not belong to $V_{i}$.
Let $k$ be the maximum integer such that $|N_{k}|>n/x$. We define $P=\{v_{1},\ldots,v_{k}\}$ to be the set of pivots. Note that since $N_{1},\dots,N_{k}$ are disjoint sets, we get that $n\geq\sum_{i=1}^{k}|N_{i}|\geq kn/x$ which implies that
$|P|\leq x.$ (1)
Moreover, note that when removing $v_{i}$, we remove $|N_{i}|$ vertices from the graph and each of those vertices has remaining degree at most $|N_{i}|$ (as we choose $v_{i}$ to be the vertex of highest remaining degree). In other words, we remove at most $|N_{i}|^{2}$ from the graph after removing $v_{i}$ and its remaining neighbors. On the other hand, after removing $v_{1},\ldots,v_{k}$ and their neighbors, the maximum degree in the graph is at most $n/x$ (by definition of $k$), and so there are at most $n^{2}/x$ edges in the graph. This implies that:
$|E_{0}|-\sum_{i=1}^{k}|N_{i}|^{2}\leq n^{2}/x.$
Given that $E_{0}$ includes $t/2$ edge disjoint matchings of size $cn$, we get that $|E_{0}|\geq tcn/2$. Plugged into the inequality above, this implies that
$\sum_{i=1}^{k}|N_{i}|^{2}\geq|E_{0}|-n^{2}/x\geq tcn/2-n^{2}/x.$ (2)
Step 2: Vertex reduction.
Take a matching $M\in\mathcal{M}^{\prime}$. Let $S_{M}$ be the set of pivots matched by $M$, i.e., $S_{M}=P\cap V(M)$. Let us define $Y_{M}:=V\setminus\cup_{v_{i}\in S}N_{i}$. From Observation 17, we get that all edges of $M$ must have both endpoints still in $Y_{M}$. Suppose for now that $M\in\mathcal{M}^{\prime}$ is chosen uniformly at random. We first argue that the expected size of $|Y_{M}|$ is relatively small (despite it preserving all edges of $M$ completely). We have
$\operatorname{\textbf{E}}_{M\sim\mathcal{M}^{\prime}}[|Y_{M}|]$ $=n-\sum_{i=1}^{k}\Pr[v_{i}\in V(M)]\cdot|N_{i}|$
$=n-\sum_{i=1}^{k}\frac{\deg_{\mathcal{M}^{\prime}}(v_{i})}{|\mathcal{M}^{\prime}|}\cdot|N_{i}|$ (As $M$ is chosen uniformly from $\mathcal{M}^{\prime}$ and $v_{i}$ is matched in $\deg_{\mathcal{M}^{\prime}}(v_{i})$ of matchings in $\mathcal{M}^{\prime}$.)
$=n-\sum_{i=1}^{k}\frac{2\deg_{\mathcal{M}^{\prime}}(v_{i})}{t}\cdot|N_{i}|$ (Since $|\mathcal{M}^{\prime}|=t/2$.)
$\leq n-\sum_{i=1}^{k}\frac{2\delta ct}{t}\cdot|N_{i}|$ (Since $\deg_{\mathcal{M}^{\prime}}(v_{i})\geq\delta ct$ as assumed in Lemma 16.)
$\leq n-\sum_{i=1}^{k}\frac{4\delta c}{t}\cdot|N_{i}|^{2}.$ (Since $|N_{i}|\leq t/2$ as $\mathcal{M}$ is the union of $t/2$ matchings.)
$\leq n-\frac{4\delta c}{t}\cdot(tcn/2-n^{2}/x)$ (By 2.)
$=n-2\delta c^{2}n+\frac{4\delta cn^{2}}{tx}$
$\leq n-2\delta c^{2}n+\frac{4\delta cn^{2}}{(4n/cx)x}$ (Since $t=4n/cx$ by definition of $x$ in Lemma 16.)
$=n-2\delta c^{2}n+\delta c^{2}n$
$=(1-\delta c^{2})n.$ (3)
Instead of the expected value of $Y_{M}$, we need some (weak) concentration. Take $\varepsilon=\delta c^{2}/2$ and note from 3 that $(1+\varepsilon)\operatorname{\textbf{E}}[|Y_{M}|]<(1+\varepsilon)(1-\delta c^{2})n<(1-\delta c^{2}/2)n$. We have
$\Pr_{M\sim\mathcal{M}^{\prime}}\Big{[}|Y_{M}|<(1+\varepsilon)\operatorname{\textbf{E}}[|Y_{M}|]<(1-\delta c^{2}/2)n\Big{]}$ $=1-\Pr_{M\sim\mathcal{M}^{\prime}}\Big{[}|Y_{M}|>(1+\varepsilon)\operatorname{\textbf{E}}[|Y_{M}|]\Big{]}$
$\geq 1-\frac{1}{1+\varepsilon}$ (By Markov’s inequality.)
$\geq\varepsilon/2=\delta c^{2}/4.$ (4)
We call a matching $M\in\mathcal{M}^{\prime}$ a good matching if $|Y_{M}|<(1-\delta c^{2}/2)n$. From 4 we get that at least $(\delta c^{2}/4)|\mathcal{M}^{\prime}|=\delta c^{2}t/8$ matchings are good.
Step 3: Grouping good matchings.
Now, we partition the good matchings in $\mathcal{M}^{\prime}$ into $2^{|P|}\leq 2^{x}$ classes depending on which subset of the pivots they match. Since there are at least $\delta c^{2}t/8$ good matchings, at least $\frac{\delta c^{2}t}{8\cdot 2^{x}}$ of them must belong to the same class $C$. Let $Y=Y_{M}$ for some $M\in C$. Note also that $Y=Y_{M^{\prime}}$ for any other matching $M^{\prime}\in C$ as well since they all match the same subset of pivots. Moreover, we have $|Y|<(1-\delta c^{2}/2)n$ since $M$ is good, and all matchings in $C$ only match vertices in $Y$ as discussed earlier (as a consequence of Observation 17). This implies that the graph $H$ on vertex set $Y$ and edge-set $C$ (i.e., including all edges of all matchings in $C$) is an $\mathsf{ORS}_{n^{\prime}}(cn,t^{\prime})$ graph for $n^{\prime}<(1-\delta c^{2}/2)n$ and $t^{\prime}\geq\delta c^{2}t/(8\cdot 2^{x})$ as desired. ∎
The next lemma follows by applying Lemma 16 after carefully pruning vertices of low degree in $\mathcal{M}^{\prime}$. Intuitively, Lemma 18 significantly reduces the number of vertices and shows that a relatively large number of matchings will have a lot of edges in the remaining graph.
Lemma 18.
Let $G$ be an $\mathsf{ORS}_{n}(cn,t)$ graph for even $t$. Then for some $\kappa\geq c^{3}/160$, there exist $\mathsf{ORS}_{n^{\prime}}(c^{\prime}n^{\prime},t^{\prime})$ graphs such that $n^{\prime}\leq n$, $c^{\prime}\geq(1+\kappa)c$ and $t^{\prime}\geq t/2^{O(n/ct)}$.
Proof.
Let $\mathcal{M}$ and $\mathcal{M}^{\prime}$ be defined for $G$ as in Lemma 16. We would like to apply Lemma 16 iteratively to prove Lemma 18. However, Lemma 16 requires a lower bound on the degrees in $\mathcal{M}^{\prime}$ which might not necessarily hold for our graph $G$. Therefore to be able to use it, we first have to prune vertices of small degrees.
Let $G^{\prime}_{0}$ be the subgraph of $G$ only including the edges of the matchings in $\mathcal{M}^{\prime}$. Let $\delta=1/4$. We iteratively remove vertices of low degree. Formally, in step $i$, if there is a vertex $v_{i}$ with degree less than $2\delta ct$ in $G^{\prime}_{i-1}$, we define the graph $G^{\prime}_{i}$ of the next step to be graph obtained after removing $v_{i}$ and its edges from $G^{\prime}_{i-1}$. Let $G^{\prime}_{k}$ be the final graph that does not have any vertex of degree smaller than $2\delta ct$.
Suppose that $k=\alpha n$; that is, we remove $\alpha n$ vertices in the process above. We consider two cases depending on the value of $\alpha$.
Case 1: $\alpha\geq\delta c^{3}/10$.
Since we remove at most $\delta ct$ edges in every step, the total number of removed edges is at most $\delta\alpha ctn$ over the $k=\alpha n$ steps of constructing $G^{\prime}_{k}$. Say a matching $M\in\mathcal{M}^{\prime}$ is damaged if a total of at least $3\delta\alpha cn$ of its edges have been removed. In other words, $M$ is damaged if less than $|M|-3\delta\alpha cn=(1-3\delta\alpha)cn$ of its edges belong to $G^{\prime}_{k}$. Since the total number of edges removed is at most $\delta\alpha ctn$ and each damaged matching has at least $3\delta\alpha cn$ of its edges removed, the total number of damaged matchings can be upper bounded by $\delta\alpha ctn/3\delta\alpha cn=t/3$. Since $|\mathcal{M}^{\prime}|=t/2$, at least $t/6$ matchings in $\mathcal{M}^{\prime}$ are not damaged, and have size at least $(1-3\delta\alpha)cn$. These matchings themselves form an $\mathsf{ORS}_{n^{\prime}}(c^{\prime}n^{\prime},t/6)$ graph with parameters
$n^{\prime}=(1-\alpha)n\leq(1-\delta c^{3}/10)n\stackrel{{\scriptstyle(\delta=1/4)}}{{=}}(1-c^{3}/40)n,$
and
$c^{\prime}n^{\prime}\geq(1-3\delta\alpha)cn=(1-3\delta\alpha)c\frac{n^{\prime}}{1-\alpha}\stackrel{{\scriptstyle(\delta=1/4)}}{{=}}\left(1-\frac{3}{4}\alpha\right)c\frac{n^{\prime}}{1-\alpha}\geq\left(1+\frac{1}{4}\alpha\right)cn^{\prime}\geq(1+c^{3}/160)cn^{\prime},$
which implies that $c^{\prime}\geq(1+c^{3}/160)c$. Taking $\kappa=c^{3}/160$ proves the lemma in this case.
Case 2 – $\alpha<\delta c^{3}/10$:
Let $G_{k}$ be the graph $G$ after removing vertices $v_{1},\ldots,v_{k}$ (the difference between $G_{k}$ and $G^{\prime}_{k}$ is that $G^{\prime}_{k}$ only includes the edges of $\mathcal{M}^{\prime}$ but $G_{k}$ includes both edges of $\mathcal{M}^{\prime}$ and $\mathcal{M}$ that do not have any endpoint removed). Since we remove a total of $\alpha n$ vertices from $G$ to obtain $G_{k}$, each matching in $\mathcal{M}^{\prime}$ and $\mathcal{M}$ loses at most $\alpha n$ edges. Therefore, $G_{k}$ is an $\mathsf{ORS}_{n^{\prime\prime}}(c^{\prime\prime}n^{\prime\prime},t^{\prime\prime})$ graph with parameters
$n^{\prime\prime}=(1-\alpha)n,\qquad c^{\prime\prime}\geq\frac{cn-\alpha n}{n^{\prime\prime}}=\frac{(c-\alpha)n}{(1-\alpha)n}\geq c-\alpha\geq(1-\delta c^{2}/10)c,\qquad t^{\prime\prime}=t.$ (5)
Additionally, by the construction of $G^{\prime}_{k}$, all vertices $v$ in $G_{k}$ satisfy
$\deg_{\mathcal{M}^{\prime}}(v)\geq 2\delta ct\geq\delta c^{\prime\prime}t^{\prime\prime}$ (Since $t^{\prime\prime}=t$ and $c^{\prime\prime}\leq\frac{cn}{n^{\prime\prime}}=\frac{cn}{(1-\alpha)n}\leq 2c$ where the last inequality follows from $\alpha<1/2$.)
This now satisfies the requirements of Lemma 16. Applying it on graph $G_{k}$, we obtain an $\mathsf{ORS}_{n^{\prime}}(c^{\prime}n^{\prime},t^{\prime})$ graph where the number of vertices satisfies
$n^{\prime}$ $<(1-\delta c^{\prime\prime 2}/2)n^{\prime\prime},$ (6)
the number of matchings satisfies
$t^{\prime}\geq\delta c^{\prime\prime 2}t^{\prime\prime}/(8\cdot 2^{4n^{\prime\prime}/c^{\prime\prime}t^{\prime\prime}})=\delta c\cdot t/2^{O(n/ct)}=t/2^{O(n/ct)},$
and finally since Lemma 16 does not change the size of matchings, we get
$c^{\prime}n^{\prime}$ $=c^{\prime\prime}n^{\prime\prime}$
$\geq\frac{c^{\prime\prime}n^{\prime}}{(1-\delta c^{\prime\prime 2}/2)}$ (By 6.)
$\geq(1+\delta c^{\prime\prime 2}/2)c^{\prime\prime}n^{\prime}$
$\geq\Big{(}1+\delta\Big{(}(1-\delta c^{2}/10)c\Big{)}^{2}/2\Big{)}\Big{(}(1-\delta c^{2}/10)c\Big{)}n^{\prime},$ (Since the RHS is minimized when $c^{\prime\prime}$ is minimized, thus we can replace $c^{\prime\prime}$ with its LB from 5.)
$\geq\Big{(}1+0.4\delta c^{2}\Big{)}\Big{(}(1-\delta c^{2}/10)c\Big{)}n^{\prime}$ (Since $\delta\big{(}(1-\delta c^{2}/10)c\big{)}^{2}/2\geq\delta(0.9c)^{2}/2>0.4\delta c^{2}$)
$>(1+0.07c^{2})cn^{\prime}.$
Dividing both sides of the inequality by $n^{\prime}$ implies $c^{\prime}\geq(1+0.07c^{2})c$. Taking $\kappa=0.07c^{2}\geq c^{3}/160$ implies the lemma. ∎
We are now ready to prove Theorem 2.
Proof of Theorem 2.
We prove that for some $d=\operatorname{poly}(1/c)$, there does not exist any $\mathsf{ORS}_{n}(cn,t)$ graph $G$ with $t\geq n/\log_{b}^{(d)}n$ where $b=2^{\beta/c}$ for some large enough constant $\beta\geq 1$. Note that the base of the iterated log is different from the statement of the theorem, but since $\log^{(x)}_{b}z=\log^{(\Theta(x\log^{*}b))}z$, this implies the theorem as well by letting $\ell=\Theta(d\cdot\log^{*}b)=\operatorname{poly}(1/c)$.
Suppose for the sake of contradiction that that there exists an $\mathsf{ORS}_{n}(cn,t)$ graph $G$ with $t\geq n/\log_{b}^{(d)}n$. Let $\kappa$ be as in Lemma 18. We iteratively apply Lemma 18 for $k=\log_{2}(1/c)/\kappa=\operatorname{poly}(c)$ steps to obtain a sequence of graphs $G_{0},G_{1},G_{2},G_{3},\ldots,G_{k}$ where $G_{0}=G$ is the original graph, $G_{1}$ is obtained by applying Lemma 18 on $G_{0}$, $G_{2}$ is obtained by applying Lemma 18 on $G_{1}$, and so on so forth.
Let us now analyze the ORS properties of graph $G_{k}$, starting with the parameter $c_{k}$. Since every application of Lemma 18 multiplies the parameter $c_{i}$ by a factor of at least $(1+\kappa)$, we have
$c_{k}$ $\geq(1+\kappa)^{k}c$
$=(1+\kappa)^{\log_{2}(1/c)/k}c$
$\geq 2^{\log_{2}(1/c)}c$ (Since $(1+\kappa)^{1/\kappa}\geq 2$ for all $0<\kappa\leq 1$)
$\geq 1.$
This implies that in graph $G_{k}$, there must be $t_{k}$ edge-disjoint matchings of size $c_{k}n_{k}\geq n_{k}$, but each matching in a graph on $n_{k}$ vertices can have size at most $n_{k}/2$. In other words, we must have $t_{k}=0$. We will obtain a contradiction by proving that $t_{k}\geq 1$.
We prove by induction that for every $i\in[d]$,
$t_{i}\geq\frac{n}{\log^{(d-i)}_{b}n}.$
For the base case $i=0$ this holds as assumed at the beginning of the proof. By choosing appropriately large $\beta$, we know by Lemma 18 that,
$t_{i}$ $\geq t_{i-1}/2^{0.5\beta(n_{i-1}/c_{i-1}t_{i-1})}$
$\geq t_{i-1}/2^{0.5\beta(n/ct_{i-1})}$ (Since $n_{i-1}\leq n$ and $c_{i-1}\geq c$ as we iteratively apply Lemma 18.)
$=t_{i-1}/b^{0.5(n/t_{i-1})}$ (Since we defined $b=2^{\beta/c}$.)
$\geq\frac{n}{\log^{(d-i+1)}_{b}n}\cdot b^{-0.5\left(n/\frac{n}{\log^{(d-i+1)}_{b}n}\right)}$ (By the induction hypothesis $t_{i-1}\geq\frac{n}{\log^{(d-i+1)}_{b}n}$.)
$=\frac{n}{\log^{(d-i+1)}_{b}n}\cdot b^{-0.5\log^{(d-i+1)}_{b}n}$
$=\frac{n}{\log^{(d-i+1)}_{b}n\cdot\sqrt{\log^{(d-i)}_{b}n}}$
$\geq\frac{n}{\log^{(d-i)}_{b}n},$
concluding the proof.
Now let $d>k=\operatorname{poly}(1/c)$. From the above, we get that $t_{k}\geq n/(\log_{b}^{(d-k)})\geq 1$, which as discussed is a contradiction. ∎
5.2 When Matchings are (Very) Large
The goal of this section is to prove that the number of matching of an $\mathsf{ORS}$ graph with $r>n/4$ is bounded by a constant value. More precisely we prove Theorem 3. An equivalent of this theorem was already known for RS graphs (see [24]), and we show that almost the same proof carries over to ORS graphs as well.
Theorem 3.
For any $c>1/4+\varepsilon$, $\mathsf{ORS}_{n}(cn,t)\leq 1/\varepsilon+1$.
First, we argue that the number of common vertices of any two matching in ORS is at most $r$, using this we can prove that a constant value bounds the number of matching since the number of subsets of length $2r$ with at most $r$ common vertices between any two sets.
Lemma 19.
Suppose $G$ is an $\mathsf{ORS}(r,t)$ graph on $n$ vertices and let $M_{1},\ldots,M_{t}$ be the corresponding ordered list of matchings as in Definition 2. Then for any $i\not=j$ in $[t]$, we have $|V(M_{i})\cap V(M_{j})|\leq r$.
Proof.
Let us assume w.l.o.g. that $i<j$ and suppose that $|V(M_{i})\cap V(M_{j})|\geq r+1$. By the pigeonhole principle, this means that there must be an edge in $M_{i}$ whose both endpoints are matched in $M_{j}$, contradicting the fact from Definition 2 that $M_{j}$ is an induced matching in $M_{1}\cup\ldots\cup M_{j}$. ∎
Now, we use a similar approach as [24] to prove that $t<1/\varepsilon$. This theorem only uses the fact that the number of common vertices of two matchings is at most $r$.
Proof of Theorem 3.
Using Lemma 19 we have that the distance between sets of vertices of any two matchings is at most $r$, this implies a constant upper bound for the size of $\mathsf{ORS}$ using bounds from coding theory. To improve the upper bound we set $v_{i}\in\{0,1\}^{n}$, a vector that represents vertices in $V(M_{i})$. Let us use $a_{i}$ to denote the number of vectors $v_{j}$ with 00 in their $i$-th coordinate and let $b_{i}$ be the number of vectors with $1$ in their $i$-th coordinate. If we set $v_{0}$ a vector of all zeros. Note that we have $a_{i}+b_{i}=t+1$. Now, by double counting we have:
$2r{t+1\choose 2}\leq\sum_{0\leq i<j\leq t}dist(v_{i},v_{j})=\sum_{i=1}^{n}a_{i}b_{i}.$ (7)
By $dist(v_{i},v_{j})$ we mean the Hamming distance which satisfies the following inequality:
$dist(v_{i},v_{j})=|V_{i}|+|V_{j}|-2|V_{i}\cap V_{j}|\geq 2r.$
Note that in 7, $a_{i}b_{i}$ maximizes when we have $a_{i}=b_{i}=(t+1)/2$ for odd $t$ and $a_{i}=(t+2)/2,b_{i}=t/2$ for even $t$. This concludes that $t<1/\varepsilon+1$
Acknowledgements
The first author thanks Sepehr Assadi, Sanjeev Khanna, Huan Li, Ray Li, Mohammad Roghani, and Aviad Rubinstein for helpful discussions about RS graphs and their applications in dynamic graphs over the years.

### v2

@@
 While determining the density of ORS (or RS) remains a hard problem in combinatorics, we prove that if the existing constructions of ORS graphs are optimal, then our algorithm runs in $n^{1/2+O(\varepsilon)}$ time for any fixed $\varepsilon>0$ which would be significantly faster than existing near-linear in $n$ time algorithms.
-Our second main contribution is a better upper bound on density of both ORS and RS graphs with linear size matchings.
-The previous best upper bound was due to a proof of the triangle-removal lemma from more than a decade ago due to Fox [Annals of Mathematics ’11].
 1 Introduction
@@
 Unfortunately there is a huge gap between existing lower and upper bounds for $\mathsf{RS}_{n}(cn)$ (and as a result also for $\mathsf{ORS}_{n}(cn)$).
-To the best of our knowledge, the best known upper bound on $\mathsf{RS}_{n}(cn)$ follows from the improved triangle-removal lemma due to Fox [23] from more than a decade ago which shows $\mathsf{RS}_{n}(cn)\leq n/2^{\Omega(\log^{*}n)}=n^{1-o(1)}$ (see also [24, 2]).
-Our second main result in this paper is the following upper bound.
+The best known upper bound on $\mathsf{RS}_{n}(cn)$ for linear size matchings follows from the improved triangle-removal lemma of Fox [23] which implies $\mathsf{RS}_{n}(cn)\leq n/\log^{(\ell)}$ for $\ell=O(\log(1/c))$ where $\log^{(x)}$ is the iterated log function.
+We note that this result is implicit in [23] and was mentioned in the paper of [24].
+To our knowledge, this upper bound does not carry over to ORS graphs (we briefly discuss this at the beginning of Section 5).
+Our second result is a similar upper bound for ORS albeit with a worse dependence on constant $c$.
 Theorem 2.
-For any $c>0$, it holds that $\mathsf{ORS}_{n}(cn)=O(n/\log^{(\ell)}n)$ for some $\ell=\operatorname{poly}(1/c)$, where $\log^{(x)}$ is the iterated log function.
-Since every RS graph is also an ORS graph with the same parameters, Theorem 2 immediately implies the following bound on the density of RS graphs:
-Corollary 3.
-For any $c>0$, it holds that $\mathsf{RS}_{n}(cn)=O(n/\log^{(\ell)}n)$ for some $\ell=\operatorname{poly}(1/c)$.
-Note that $\log^{(\ell)}n\gg 2^{\Omega(\log^{*}n)}$ for any fixed $c$. As such, Corollary 3 also improves the upper bound for RS graphs after more than a decade since the work of [23].
+For any $c>0$, it holds that $\mathsf{ORS}_{n}(cn)=O(n/\log^{(\ell)}n)$ for some $\ell=\operatorname{poly}(1/c)$.
+Since every RS graph is also an ORS graph with the same parameters, Theorem 2 immediately implies the same upper bound for $\mathsf{RS}_{n}(cn)$.
+Note that this implication is not a new result, but the proof is very different from the abovementioned upper bound.
 1.2 Perspective: ORS vs RS
-Summarizing the above-mentioned bounds, the state-of-the-art bounds for the values of $\mathsf{RS}_{n}(cn)$ and $\mathsf{ORS}_{n}(cn)$ can be summarized as
+Summarizing the above-mentioned bounds, we have
 […]
@@
 Proof.
-Algorithm 1: RandomSamplingAlgorithm$(G[U],\varepsilon,\mu)$   Parameters: $\varepsilon\in(0,1]$.
-$M\leftarrow\emptyset$.
-$M_{1},...,M_{1/(2\varepsilon)}\leftarrow\emptyset$.
-$S\leftarrow\emptyset$ $b\leftarrow 1$.
-$c=0$.
-while $|M|<\varepsilon\mu$ do 1       if $b>\varepsilon c\cdot\log n$ then 2             $M_{c+1}\leftarrow M$.
-$M\leftarrow\emptyset$.
-$c\leftarrow c+1$.
-3      for vertex $v\in U\setminus V(M)$ in random order do 4             $S\leftarrow$ Sample $b$ vertices in $U$.
-if $\exists u\in S$ such that $u\in U$ and $u\notin V(M)$ and $u$ is a neighbor of $v$ then 5                   $M\leftarrow(v,u)\cup M$.
-6             7      $b\leftarrow 2b$ return $M$
-Let us consider Section 4.1, First we prove that this algorithm terminates.
-We need to prove that there exists $c$ such that sampling vertices with a budget $b$ in the interval of $[\varepsilon c\cdot\log n,\varepsilon(c+1)\cdot\log n]$ outputs a matching of size at least $\varepsilon\mu$.
+[…]
+Let us consider Section 4.1, first we prove that this algorithm will provide a matching $M_{c}$ of size a least $\varepsilon\mu/2$ by sampling vertices with budget $b\in[2^{\varepsilon c\cdot\log n},2^{\varepsilon(c+1)\cdot\log n}]$.
 Let us assume there are no such intervals.
-This means that for any integer $c\in[0,1/(2\varepsilon)]$, running the algorithm for $b$ in the interval $[\varepsilon c\cdot\log n,\varepsilon(c+1)\cdot\log n]$ outputs a matching with less than $\varepsilon\mu$ edges.
-Note that for each edge $e$ in $M_{1}\cup\dots\cup M_{1/(2\varepsilon)}$, at most 2 edges in the certificate matching $\mu$ have common vertices with $e$.
+This means that for any integer $c\in[0,1/(2\varepsilon)]$, running the algorithm for $b$ in the interval $[2^{\varepsilon c\cdot\log n},2^{\varepsilon(c+1)\cdot\log n}]$ outputs a matching with less than $\varepsilon\mu$ edges.
+Note that for each edge $e$ in $M_{1}\cup\dots\cup M_{1/(2\varepsilon)}$, at most 2 edges in the maximum matching of $G[U]$ have common vertices with $e$.
 Summing over all the $1/(2\varepsilon)$ intervals we have:
-$2\sum_{c=1}^{1/(2\varepsilon)}{|M_{i}|}<2\cdot\frac{1}{2\varepsilon}\cdot\varepsilon\mu=\mu.$
-The fact that we could not find a matching of size $\mu$ is a contradiction with the lemma statement ensuring the existence of such matching.Therefore, we have at least the size of one of the matchings $M_{1},...,M_{1/(2\varepsilon)}$ is $\varepsilon\mu$.
+$22\sum_{c=1}^{1/(2\varepsilon)}{|M_{i}|}<2\cdot\frac{1}{2\varepsilon}\cdot\varepsilon\mu/2=\mu/$ 2 .
+The fact that we could not find a matching of size $\mu/2$ is a contradiction with the lemma statement ensuring that the size of maximum matching is at least $\mu/2$.
+Note that by exhausting the algorithm with budget of $n$, we will go through all the edges and output a maximal matching.
+Note that any maximal matching in this graph should be of size at least $\mu/2$.Therefore, we have at least the size of one of the matchings in$\{M_{1},...,M_{1/(2\varepsilon)}\}$ is $\varepsilon\mu/2$.
 Now, we prove the statement of the algorithm.
-We assume that this algorithm terminates at $\varepsilon(c-1)\cdot\log n\leq b<\varepsilon c\cdot\log n$, for a constant $\delta$, we spend $n[b+b/2+..+1]<\delta nb\leq\delta\varepsilon c\cdot n\log n<\delta\varepsilon cn^{1+\varepsilon}$ time to find the matching $M_{c}$ such that $|M_{c}|\geq\varepsilon\mu$.
-Note that if it takes at most $n^{2+\varepsilon}/d$ time to find the matching of size $\varepsilon n$, then we have $n^{2+\varepsilon}/d<\delta\varepsilon cn^{1+\varepsilon}$ therefore, $d>\frac{n}{\delta\varepsilon c}$.
-For budget $b$, we sample vertices for all the vertices, we argue that for the vertices that are not in $M$, and after sampling $b$ vertices we could not find their neighbor then the degree of vertices out of the matching is less than $n/b$.
-Assuming that the degree of vertex $v$ is $d_{v}$, with probability $(1-\frac{d_{v}}{n})^{b}$ we sampled a neighbor, and this probability converges to $1/e$ for $d_{v}>O(n/b)$.
+We assume that this algorithm terminates at $b$ such that $2^{\varepsilon(c-1)\cdot\log n}\leq b<2^{\varepsilon c\cdot\log n}$, for a constant $\delta$, we spend $n[b+b/2+..+1]<\delta nb\leq\delta n\cdot 2^{\varepsilon c\log n}<\delta n^{1+\varepsilon c}$ time to find the matching $M_{c}$ such that $|M_{c}|\geq\varepsilon\mu/2$.
+Note that if it takes at most $n^{2+\varepsilon}/d$ time to find the matching of size $\varepsilon\mu/2$, then we have $n^{2+\varepsilon}/d<\delta n^{1+\varepsilon c}$ therefore, $d>\frac{n^{1+\varepsilon-\varepsilon c}}{\delta}\geq\frac{n^{1+\varepsilon}}{\delta}$.
+For budget $b$, we sample vertices for all the vertices, we argue that for the vertices that are not in $M$, and after sampling $b$ vertices we could not find their neighbor then the degree of vertices out of the matching is less than $O(n/b)$.
+Assuming that the degree of vertex $v$ is $d_{v}=c\cdot\frac{n}{b}$, with probability $(1-\frac{d_{v}}{n})^{b}=(1-c\cdot\frac{1}{b})^{b}\approx(1/e)^{c}$ we sampled a neighbor.
 Therefore in this case with a constant probability we already matched $v$.
-This implies that if we find the first edge of this matching at a budget of $2b$, we get that the degree of any vertex out of the matching we found until this point is less than $O(n/b)$.
-Therefore, the total sum of the degree of the matching is less than $O(\varepsilon\mu\cdot n/b)<O(\varepsilon\mu n/((c-1)\log n))<O(\delta\varepsilon^{2}cdn/((c-1)\log n))<O(nd)$.
+This implies that if we find the first edge of this matching at a budget of $2b$, we get that the degree of any vertex out of the matching we found until this point is less than $c\cdot\frac{n}{b}$ for any constant $c$.
+Picking $c$ small enough, the total sum of the degree of the matching is at most $\varepsilon\mu c\cdot\frac{n}{b}<\varepsilon\mu n^{(1-\varepsilon(c-1))}=O(nd)$.
 ∎
@@
 This means we can output a matching after each $\varepsilon n^{\prime}$ update and get an $\varepsilon$-approximate matching.
-Observation 10.
+Observation 9.
 Given a graph $G=(V,E)$ with $n$ vertices, if the maximum matching size is $O(\mu)$ and we provide an $\varepsilon$-approximate matching $M^{\prime}$, after $\varepsilon^{2}\mu$ updates in the dynamic model, $M^{\prime}$ or one of its subsets remains an $\varepsilon$-approximate matching for the updated graph.
@@
 This algorithm works in phases, each phase represents the number of updates we want to use the matchings we found in the last updates.
-Algorithm 2: Algorithm for Dynamic Model in Contracted Graphs   Parameters: $\varepsilon\in(0,1]$, $t$ the threshold of the algorithm.
-$G^{\prime}\leftarrow$ the contracted graph.
-$S_{1}\leftarrow\emptyset$.
-$S_{2}\leftarrow\emptyset$.
-while  edge updates $<t$  do 1       if There are $\varepsilon^{2}n^{\prime}$new updates then 2             $S_{1}\leftarrow S_{1}\cup$ Insertions.
-$M\leftarrow\emptyset$.
-while Proposition 4 outputs a subset $U$ that has a matching of size at least $\mu=\varepsilon n^{\prime}$  do 3                   $M^{\prime}\leftarrow$ a maximal matching found by a greedy algorithm on edges of $S_{1}\cup S_{2}$ in $U$.
-For each edge in this set, we check if it exists in the dynamic graph.
-if  $|M^{\prime}|>\varepsilon\mu$  then 4                        $M\leftarrow M^{\prime}$.
-5                  else 6                         $M_{U}\leftarrow$ $RandomSamplingAlgorithm(G[U\setminus V(M^{\prime})],\frac{\varepsilon}{1-2\varepsilon},(1-2\varepsilon)\mu).$ $M\leftarrow M\cup M_{U}$.
-7                  $S_{2}\leftarrow S_{2}\cup M_{U}.$ Output $M$.
-8             9
+[…]
 First, we prove that this algorithm obtains a $\varepsilon$-approximation matching.
@@
 Proof.
-Note that we proved in Observation 10 that if the maximal matching size is of order $n^{\prime}$, then we do not need to output a matching in $\varepsilon^{2}n^{\prime}$ updates.
+Note that we proved in Observation 9 that if the maximal matching size is of order $n^{\prime}$, then we do not need to output a matching in $\varepsilon^{2}n^{\prime}$ updates.
 By Section 4.2, set $S_{1}$ contains edge insertions, and set $S_{2}$ contains all the matchings we found earlier in the algorithm.
@@
 We will run $O(\log^{2}n)$ parallel versions of Section 4.2 each for a maintained contracted graphs that we maintain. More formally we introduce Section 4.2:
-Algorithm 3: Main   Parameters: $\varepsilon\in(0,1]$.
-$k\leftarrow\log n$.
-$n_{1},n_{2},\dots,n_{k}\leftarrow 2,4,\dots,2^{k}$.
-$t\leftarrow$ threshold of algorithm Section 4.2.
-$\mathcal{(}C)\leftarrow$ set of maintained $O(\log n)$ contracted graphs for each $n_{c}\in\{n_{1},\dots,n_{k}\}$ using Proposition 7.
-$\mathcal{P}\leftarrow$ set of maintained parallel edges between two contracted sets of vertices after each update.
-Run a new phase Section 4.2 for each contraction after $t$ updates.
-For each contraction $c$ with $c_{n}$ vertices, run Proposition 5 on $c$ after $\varepsilon^{2}n_{c}$ updates to see whether $c$ preserves the maximum matching in $G$.
-Find the contraction $c$ preserving the largest matching of $G$ and use the output of Section 4.2 for the corresponding contraction, to output an $(1-\varepsilon)$-approximate matching.
+[…]
 Now, we prove that this algorithm outputs a $(1-\varepsilon)$-approximation matching.
@@
 In this section we prove upper bounds on the density of ORS graphs.
-Our main result is a proof of Theorem 2 which we present in Section 5.1.
+Our main result of this section is a proof of Theorem 2 which we present in Section 5.1.
 We then show that a much stronger upper bound can be proved if matchings cover more than half of vertices in Section 5.2.
+Before presenting our proof, let us first briefly discuss how the triangle removal lemma is useful for upper bounding density of RS graphs and why it does not seem to help for upper bounding ORS.
+The triangle removal lemma states that so long as the number of triangles in a graph are $o(n^{3})$, then we can make the graph triangle free by removing $o(n^{2})$ of its edges.
+Suppose we have a bipartite RS graph with induced matchings $M_{1},\ldots,M_{k}$.
+Add $k$ vertices $v_{1},\ldots,v_{k}$ to this graph and connect each $v_{i}$ to all the vertices matched by $M_{i}$.
+Now, crucially, because each $M_{i}$ is an induced matching, each $v_{i}$ is part of exactly $|M_{i}|$ triangles, one for each edge of $M_{i}$.
+As such, the total number of triangles can be upper bounded by $O(kn)=O(n^{2})$ which is well within the regime that one can apply triangle removal lemma.
+However, because the matchings in ORS are not induced matchings of the whole graph, the number of triangles cannot simply be upper bounded by $O(n^{2})$ after adding the auxiliary vertices.
+In fact, our upper bound completely deviates from this approach and does not rely on the triangle removal lemma.
 5.1 Linear Matchings
 The following lemma, which is one of our main tools in proving Theorem 2, shows that so long as vertex degrees are larger than a threshold, we can reduce the number of vertices rather significantly without hurting the size of a relatively large number of induced matchings.
+Our proof of Lemma 15 builds on the techniques developed for upper bounding RS graphs with matchings of size very close to $n/4$ in [24] that we extend to ORS graphs with balanced degrees.
 Lemma 15.
@@
 Let us partition these matchings into two subsets
-[…]
+=$\mathcal{M}=\{M_{1},\ldots,M_{t/2}\}\qquad\text{and}\qquad\mathcal{M}^{\prime}=\{M_{t/2+1},\ldots,M_{t}\}.$
 Suppose that for every vertex $v$, it holds that $\deg_{\mathcal{M}^{\prime}}(v)\geq\delta ct$ for some $\delta>0$.
@@
 Let us start with an observation that is extremely helpful in proving Lemma 15.
-Observation 17.
+Observation 16.
 Take a matching $M^{\prime}\in\mathcal{M^{\prime}}$ and take any vertex $v$ matched by $M^{\prime}$.
@@
 Let us define $Y_{M}:=V\setminus\cup_{v_{i}\in S}N_{i}$.
-From Observation 17, we get that all edges of $M$ must have both endpoints still in $Y_{M}$.
+From Observation 16, we get that all edges of $M$ must have both endpoints still in $Y_{M}$.
 Suppose for now that $M\in\mathcal{M}^{\prime}$ is chosen uniformly at random.
@@
 Note also that $Y=Y_{M^{\prime}}$ for any other matching $M^{\prime}\in C$ as well since they all match the same subset of pivots.
-Moreover, we have $|Y|<(1-\delta c^{2}/2)n$ since $M$ is good, and all matchings in $C$ only match vertices in $Y$ as discussed earlier (as a consequence of Observation 17).
+Moreover, we have $|Y|<(1-\delta c^{2}/2)n$ since $M$ is good, and all matchings in $C$ only match vertices in $Y$ as discussed earlier (as a consequence of Observation 16).
 This implies that the graph $H$ on vertex set $Y$ and edge-set $C$ (i.e., including all edges of all matchings in $C$) is an $\mathsf{ORS}_{n^{\prime}}(cn,t^{\prime})$ graph for $n^{\prime}<(1-\delta c^{2}/2)n$ and $t^{\prime}\geq\delta c^{2}t/(8\cdot 2^{x})$ as desired.

### v3

@@
 Until recently, the fastest known algorithm for this problem required $\Theta(n)$ time per update where $n$ is the number of vertices.
-This bound was slightly improved to $n/(\log^{*}n)^{\Omega(1)}$ by Assadi, Behnezhad, Khanna, and Li [STOC’23] and very recently to $n/2^{\Omega(\sqrt{\log n})}$ by Liu [ArXiv’24].
+This bound was slightly improved to $n/(\log^{*}n)^{\Omega(1)}$ by Assadi, Behnezhad, Khanna, and Li [STOC’23] and very recently to $n/2^{\Omega(\sqrt{\log n})}$ by Liu [FOCS’24].
 Whether this can be improved to $n^{1-\Omega(1)}$ remains a major open problem.
-In this paper, we present a new algorithm that maintains a $(1-\varepsilon)$-approximate maximum matching.
-The update-time of our algorithm is parametrized based on the density of a certain class of graphs that we call Ordered Ruzsa-Szemerédi (ORS) graphs, a generalization of the well-known Ruzsa-Szemerédi graphs.
-While determining the density of ORS (or RS) remains a hard problem in combinatorics, we prove that if the existing constructions of ORS graphs are optimal, then our algorithm runs in $n^{1/2+O(\varepsilon)}$ time for any fixed $\varepsilon>0$ which would be significantly faster than existing near-linear in $n$ time algorithms.
+In this paper, we introduce Ordered Ruzsa-Szemerédi (ORS) graphs (a generalization of Ruzsa-Szemerédi graphs) and show that the complexity of dynamic matching is closely tied to them.
+For $\delta>0$, define $\mathsf{ORS}(\delta n)$ to be the maximum number of matchings $M_{1},\ldots,M_{t}$, each of size $\delta n$, that one can pack in an $n$-vertex graph such that each matching $M_{i}$ is an induced matching in subgraph $M_{1}\cup\ldots\cup M_{i}$.
+We show that there is a randomized algorithm that maintains a $(1-\varepsilon)$-approximate maximum matching of a fully dynamic graph in
+$\widetilde{O}\left(\sqrt{n^{1+\varepsilon}\cdot\mathsf{ORS}(\Theta_{\varepsilon}(n))}\right)$
+amortized update-time.
+While the value of $\mathsf{ORS}(\Theta(n))$ remains unknown and is only upper bounded by $n^{1-o(1)}$, the densest construction known from more than two decades ago only achieves $\mathsf{ORS}(\Theta(n))\geq n^{1/\Theta(\log\log n)}=n^{o(1)}$ [Fischer et al.
+STOC’02].
+If this is close to the right bound, then our algorithm achieves an update-time of $\sqrt{n^{1+O(\varepsilon)}}$, resolving the aforementioned longstanding open problem in dynamic algorithms in a strong sense.
 1 Introduction
@@
 Background on Dynamic Matching:
-The dynamic matching problem has received a lot of attention over the last two decades [33, 4, 5, 32, 27, 18, 11, 12, 17, 16, 35, 21, 1, 13, 9, 10, 37, 14, 15, 34, 29, 26, 8, 20, 7, 19].
+The dynamic matching problem has received a lot of attention over the last two decades [39, 8, 9, 38, 32, 23, 16, 17, 22, 21, 41, 26, 2, 18, 13, 15, 43, 19, 20, 40, 34, 31, 12, 25, 11, 7, 24].
 There is a relatively simple algorithm that maintains a $(1-\varepsilon)$-approximate maximum matching, for any fixed $\varepsilon>0$, in just $O(n)$ time per update (see e.g.
@@
 Open Problem 1.
-Is it possible to maintain a $(1-\varepsilon)$-approximte maximum matching in a fully dynamic graph, for any fixed $\varepsilon>0$, in $n^{1-\Omega(1)}$ update-time?
-We note that there is an orthogonal line of work on fully dynamic algorithms that instead of maintaining the edges of the matching, maintain only its size [7, 20, 19].
+Is it possible to maintain a $(1-\varepsilon)$-approximate maximum matching in a fully dynamic graph, for any fixed $\varepsilon>0$, in $n^{1-\Omega(1)}$ update-time?
+We note that there is an orthogonal line of work on fully dynamic algorithms that instead of maintaining the edges of the matching, maintains only its size [11, 25, 24].
 For this easier version of the problem, Bhattacharya, Kiss, and Saranurak [24] positively resolved the open problem above.
@@
 Let $\varepsilon>0$ be fixed.
-There is a fully dynamic algorithm that maintains the edges of a $(1-\varepsilon)$-approximate maximum matching in $O\left(\sqrt{n^{1+\varepsilon}\cdot\mathsf{ORS}_{n}(\Theta(\varepsilon^{2}n))}\operatorname{poly}(\log n)\right)$ amortized update-time.
+There is a fully dynamic algorithm that maintains the edges of a $(1-\varepsilon)$-approximate maximum matching in $O\left(\sqrt{n^{1+\varepsilon}\cdot\mathsf{ORS}_{n}(\Theta_{\varepsilon}(n))}\operatorname{poly}(\log n)\right)$ amortized update-time.
 The algorithm is randomized but works against adaptive adversaries.
@@
 If this turns out to be the right bound, Theorem 1 implies a $(1-\varepsilon)$-approximation in $\widetilde{O}(n^{1/2+\varepsilon})$ time, an almost quadratic improvement over prior near-linear in $n$ algorithms of [6, 36].
-In fact, we note that so long as $\mathsf{ORS}_{n}(cn)$ is moderately smaller than $n$ (say $\mathsf{ORS}_{n}(\Theta(\varepsilon^{2}n))\ll n^{1-2\varepsilon}$) Theorem 1 implies a truly sublinear in $n$ update-time algorithm, positively resolving 1.
+In fact, we note that so long as $\mathsf{ORS}_{n}(cn)$ is moderately smaller than $n$ (say $\mathsf{ORS}_{n}(\Theta_{\varepsilon}(n))\ll n^{1-\Omega(1)}$) Theorem 1 still implies a truly sublinear in $n$ update-time algorithm, positively resolving 1.
 Finally, we note that there is a long body of work on proving conditional lower bounds for dynamic problems.
@@
 Adapting these lower bounds to the $(1-\varepsilon)$-approximate maximum matching problem has remained open since then.
-Our Theorem 1 implies that proving such lower bounds either requires a strong lower bound of near-linear on $\mathsf{ORS}_{n}(\varepsilon n)$, or requires a conjecture that implies this.
-Contribution 2: Better Upper Bounds for ORS and RS:
+Our Theorem 1 implies that proving such lower bounds either requires a strong lower bound of near-linear on $\mathsf{ORS}_{n}(\Theta(n))$, or requires a conjecture that implies this.
+Contribution 2: Upper Bounding ORS.
 Unfortunately there is a huge gap between existing lower and upper bounds for $\mathsf{RS}_{n}(cn)$ (and as a result also for $\mathsf{ORS}_{n}(cn)$).
-The best known upper bound on $\mathsf{RS}_{n}(cn)$ for linear size matchings follows from the improved triangle-removal lemma of Fox [23] which implies $\mathsf{RS}_{n}(cn)\leq n/\log^{(\ell)}$ for $\ell=O(\log(1/c))$ where $\log^{(x)}$ is the iterated log function.
+The best known upper bound on $\mathsf{RS}_{n}(cn)$ for linear size matchings follows from the improved triangle-removal lemma of Fox [28] which implies $\mathsf{RS}_{n}(cn)\leq n/\log^{(\ell)}n$ for $\ell=O(\log(1/c))$ where $\log^{(x)}$ is the iterated log function.
 We note that this result is implicit in [28] and was mentioned in the paper of [29].
@@
 Since every RS graph is also an ORS graph with the same parameters, Theorem 2 immediately implies the same upper bound for $\mathsf{RS}_{n}(cn)$.
-Note that this implication is not a new result, but the proof is very different from the abovementioned upper bound.
+Note that this implication is not a new result, but the proof is different from that of [28] and is closer to the arguments in [29].
 1.2 Perspective: ORS vs RS
@@
 Open Problem 2.
-Does it hold that $\mathsf{ORS}_{n}(cn)=\operatorname{poly}(\mathsf{RS}_{n}(O(cn)))$?
+Is it true that for fixed $\varepsilon>0$,
+$\mathsf{ORS}_{n}(\varepsilon n)\leq\operatorname{poly}(\mathsf{RS}_{n}(\Theta_{\varepsilon}(n)))?$
 In the event that the answer to 2 is positive and $\mathsf{RS}_{n}(cn)=n^{o(1)}$ for any fixed $c>0$, we also get that $\mathsf{ORS}_{n}(cn)=n^{o(1)}$.
 Therefore Theorem 1 would imply an $n^{1/2+O(\varepsilon)}$ time algorithm in this case.
-In the event that the answer to 2 is negative, one might wonder whether we can improve Theorem 1 by parametrizing it based on $\mathsf{RS}$ instead of $\mathsf{ORS}$.
+In the event that the answer to 2 is negative, one might wonder whether we can improve Theorem 1 by parametrizing it based on RS instead of ORS.
 Put differently, suppose that $\mathsf{ORS}_{n}(cn)=n^{1-o(1)}$ and $\mathsf{RS}_{n}(cn)=n^{o(1)}$.
 Can we somehow utilize the sparsity of RS graphs (instead of ORS graphs) in this case to improve existing dynamic matching algorithms?
-We start Section 2 by providing an input construction which informally shows ORS is the right parameter for Theorem 1 even if RS is much sparser.
+We start Section 2 by providing an input construction which informally shows ORS is the right parameter for Theorem 1 even if RS graphs turn out to be much sparser.
+1.3 Connections and Recent Developments
+Connection to Sublinear Time Algorithms:
+Our algorithm crucially relies on sublinear time algorithms; particularly the algorithm of the first author in [10] for estimating the size of maximum matching.
+While sublinear-time algorithms have also played a crucial role in the dynamic matching algorithms of [11, 25, 24], these algorithms maintain only the size of the matching as opposed to its edges.
+Our algorithm, on the other hand, maintains the matching explicitly yet relies on matching size estimators crucially.
+On a high level, our algorithm can tolerate the time needed to find a large matching in a specific induced subgraph so long as a large matching is guaranteed to exist there.
+We use the fast sublinear-time matching algorithm of [10] to first verify existence of a large matching, then spend the time needed if this matching is large.
+Connection to a Lower Bound of Liu [36]:
+In his very recent work, Liu [36] showed that under the approximate OMv conjecture, there is no algorithm that maintains a $(1-\varepsilon)$-approximate maximum matching in $n^{1-\Omega(1)}\operatorname{poly}(1/\varepsilon)$ update-time.
+This might, at the first glance, appear to contradict Theorem 1 which runs in $n^{1-\Omega(1)}$ time if it so happens that $\mathsf{ORS}(\Theta(n))=n^{1-\Omega(1)}$.
+However, we note that there is no contradiction even if approximate OMv conjecture turns out to be true and $\mathsf{ORS}(\Theta(n))=n^{1-\Omega(1)}$ at the same time!
+The reason is that when parameter $\varepsilon$ is sufficiently sub-constant, dense constructions of RS (and consequently ORS) with $\Theta(n^{2})$ edges already exist [1].
+So our Theorem 1 is only potentially useful for fixed (or mildly sub-constant) $\varepsilon$.
+On the other hand, the reduction of [36] sets $\varepsilon=n^{-\Omega(1)}$, thus only targets the sub-constant regime.
+Interestingly, while the result of [36] removes the combinatorial structure of the dynamic matching problem and reduces it to a purely algebraic question for the sub-constant regime of $\varepsilon$, our work shows that the fixed regime of $\varepsilon$ is, in fact, a completely combinatorial question.
+Recent Developments:
+After the first version of this paper appeared online, the nice follow-up works of Assadi and Khanna [4] and Kiss [35] improved our update-time from $O(\sqrt{n^{1+o(1)}\mathsf{ORS}(\Theta(n)})$ to $O(n^{o(1)}\mathsf{ORS}(\Theta(n)))$.
+Importantly, their algorithms run in $n^{o(1)}$ update-time under $\mathsf{ORS}(\Theta(n))=n^{o(1)}$, showing that ORS graphs fully characterize the hardness of the dynamic matching problem.
 2 Our Techniques
+In this section, we provide an informal overview of our algorithm for Theorem 1 as well as the upper bound of Theorem 2.
 Before describing the intuition behind our algorithm of Theorem 1, let us start with a sequence of updates that, in a sense, explains why existence of dense ORS graphs would make it challenging to maintain a $(1-\varepsilon)$-approximate maximum matching in a fully dynamic setting.
-We then proceed to show that this input construction is essentially the only instance that we do not know how to solve fast.
-An input construction based on ORS graphs:
+Why ORS graphs are seemingly hard:
 Consider a fully dynamic input graph $G=(V,E)$ that is composed of two types of vertices: the ORS vertices $V_{ORS}\subseteq V$ which is a subset of $n$ vertices, and the singleton vertices $V_{S}$ which is a subset of $(1-2\varepsilon)n$ vertices.
@@
 Hence, the amortized update-time of this algorithm must be at least $\Omega(nt/n)=\Omega(\mathsf{ORS}_{n}(\varepsilon n))$.
-The input construction above implies that to maintain a $(1-\varepsilon)$-approximation of maximum matching, either we have to find a way to identify induced matchings of an ORS graph fast (without scanning the neighbors of each vertex) or have to parameterize our algorithm’s update-time by $t$, the density of ORS graphs.
+The input construction above implies that to maintain a $(1-\varepsilon)$-approximation of maximum matching, either we have to find a way to identify induced matchings of an ORS graph fast (without scanning the neighbors of each vertex) which appears extremely challenging, or we have to parameterize our algorithm’s update-time by $t$, the density of ORS graphs.
+We take the latter approach in this work.
 Overview of our algorithm for Theorem 1:
 Let us for this informal overview of our algorithm assume that the maximum matching size is at least $\Omega(n)$.
-Having this assumption allows us to find the matching once, do nothing for the next $\varepsilon n$ updates, and then repeat without hurting the size of the approximate matching that we find by more than a $1+O(\varepsilon)$ factor.
+Having this assumption (which comes w.l.o.g.
+as stated by Proposition 5) allows us to find the matching once, do nothing for the next $\varepsilon n$ updates, and then repeat without hurting the size of the approximate matching that we find by more than a $1+O(\varepsilon)$ factor.
 As it is standard by now, to find a $(1-\varepsilon)$-approximate matching it suffices to design an algorithm that given a subset $U\subseteq V$, finds a constant approximate maximum matching in $G[U]$.
@@
 when it is an induced matching.
-Then instead of charging a matching in $G[U]$ that is an induced matching, we charge this matching of average degree at most $d$ inside.
-Let $M_{1},\ldots,M_{t}$ be the matchings that we charge and let $d_{i}$ be the average degree of the $i$-th matching and suppose that these matchings are edge-disjoint.
-We show in our update-time analysis that $\sum_{i=1}^{t}1/d_{i}$ can be at most $\mathsf{ORS}(\varepsilon^{2}n)$, and therefore the total time spent by our algorithm can be upper bounded by $n^{2}\mathsf{ORS}(\varepsilon^{2}n)$.
-For the second problem, or in other words, to ensure that the matchings that we charge are edge-disjoint, we maintain a set $S$ and add all edges of any matching that we charge to this set.
-Thereafter, before solving $G[U]$, we first go over the edges stored in this set and see whether they can be used to find a large matching in $G[U]$.
+Then instead of charging a matching in $G[U]$ that is an induced matching, we charge this matching of average degree at most $d$ inside, which we call a certificate.
+Let $M_{1},\ldots,M_{t}$ be the certificates that we charge and let $d_{i}$ be the average degree of the $i$-th matching and suppose that these matchings are edge-disjoint.
+We show in our update-time analysis that $\sum_{i=1}^{t}1/d_{i}$ can be at most $\mathsf{ORS}(\Theta(n))$, and therefore the total time spent by our algorithm during a phase can be upper bounded by $n^{2}\mathsf{ORS}(\Theta(n))$.
+For the second problem, or in other words, to ensure that the certificate matchings that we charge are edge-disjoint, we maintain a set $H_{\text{cert}}$ and add all edges of any matching that we charge to this set.
+Thereafter, before solving $G[U]$, we first go over the edges stored in this certificate set and see whether they can be used to find a large matching in $G[U]$.
 If they do, we do not run the random sampling algorithm discussed above.
 If not, the matching that we find must be edge-disjoint.
-We have to be careful that we do not make $S$ too dense though as we spend linear time in the size of $S$.
-Our final algorithm resets $S$ after a certain number of updates.
+We have to be careful that we do not make $H_{\text{cert}}$ too dense though as we spend linear time in the size of $H_{\text{cert}}$.
+Our final algorithm resets $H_{\text{cert}}$ after a certain number of updates.
+We note that our informal discussion of this section hides many (important) details of the final algorithm that we formalize in Section 4.2.
 Overview of our upper bound in Theorem 2:
 To obtain our upper bound of Theorem 2, we partition the matchings into two subsets, $\mathcal{M}$ and $\mathcal{M}^{\prime}$, based on their order in the sequence.
-A key insight is the following: take a vertex $v$ and suppose that it is matched by some matching $M$ in $\mathcal{M}^{\prime}$.
+A key insight (a variant of which was used in the RS upper bound of [29]) is the following: take a vertex $v$ and suppose that it is matched by some matching $M$ in $\mathcal{M}^{\prime}$.
 If we remove all neighbors of vertex $v$ in $\mathcal{M}$, then it can be proved that no vertex of matching $M$ is removed because otherwise there must be an edge from $\mathcal{M}$ that matches two vertices of $M$, violating the inducedness property.
@@
 We assume that each edge update is issued by an adaptive adversary that can see our algorithm’s previous outputs and can accordingly decide on the next update.
+We use $\mu(G)$ to denote the maximum matching size in G.
 We say an algorithm has amortized update-time $U$ if the total time spent after $T$ updates is $U\cdot T$ for some sufficiently large $T=\operatorname{poly}(n)$.
@@
 The following proposition, implied by the streaming algorithm of McGregor [37] (see also [24] for its dynamic adaptation), shows that to find a $(1-\varepsilon)$-approximate matching, it suffices to solve a certain induced matching problem a constant number of times.
-Proposition 3 ([31]).
+Proposition 3 (Approximation Boosting Framework [37, 24]).
 Let $G=(V,E)$ be any (possibly non-bipartite) $n$-vertex graph.
-Suppose that for some parameters $\varepsilon\in(0,1)$ we have an algorithm $\mathcal{A}$ that provided any vertex subset $U\subseteq V$ with $\mu(G[U])\geq\varepsilon n$, finds a matching of size at least $\varepsilon\mu(G[U])$ in $G[U]$.
-Then there is an algorithm that finds a matching of size at least $\mu(G)-O(\varepsilon n)$ in $G$ by making $\widetilde{O}_{\varepsilon}(1)$ adaptive calls to $\mathcal{A}$ where preparing the subset $U$ for each call can be done in $\widetilde{O}_{\varepsilon}(n)$ time.
+Suppose that for any parameter $\delta\in(0,1)$ we have an algorithm $\mathcal{A}(G,U,\delta)$ that provided any vertex subset $U\subseteq V$ with $\mu(G[U])\geq\delta\cdot n$, finds a matching of size at least $\operatorname{poly}(\delta)\cdot\mu(G[U])$ in $G[U]$.
+Then for any $\varepsilon\in(0,1)$, there is an algorithm that finds a matching of size at least $\mu(G)-\varepsilon n$ in $G$ by making $t=O_{\varepsilon}(1)$ adaptive calls
+$\mathcal{A}(G,U_{1},\delta_{1}),\ldots,\mathcal{A}(G,U_{t},\delta_{t})$
+to algorithm $\mathcal{A}$.
+The value of $\delta_{i}$ in each of these calls is just a function of $\varepsilon$, $\delta_{i}\leq\varepsilon/2$, and preparing the vertex subsets $U_{1},\ldots,U_{t}$ can be done in $\widetilde{O}_{\varepsilon}(n)$ total time.
 We also use the following sublinear-time algorithm of Behnezhad [10] for estimating the size of maximum matching.
-We note that even though we use this algorithm, we maintain the edges of the maximum matching explicitly.
+We note that even though we use this algorithm in a crucial, our final dynamic algorithm does not maintain just the size, but rather the edges of the matching, explicitly.
 Proposition 4 ([10]).
@@
 $0.5\mu(G)-\varepsilon n\leq\widetilde{\mu}\leq\mu(G).$
-Finally, we use the following vertex sparsification idea which was adapted to the dynmaic setting in the work of Kiss [29].
-Definition 5.
-Let $V$ be a vertex set and denote $k\leq|V|/2$ be a parameter.
-A set $\mathcal{P}=\{P_{1},\ldots,P_{\ell}\}$ of $\ell$ partitionings of $V$ is a $(\ell,k,\varepsilon)$-vertex sparsifier if the following hold:
-  1. 1.
-    Each partitioning $P_{i}$ partitions $V$ into $O(k/\varepsilon^{2})$ subsets.
-  2. 2.
-    For each subset $S\subseteq V$ of size $2k$, there is at least one partitioning $P_{i}$ such that there are at least $(1-\varepsilon)|S|$ subsets in $P_{i}$ that have exactly one vertex of $S$.
-Proposition 6 ([29]).
-Let $\mathcal{P}$ include $\ell=\lceil 512\log n/\varepsilon^{2}\rceil$ random partitionings $P_{1},\ldots,P_{\ell}$ of $V$, where each $P_{i}$ partitions $V$ into at least $8k/\varepsilon$ subsets randomly.
-Then, with probability $1-1/\operatorname{poly}(n)$, $\mathcal{P}$ is a $(\ell,k,\varepsilon)$-vertex sparsifier of $V$.
-The notion of graph contraction here refers to the process of simplifying the graph based on the partitionings provided by the vertex sparsifiers.
-Each set of partitions effectively represents a contracted version of the graph, where the vertices within the same subset can be considered as a single entity for the purpose of matching.
-This contraction not only reduces the overall size of the graph but also retains the essential structural properties that influence the size and composition of maximum matchings.
-Definition 7.
-Denote $M$ be a given matching with $2r$ vertices in graph $G$.
-Let us assume the $d_{v}$ for a vertex $v$ in the matching is the number of neighbors of $v$ in $V(M)$.
-we define average degree of $M$ as $\sum_{v\in V(M)}d_{v}/2r$.
-4 Dynamic Approximate Matching
-We present our dynamic matching algorithm in this section and prove Theorem 1.
-Let us start with an overview of how our proposed algorithm is built.
-We randomly sample edges from graph $G$ to find a matching with $\varepsilon$-approximation, which is described in Section 4.1.
-We propose that if we need many samples to find this matching then the average degree of this matching is small in the remaining edges (what we mean by the remaining edges is that, when we output a matching, the adversary can remove its edges from the graph, while searching for a new matching we should discard the edges found before), resulting in the matching to be induced in the remaining graph.
-In this case, we charge the cost of the algorithm to ORS, with characteristics similar to the matching we found.
-Otherwise, we will succeed in finding a matching, relatively fast.
-Note that, in this model, we can argue that if the matching size in the graph is $\mu$ then the matching size will not change after $\varepsilon^{2}n$ updates.
-We will find a matching once in $\varepsilon^{2}n$ edge updates to optimize amortized running time.
-This motivates us to run Section 4.1 to find a matching after $\varepsilon^{2}n$ updates.
-To keep track of the large matchings we already found we introduce Section 4.2.
-This algorithm simplifies our problem where the matching size does not change.
-If we maintain this algorithm for different matching sizes, knowing the approximate matching we will know which maintained algorithm to look at to find the $\varepsilon$-approximate matching.
-To boost our approximation to $(1-\varepsilon)$ we use Proposition 3 to output the subsets of $U$ that we need to search for the matchings of size $\varepsilon\mu$.
-We use the subsets in Section 4.1 which is a $\varepsilon$-approximation algorithm.
-Note that we provide a multiplicative approximation maximum matching algorithm for the dynamic model, which leads to the fact that we need to handle the cases where the matching size is not of order $n$.
-This adds some layers to the intuition we stated above.
-  1. 1.
-    If the matching size is $n^{\prime}$ we output a matching after $\varepsilon^{2}n^{\prime}$ updates.
-This means we need to handle matching with similar sizes together, we will have $\log n$ partition to keep track of the matchings outputted by Section 4.1 of almost the same size simultaneously.
-  2. 2.
-    To handle the graph with small matchings of $O(n^{\prime})$ we need to contract the graph to one with $n^{\prime}$ vertices.
-We build a contracted graph by creating $n^{\prime}$ buckets of $O(n/n^{\prime})$ vertices, the vertices are distributed randomly in buckets.
-Since we want to preserve the edges of the maximal matching we need to randomly maintain $O(\log n)$ such contractions for $\log n$ values of $n^{\prime}$.
-  3. 3.
-    The fact that the edges of the matching found in the earlier steps of Section 4.2, could be removed by the adversary leads us to a definition of the degree of matching that only takes into account the matchings that we will be finding in the next rounds (we do not care about the edges removed from the graph in the dynamic steps).
-This concept leads us to define the set of matchings named ORS, where every matching $M$ is induced in the set of matchings we observe after $M$.
-4.1 Random Sampling Algorithm
-Here we introduce a sampling algorithm that given a graph $G$, with a matching of size $\mu$ can output a matching of size $\varepsilon\mu$.
-We bound the algorithm running time with the average degree of the outputted matching.
-The way the algorithm works is that for each vertex we sample $b$ vertices and check if the edge is in the graph using the adjacency matrix of the graph.
-We add the edges to the matching if none of the endpoints have been used previously.
-Note that by starting with a small $b$ for the budget, we ensure to find the edges from dense parts of the graph which is easier to find.
-As we go further in the algorithm we make the budget larger to find edges in more sparse parts of the graph.
-Now, we state the random sampling algorithm and prove its efficiency.
+Finally, we use the following proposition which turns an additive approximation into a multiplicative approximation.
+The proof is based on a vertex sparsification idea for matchings [5, 14] which was adapted to the dynamic setting in the work of Kiss [34].
+Proposition 5 ([34, 5]).
+Suppose there is an adaptive algorithm $\mathcal{A}$, that for any parameter $\varepsilon>0$ and a fully dynamic $n$-vertex graph $G=(V,E)$, maintains a matching of size $\mu(G)-\varepsilon n$ in $Q(n,\varepsilon)$ amortized time.
+Then there is an algorithm that maintains a multiplicative $(1-\varepsilon)$-approximation maximum matching of $G$ in $\operatorname{poly}(\log(n),\varepsilon)\cdot Q(n,\varepsilon^{2})$ amortized time per update.
+4 Our Dynamic Algorithm for Approximate Maximum Matching
+In this section, we present our algorithm and prove Theorem 1.
+We start in Section 4.1 with a static potentially sublinear-time algorithm that is one of the main building blocks of our final algorithm.
+We then formalize our dynamic algorithm in Section 4.2.
+In Section 4.3 we prove correctness of our dynamic algorithm and analyze its running time in Section 4.4.
+4.1 A Static Potentially-Sublinear Time Algorithm via Random Sampling
+In this section, we present a potentially sublinear time algorithm that given an $n$ vertex graph $G=(V,E)$, finds a matching $M$ that is by an additive factor of at most $\varepsilon n$ smaller than the maximum matching of $G$.
+The algorithm, in addition, returns a certificate $M_{C}$, which is another matching of $G$ that we use to bound the running time of the algorithm.
+Particularly, denoting by $d$ the average degree in the subgraph induced on the vertices of the certificate (i.e., graph $G[V(M_{C})]$), the running time of the algorithm overall is $\widetilde{O}(n^{2+2\varepsilon}/d)$.
+If the certificate $M_{C}$ is close to an induced matching (i.e., $d$ is small), then the algorithm is not better than the trivial algorithm which reads all the edges in quadratic in $n$ time.
+On the other hand, if $G[V(M_{C})]$ is dense, the algorithm runs in sublinear time.
+This algorithm will be a crucial component of our final dynamic algorithm.
+The following is the main lemma of this section.
 Lemma 6.
-Given a graph $G=(V,E)$, vertex subset $U\subseteq V$, and parameters $\varepsilon\in(0,1)$, such that maximal matching of $G[U]$ is $\mu\geq\varepsilon n$, there is an algorithm that finds a matching $M\subseteq G[U]$ of size at least $\varepsilon\mu$.
-The algorithm runs in time $O(n^{2+\varepsilon}/d)$ where $d=|E\cap V^{2}(M)|/O(n)$.
+Let $G=(V,E=E_{\text{dense}}\cup E_{\text{sparse}})$ be a given $n$-vertex graph where we have adjacency matrix access to $E_{\text{dense}}$, adjacency list access to $E_{\text{sparse}}$, and $E_{\text{dense}}$ and $E_{\text{sparse}}$ may share edges.
+For any parameter $\varepsilon\in(0,1)$, there is an algorithm $\textsc{MatchAndCertify}(E_{\text{dense}},E_{\text{sparse}},\varepsilon)$ that returns a $(1,\varepsilon n)$-approximate maximum matching $M$ of $G$ and a certificate $M_{C}$.
+Let $T$ be the running time of the algorithm.
+Then exactly one of the following two conditions holds:
+  1. $(C1)$
+    $M_{C}=\emptyset$ and $T=\widetilde{O}_{\varepsilon}(n+|E_{\text{sparse}}|)$.
+  2. $(C2)$
+    $M_{C}$ is a matching in $E_{\text{dense}}$ of size $\Omega_{\varepsilon}(n)$ where $M_{C}\cap E_{\text{sparse}}=\emptyset$.
+Additionally, $T=\widetilde{O}_{\varepsilon}(|E_{\text{sparse}}|+n^{2+\varepsilon}/d)$ where $d:=|E_{\text{dense}}\cap V^{2}(M_{C})|/|M_{C}|$ is the average degree of $E_{\text{dense}}[V(M_{C})]$.
+Towards proving Lemma 6, we first prove the following Lemma 7 which solves a slightly simpler subproblem that finds a large matching in a given induced subgraph of $G$.
+We will later prove Lemma 6 by combining Lemma 7 with Proposition 3.
+Lemma 7.
+Let $G=(V,E)$ be a given $n$-vertex graph to which we have adjacency matrix access.
+Let $U\subseteq V$ be a given vertex subset, and let $\delta\in(0,1)$ be a given parameter such that $\mu(G[U])\geq\delta n$.
+There is an algorithm that finds a matching $M\subseteq G[U]$ of size at least $\delta^{2}n$.
+The algorithm runs in $\widetilde{O}(n^{2+2\delta}/d)$ time where $d=|E\cap V^{2}(M)|/|M|$.
 Proof.
+The algorithm is formalized below as Section 4.1.
 […]
-Let us consider Section 4.1, first we prove that this algorithm will provide a matching $M_{c}$ of size a least $\varepsilon\mu/2$ by sampling vertices with budget $b\in[2^{\varepsilon c\cdot\log n},2^{\varepsilon(c+1)\cdot\log n}]$.
-Let us assume there are no such intervals.
-This means that for any integer $c\in[0,1/(2\varepsilon)]$, running the algorithm for $b$ in the interval $[2^{\varepsilon c\cdot\log n},2^{\varepsilon(c+1)\cdot\log n}]$ outputs a matching with less than $\varepsilon\mu$ edges.
-Note that for each edge $e$ in $M_{1}\cup\dots\cup M_{1/(2\varepsilon)}$, at most 2 edges in the maximum matching of $G[U]$ have common vertices with $e$.
-Summing over all the $1/(2\varepsilon)$ intervals we have:
-$22\sum_{c=1}^{1/(2\varepsilon)}{|M_{i}|}<2\cdot\frac{1}{2\varepsilon}\cdot\varepsilon\mu/2=\mu/$ 2 .
-The fact that we could not find a matching of size $\mu/2$ is a contradiction with the lemma statement ensuring that the size of maximum matching is at least $\mu/2$.
-Note that by exhausting the algorithm with budget of $n$, we will go through all the edges and output a maximal matching.
-Note that any maximal matching in this graph should be of size at least $\mu/2$.Therefore, we have at least the size of one of the matchings in$\{M_{1},...,M_{1/(2\varepsilon)}\}$ is $\varepsilon\mu/2$.
-Now, we prove the statement of the algorithm.
-We assume that this algorithm terminates at $b$ such that $2^{\varepsilon(c-1)\cdot\log n}\leq b<2^{\varepsilon c\cdot\log n}$, for a constant $\delta$, we spend $n[b+b/2+..+1]<\delta nb\leq\delta n\cdot 2^{\varepsilon c\log n}<\delta n^{1+\varepsilon c}$ time to find the matching $M_{c}$ such that $|M_{c}|\geq\varepsilon\mu/2$.
-Note that if it takes at most $n^{2+\varepsilon}/d$ time to find the matching of size $\varepsilon\mu/2$, then we have $n^{2+\varepsilon}/d<\delta n^{1+\varepsilon c}$ therefore, $d>\frac{n^{1+\varepsilon-\varepsilon c}}{\delta}\geq\frac{n^{1+\varepsilon}}{\delta}$.
-For budget $b$, we sample vertices for all the vertices, we argue that for the vertices that are not in $M$, and after sampling $b$ vertices we could not find their neighbor then the degree of vertices out of the matching is less than $O(n/b)$.
-Assuming that the degree of vertex $v$ is $d_{v}=c\cdot\frac{n}{b}$, with probability $(1-\frac{d_{v}}{n})^{b}=(1-c\cdot\frac{1}{b})^{b}\approx(1/e)^{c}$ we sampled a neighbor.
-Therefore in this case with a constant probability we already matched $v$.
-This implies that if we find the first edge of this matching at a budget of $2b$, we get that the degree of any vertex out of the matching we found until this point is less than $c\cdot\frac{n}{b}$ for any constant $c$.
-Picking $c$ small enough, the total sum of the degree of the matching is at most $\varepsilon\mu c\cdot\frac{n}{b}<\varepsilon\mu n^{(1-\varepsilon(c-1))}=O(nd)$.
+Note that if the algorithm returns a matching $M_{i}$, then it is guaranteed that the matching has our desired size of at least $\delta^{2}n$.
+So it suffices to prove that at some point $M_{i}$ is as large as we want.
+To see this, note that once the budget $b$ reaches $n$, we will go through all the edges, indicating that $M$, the union of matching $M_{1},M_{2},\dots,M_{1/(2\delta)}$, is a maximal matching of $G[U]$.
+Since the size of a maximal matching is at least half the size of a maximum matching, we have $|M|\geq\mu(G[U])/2\geq\delta n/2$.
+This means that at least one of the matchings in $\{M_{1},...,M_{1/(2\delta)}\}$ is of size at least $\frac{\delta^{2}n/2}{1/(2\delta)}=\delta^{2}n$.
+Next, we focus on relating the running time of the algorithm to the parameter $d$, i.e., the average degree of the output matching of Section 4.1.
+Let us first introduce some notation.
+Let $U^{\prime}_{i}$ be the value of the set $U^{\prime}$ at the end of iteration $i$.
+Additionally, for every vertex $v$, let $d^{i}_{v}$ be the degree of vertex $v$ in graph $G[U^{\prime}_{i}]$.
+We start by proving the following useful claim.
+Claim 8.
+It holds with probability $1-1/n$ for every $v\in U^{\prime}_{i}$ and every $i$ that $d^{i}_{v}=O(n^{1-2\delta(i-1)}\log n)$.
+Proof.
+Fix $v$ and $i$.
+We prove that either $v\not\in U^{\prime}_{i}$ or $d^{i}_{v}=O(n^{1-2\delta(i-1)}\log n)$ with probability at least $1-1/n^{3}$.
+A union bound over the choices of $v$ and $i$ completes the proof.
+Consider iteration $i$ of the algorithm and suppose $v\in U^{\prime}_{i}$.
+This means that in this iteration we attempt to match vertex $v$ by sampling $b=n^{2\delta(i-1)}$ vertices in $U^{\prime}\supseteq U^{\prime}_{i}$ and checking if any of them is a neighbor of $v$.
+Let $d^{\prime\prime}$ be the number of neighbors of $v$ in $U^{\prime}$ when processing $v$ in iteration $i$.
+We have
+$\Pr[v\text{ not matched by }M_{i}]\leq\left(1-\frac{d^{\prime\prime}}{|U^{\prime}|}\right)^{b}\leq\left(1-\frac{d^{\prime\prime}}{n}\right)^{b}\leq\exp\left(-\frac{d^{\prime\prime}b}{n}\right)=\exp\left(-\frac{d^{\prime\prime}}{n^{1-2\delta(i-1)}}\right).$
+Now consider two cases.
+If $d^{\prime\prime}\geq 3n^{1-2\delta(i-1)}\log n$, then $v$ is matched in $M_{i}$ with probability at least $1-1/n^{3}$ and thus will be removed from $U^{\prime}$.
+Otherwise, since $d^{i}_{v}\leq d^{\prime\prime}$, we have $d^{i}_{v}=O(n^{1-2\delta(i-1)}\log n)$.
+Thus, in either case, the statement holds.
 ∎
-4.2 The Overall Algorithm via Certificates and Caching Them
-In the next step, we aim to use Section 4.1 efficiently to output an approximate matching in the dynamic model.
-First, we prove that given the matching size is equal to $O(n^{\prime})$, after $\varepsilon^{2}n^{\prime}$ updates the matching size does not change substantially.
-This means we can output a matching after each $\varepsilon n^{\prime}$ update and get an $\varepsilon$-approximate matching.
-Observation 9.
-Given a graph $G=(V,E)$ with $n$ vertices, if the maximum matching size is $O(\mu)$ and we provide an $\varepsilon$-approximate matching $M^{\prime}$, after $\varepsilon^{2}\mu$ updates in the dynamic model, $M^{\prime}$ or one of its subsets remains an $\varepsilon$-approximate matching for the updated graph.
+Let $M_{i}$ be the matching that Section 4.1 returns.
+Note that in iteration $j$ we have $b=n^{2(j-1)\delta}$ and the algorithm takes $O(nb)=O(n^{1+2(j-1)\delta})$ time.
+So the running time of the algorithm until termination is $O(n)+O(n^{1+2\delta})+\ldots+O(n^{1+2(i-1)\delta})=O(n^{1+2(i-1)\delta})$.
+Since every vertex $v$ matched by $M_{i}$ must belong to $U^{\prime}_{i-1}$, applying Claim 8 gives $d^{i-1}_{v}=O(n^{1-2\delta(i-2)}\log n)$.
+This implies that the total number of edges in $G[V(M_{i})]$ can be upper bounded by $|M|\cdot O(n^{1-2\delta(i-2)}\log n)$.
+Letting $d=\Theta(n^{1-2\delta(i-2)}\log n)$, the running time can be re-written as
+$O(n^{1+2(i-1)\delta})=O(n^{1+2(i-1)\delta}d/d)=O(n^{1+2(i-1)\delta+1-2\delta(i-2)}\log n/d)=O(n^{2+2\delta}\log n/d),$
+which is the claimed bound. ∎
+Next, we apply Proposition 3 on Lemma 7 to complete the proof of Lemma 6.
+Proof of Lemma 6.
+For any graph $H$, let GreedyMatching($H$) be a greedy algorithm that goes over all the edges of $H$ one by one and adds an edge $(u,v)$ to the matching if $u$ and $v$ are not matched already.
+Note that the output of this algorithm is a maximal matching in $H$.
+The algorithm MatchAndCertify proceeds as follows.
+Our goal is to run the algorithm of Proposition 3 to find a matching of size $\mu(G)-\varepsilon n$ in $G$.
+This algorithm makes $t=O_{\varepsilon}(1)$ adaptive calls
+$\mathcal{A}(G,U_{1},\delta_{1}),\dots,\mathcal{A}(G,U_{t},\delta_{t})$
+to an algorithm $\mathcal{A}$ that outputs a matching of size $\operatorname{poly}(\delta_{i})\cdot\mu(G[U_{i}])$ in $G[U_{i}]$ provided that $\mu(G[U_{i}])\geq\delta_{i}n$.
+Here $U_{i}$ is a subset of $V$ and parameter $\delta_{i}\leq\varepsilon$ is a function of $\varepsilon$.
+In order to use Proposition 3, we have to specify how the algorithm $\mathcal{A}$ is implemented.
+We implement $\mathcal{A}(G,U_{i},\delta_{i})$ as follows.
+Let $G_{i}:=G[U_{i}]\cap E_{\text{sparse}}$.
+As the first step, we call $\textsc{GreedyMatching}(G_{i})$ to obtain matching $M_{i}$ in $O(|E_{\text{sparse}}|)$ time.
+If $|M_{i}|\geq\delta^{2}_{i}\cdot n/8$, then we have a large enough matching in $G[U_{i}]$, and $\mathcal{A}$ terminates by outputting $M_{i}$ and we set $d_{i}=\infty$.
+Otherwise, we run $\textsc{RandomSampling}(G[U_{i}]\cap E_{\text{dense}},\delta_{i})$ from Lemma 7.
+Note that Lemma 7 requires adjacency matrix access to its input graph which we can provide, given the guarantee of the lemma that we have adjacency matrix access to $E_{\text{dense}}$ and that the set $U_{i}$ is given explicitly.
+Since GreedyMatching returns a 1/2-approximate maximum matching, we get that
+$\mu(G[U_{i}]\cap E_{\text{dense}})\geq\mu(G[U_{i}])-\mu(G[U_{i}]\cap E_{\text{sparse}})\geq\delta_{i}n-2|M_{i}|=3\delta_{i}n/4.$
+Therefore, algorithm RandomSampling outputs a matching $M_{C_{i}}$ of size at least $169\delta^{2}_{i}n/$ 16 in $G[U_{i}]\cap E_{\text{dense}}$.
+By Lemma 7, this runs in $\widetilde{O}(n^{2+2\delta_{i}}/d_{i})=\widetilde{O}(n^{2+\varepsilon}/d_{i})$ time where $d_{i}=|E\cap V^{2}(M_{C_{i}})|/|M_{C_{i}}|$.
+From our discussion above, in the $i$-th call to $\mathcal{A}$, this algorithm outputs a matching of size at least $\min\{\delta_{i}^{2}n/8,9\delta_{i}^{2}n/16\}=\delta_{i}^{2}n/8$.
+Therefore, the algorithm of Proposition 3 outputs a matching $M$ of size $\mu(G)-\varepsilon n$.
+It remains to analyze the running time by specifying the outputs $d$ and $M_{C}$ and showing that at least one of the conditions $(C1)$ or $(C2)$ hold.
+If we never call the RandomSampling algorithm, we simply set $d=0$ and $M_{C}=\emptyset$ as all calls to $\mathcal{A}$.
+The total running time spent by Proposition 3 preparing the vertex subsets $U_{i}$ for each $i\in[t]$, and the running time of the calls to the GreedyMatching is
+$\widetilde{O}_{\varepsilon}(n)+\sum_{i\in[t]}O(|E(G_{i})|)=\widetilde{O}_{\varepsilon}(n)+t|E_{\text{sparse}}|=\widetilde{O}_{\varepsilon}(n+|E_{\text{sparse}}|).$
+Thus, condition $(C1)$ holds.
+Otherwise, we let $j=\operatorname*{arg\,min}_{i\in[t]}d_{i}$, set $d:=d_{j}$ and return the certificate $M_{C}=M_{C_{j}}\setminus E_{\text{sparse}}$. We now show that condition $(C2)$ holds.
+Let us now bound the running time.
+This is the cost of preparing the vertex subsets of Proposition 3, the running time of the calls to the GreedyMatching, and the running time of the calls to RandomSampling.
+The total running time is therefore
+$\widetilde{O}_{\varepsilon}(n)+t|E_{\text{sparse}}|+\widetilde{O}\left(\sum_{i=1}^{t}\frac{n^{2+\varepsilon}}{d_{i}}\right)=\widetilde{O}_{\varepsilon}\left(|E_{\text{sparse}}|+\frac{n^{2+\varepsilon}}{d}\right).$
+The bound above is the one required by $(C2)$, except that we defined $d=d_{j}=|E\cap V^{2}(M_{C_{j}})|/|M_{C_{j}}|$ whereas the certificate $M_{C}$ is a subset of $M_{C_{j}}$.
+To fix this, in what follows we prove that $d$ is at least a third of the value of the right parameter $d^{\prime\prime}:=|E\cap V^{2}(M_{C})|/|M_{C}|$, which completes the proof.
+The fact that we run RandomSampling on $G[U_{j}]$ implies that $\textsc{GreedyMatching}(G_{j})$ did not return a large enough matching.
+This implies that $\mu(G_{j})\leq\delta^{2}_{j}n/4$.
+Therefore, once we remove the edges of $E_{\text{sparse}}$ from $M_{C_{j}}$, we reduce the size of this matching by at most $\delta^{2}_{j}n/4$.
+Combined with $|M_{C_{j}}|\geq 9\delta_{i}^{2}n/16$, this implies that $|M_{C}|\geq|M_{C_{j}}|-\delta_{j}^{2}/4\geq 7|M_{C_{j}}|/16$.
+Hence
+$d^{\prime\prime}=\frac{|E\cap V(M_{C})|}{|M_{C}|}\leq\frac{|E\cap V(M_{C_{j}})|}{7|M_{C_{j}}|/16}<3d,$
+completing the proof as discussed above. ∎
+4.2 Our Dynamic Algorithm via Lemma 6
+In this section, we formalize our dynamic algorithm for Theorem 1 based on the static MatchAndCertify algorithm of Lemma 6. Our algorithm is formalized as Section 4.2.
+We prove in Section 4.3 that our algorithm maintains an additive $(1,\varepsilon n)$-approximation of the graph at anytime.
+Lemma 9 (Correctness of Section 4.2).
+At any point, the output matching $M_{G}$ of Section 4.2 is a $(1,\varepsilon n)$ approximate maximum matching of graph $G$.
+Later in Section 4.4, we analyze the running time of the algorithm and prove the following bound on it.
+Lemma 10 (Runtime of Section 4.2).
+Section 4.2 runs in time
+$\widetilde{O}_{\varepsilon}\left(\frac{t}{n}+\frac{n^{2+\varepsilon}\cdot\mathsf{ORS}(\Theta_{\varepsilon}(n))}{t}\right),$
+where $t$ is the threshold of the algorithm.
+The combination of the two lemmas above with Proposition 5 implies Theorem 1.
+Proof of Theorem 1.
+The running time of Section 4.2 is analyzed in Lemma 10, which gives a boundo f
+$\widetilde{O}_{\varepsilon}\left(\frac{t}{n}+\frac{n^{2+\varepsilon}\cdot\mathsf{ORS}_{n}(\Theta_{\varepsilon}(n))}{t}\right),$
+where $t$ is a threshold parameter. To achieve the amortized time, we balance the two terms in the running time expression by setting
+$t=\sqrt{n^{3+\varepsilon}\cdot\mathsf{ORS}_{n}(\Theta_{\varepsilon}(n))}.$
+By substituting this value of $t$ we have
+$\widetilde{O}_{\varepsilon}\left(\frac{\sqrt{n^{3+\varepsilon}\cdot\mathsf{ORS}_{n}(\Theta_{\varepsilon}(n))}}{n}+\frac{n^{2+\varepsilon}\cdot\mathsf{ORS}(\Theta_{\varepsilon}(n))}{\sqrt{n^{3+\varepsilon}\cdot\mathsf{ORS}_{n}(\Theta_{\varepsilon}(n))}}\right)=\widetilde{O}_{\varepsilon}\left(\sqrt{n^{1+\varepsilon}\cdot\mathsf{ORS}(\Theta_{\varepsilon}(n))}\right).$
+By Lemma 9, Section 4.2 maintains a matching of size $\mu(G)-\varepsilon n$ at any point.
+By applying Proposition 5, we can boost this additive approximation to a multiplicative $(1-\varepsilon)$-approximation while keeping the update time $\widetilde{O}_{\varepsilon}\left(\sqrt{n^{1+\varepsilon}\cdot\mathsf{ORS}(\Theta_{\varepsilon}(n))}\right).\qed$
+[…]
+4.3 Correctness of Section 4.2
+In this section, we prove Lemma 9.
+Proof of Lemma 9.
+The algorithm works as follows.
+$c_{\text{updates}}$ is a counter for the number of updates.
+Let a phase be $t$ consecutive updates starting with $c_{\text{updates}}=0$.
+The algorithm maintains the fully dynamic graph $G$.
+All the insertions throughout a phase are added in $G_{\text{add}}$ (if an inserted edge is deleted, we delete it from $G_{\text{add}}$ too).
+All the deletions throughout a phase are added to $G_{\text{del}}$ (if a deleted edge is later inserted, we delete it from $G_{\text{del}}$ too).
+First, we prove that MatchAndCertify outputs a $(1,\varepsilon n)$-approximate matching of graph $G$.
+To do this, we prove that the union of $E_{\text{dense}}=G-G_{\text{add}}$ and $E_{\text{sparse}}=G_{\text{add}}\cup H_{\text{cert}}-G_{\text{del}}$ is exactly $G$.
+To see this, note that
+$E_{\text{dense}}\cup E_{\text{dense}}=(G-G_{\text{add}})\cup(G_{\text{add}}\cup H_{\text{cert}}-G_{\text{del}})=G\cup(H_{\text{cert}}-G_{\text{del}})=G.$
+The second inequality holds because $G_{\text{add}}\cap G_{\text{del}}=\emptyset$.
+The last equality holds because any edge $e\in H_{\text{cert}}$ either belongs to $G$ or is deleted at some point, which means $H_{\text{cert}}-G_{\text{del}}\subseteq G$.
+By Lemma 6, this implies that MatchAndCertify outputs a $(1,\varepsilon n/2)$-approximate matching $M$ for $G$.
+Once $M$ is computed, we do not change $M$ within the next $\varepsilon n/2$ updates unless by applying the deletions on it.
+Each such deletion reduces the size of $M$ by at most 1.
+Therefore, after $\varepsilon n/2$ updates, we have $|M|\geq\mu(G)-\varepsilon n$, thus $M$ remains a $(1,\varepsilon n)$-approximate matching for $G$ throughout.
+∎
+4.4 Runtime Analysis of Theorem 1
+Our goal in this section is to prove Lemma 10.
+Namely, we want to bound the running time of Section 4.2 by relating it to density of ORS graphs.
+The following Lemma 11 is our main tool for this connection to ORS graphs.
+Lemma 11.
+Let $G=(V,E=M_{1}\cup\ldots\cup M_{t})$ be an $n$-vertex graph where $M_{1},\ldots,M_{t}$ are edge-disjoint matchings of size at least $r$ each.
+For any $i\in[t]$, define $G_{i}=(V,E_{1}\cup\ldots\cup E_{i})$ and let $d_{i}$ be $|E[G_{i}]\cap V^{2}(M_{i})|/|M_{i}|$.
+Then for any $\alpha\in(0,1)$, it holds that
+$\sum_{i=1}^{t}\frac{1}{d_{i}}=O\left(\frac{1}{\alpha^{2}}\cdot\mathsf{ORS}_{n}(r(1-\alpha))\cdot\log n\right).$
+Let us first prove Lemma 10 via Lemma 11. The proof of Lemma 11 is presented afterwards.
+Proof of Lemma 10.
+All the steps of Section 4.2 can be implemented in $\widetilde{O}(1)$ time per update except for Section 4.2 where we spend $\widetilde{O}(t)$ time and Section 4.2 where we call MatchAndCertify.
+Since we call Section 4.2 every $t$ updates, the amortized cost of this line is still $\widetilde{O}(1)$.
+It remains to analyze the cost of Section 4.2.
+Note that Section 4.2 is called every $\varepsilon n/2$ updates, so its cost will have to be amortized over these $\varepsilon n/2$ updates.
+Let us first discuss the time needed to prepare the inputs to MatchAndCertify.
+This input consists of graphs $E_{\text{dense}}=G-G_{\text{add}}$ and $E_{\text{sparse}}=G_{\text{add}}\cup H_{\text{cert}}-G_{\text{del}}$.
+For $E_{\text{dense}}$ we only need to provide adjacency matrix query access.
+Since each such query can be answered by making one adjacency matrix query to $G$ and one to $G_{\text{add}}$, we do not need to spend any time preparing $E_{\text{dense}}$.
+For $E_{\text{sparse}}$, we go over the edges of $G_{\text{add}},H_{\text{cert}},G_{\text{del}}$ one by one, but since these graphs are reset after $t$ updates, they only contain at most $t$ edges and the time spent is $O(t)$ which amortized over $\varepsilon n$ updates adds up to $O_{\varepsilon}(t/n)$.
+Now, let us compute the running time of the algorithm based on the total calls to the MatchAndCertify.
+Each call to this algorithm, by Lemma 6 satisfies one of the two conditions $(C1)$ and $(C2)$.
+Each call that satisfies $(C1)$ runs in $\widetilde{O}_{\varepsilon}(n+|E_{\text{sparse}}|)$ time.
+Each call that satisfies $(C2)$ runs in $\widetilde{O}_{\varepsilon}(|E_{\text{sparse}}|+n^{2+\varepsilon}/d)$ time.
+Let us for now ignore the second term in the latter running time.
+The rest of the computation, amortized over $\varepsilon n$ updates, sums up to $\widetilde{O}_{\varepsilon}(|E_{\text{sparse}}|/n)=\widetilde{O}_{\varepsilon}(t/n)$.
+Let us define a phase of the algorithm to be $t$ consecutive updates while $c_{\text{updates}}<t$.
+For analyzing the second term in the running time of the calls to MatchAndCertify that satisfy condition $(C2)$ of Lemma 6, we analyze their total cost over the whole phase and amortize this over $t$.
+Note that this is very different from our analysis before that amortizes the cost over a shorter sequence of $\varepsilon n$ updates.
+Take the $i$-th call to Lemma 6 within the phase that satisfies $(C2)$ and let $M_{C_{i}}$ be its certificate, which is a matching in $E_{\text{dense}}$ and let $d_{i}:=|E_{\text{dense}}\cap V^{2}(M_{C_{i}})|/|M_{C_{i}}|$ be the average degree of $M_{C_{i}}$ in $E_{\text{dense}}$ the moment that it is returned.
+Note that the total cost the part of the algorithm that we ignored above can be written as
+$\sum_{i}\widetilde{O}_{\varepsilon}(n^{2+\varepsilon}/d_{i})=\widetilde{O}_{\varepsilon}(n^{2+\varepsilon})\cdot\sum_{i}1/d_{i}.$ (1)
+In what follows, we upper bound the sum $\sum_{i}1/d_{i}$.
+Note that the set $E_{\text{dense}}$ used in the definition of $d_{i}$ does not remain the same throughout a phase, and particularly differs for each $d_{i}$.
+But here is the crucial observation.
+Since $E_{\text{dense}}=G-G_{\text{add}}$, the set $E_{\text{dense}}$ is decremental throughout the phase.
+This is because every edge inserted is inserted to both $G$ and $G_{\text{add}}$, thus $G-G_{\text{add}}$ remains unchanged, while deletions from $G$ will be deleted from $E_{\text{dense}}$ as well.
+Now define $H_{\geq i}:=M_{C_{i}}\cup M_{C_{i+1}}\cup\ldots\cup M_{C_{k}}$.
+From our earlier discussion that the $M_{C_{i}}$’s are matchings in a decremental graph, all edges in $H_{\geq i}$ were present when certificate $M_{C_{i}}$ was outputted by Lemma 6.
+Hence, for the average degree $d^{\prime}_{\geq i}$ of $M_{C_{i}}$ in $H_{\geq i}$ we have
+$d^{\prime}_{\geq i}:=\frac{|E(H_{\geq i})\cap V(M_{C_{i}})^{2}|}{|M_{C_{i}}|}\leq\frac{|E_{\text{dense}}\cap V(M_{C_{i}})^{2}|}{|M_{C_{i}}|}=d_{i},$ (2)
+where here $E_{\text{dense}}$ is the input to Lemma 6 at the step where certificate $M_{C_{i}}$ was returned.
+Therefore, applying Lemma 11 on the sequence $M_{C_{k}},M_{C_{k-1}},\ldots,M_{C_{1}}$ (note the reverse order), we have
+[…]
+It should be noted that our Lemma 11 requires the matchings in the sequence to be edge-disjoint.
+This is indeed satisfied by our algorithm, since once a certificate $M_{C_{i}}$ is returned, we add its edges to $H_{\text{cert}}$ and these edges will never belong to $E_{\text{dense}}$ for the rest of the phase.
+Plugged into 1, this bounds the total (remaining) running time of a phase by
+$n^{2+\varepsilon}\cdot\widetilde{O}(\mathsf{ORS}_{n}(\Theta_{\varepsilon}(n)).$
+Amortized over $t$, the number of updates of the phase, this sums up to
+$\frac{n^{2+\varepsilon}\cdot\widetilde{O}(\mathsf{ORS}_{n}(\Theta_{\varepsilon}(n))}{t}.$
+Combined with the earlier bound, we arrive at the claimed amortized update time of
+$\widetilde{O}_{\varepsilon}\left(\frac{t}{n}+\frac{n^{2+\varepsilon}\cdot\mathsf{ORS}_{n}(\Theta_{\varepsilon}(n))}{t}\right).\qed$
+We now turn to prove Lemma 11. First we prove the following auxiliary claim that helps with the proof of Lemma 11.
+Claim 12.
+Let $G=(V,E=M_{1}\cup\ldots\cup M_{t})$ be an $n$-vertex graph where $M_{1},\ldots,M_{t}$ are edge-disjoint matchings of size at least $r$ each.
+For any $i\in[t]$, define $G_{i}=(V,E_{1}\cup\ldots\cup E_{i})$ and let $d_{i}$ be $|E[G_{i}]\cap V^{2}(M_{i})|/|M_{i}|$.
+Let $d$ be an upper bound on $d_{i}$ for all $i\in[t]$.
+Then for any $\alpha\in(0,1)$, it holds that
+$t=O\left(\frac{d}{\alpha^{2}}\cdot\mathsf{ORS}_{n}(r(1-\alpha))\right).$
 Proof.
-We assume that in a graph $G$ the maximal matching size is $c\mu$ where $c$ is a constant, after $\varepsilon^{2}\mu$ updates in the graph the size of the maximal matching is between $(c-\varepsilon^{2})\mu,(c+\varepsilon^{2})\mu$.
-This implies that if we have an algorithm that outputs an $\varepsilon c\mu$ matching, after $\varepsilon^{2}\mu$ updates, the $\varepsilon$ approximate matching is preserved.
-This is because we lose at most $\varepsilon^{2}$ fraction of the outputted matching.
-∎
-Note that to use Section 4.1 we need to know about the existence of a matching $\mu$, the algorithm described in Proposition 4, given a graph with $n$ vertices, outputs if there is matching of size $O(n)$.
-Using this algorithm alone will not help us with the cases where the matching size is not of order $n$.
-This motivates us to find an algorithm that reduces the number of vertices on the graph while maintaining the maximum matching size.
-Proposition 6 will maintain $O(\log n)$ contraction on the graph for an integer $k\in[1,\log n]$, such that for any matching $M$ with high probability, there is a contraction preserving a $1-\varepsilon$ approximation of $M$.
-After running the reduction on the graph, we find the contraction that preserves the maximum matching.
-Let us assume that $G^{\prime}$ is the contracted graph, we run Section 4.1 on $G^{\prime}$ to output the $\varepsilon$ approximation matching.
-Note that throughout the changes in the dynamic model, both the matching size and the contracted graph that preserves the maximum matching, change.
-An edge insertion in $G$ is a weight addition to an edge in graph $G^{\prime}$, and edge deletion is a weight subtraction in $G^{\prime}$.
-In the following algorithms, when we find a matching in $G^{\prime}$, we do not care about the weight of edges, an edge in $G^{\prime}$ exists if its weight is more than 0, otherwise, This edge does not exist in $G^{\prime}$.
-First, we solve a simplified version of the problem where the matching size does not change throughout the updates, and a single contracted graph $G^{\prime}$ preserves matching edges in the dynamic model.
-We prove that there exists an efficient algorithm maintaining a $(1-\varepsilon)$-approximation matching in the dynamic model.
-We introduce Section 4.2.
-We assume that the updates in this algorithm are in the case that the matching size in $O(n^{\prime})$, and $G^{\prime}$ preserves the maximum matching of the graph.
-This algorithm works in phases, each phase represents the number of updates we want to use the matchings we found in the last updates.
+Let $\varepsilon=\alpha/3$ and $\delta=\varepsilon^{2}$.
+Let $H=M_{1}\cup...\cup M_{t}$.
+Consider the following process for constructing a graph $H_{3}$, that we eventually show is an appropriate ORS graph.
+  • Step 1. Let us define $x_{e}$ for any edge $e=(u,v)$ appearing in matching $M_{i}$ as follows
+    $x_{e}:=\sum_{j>i:u,v\in V^{2}(M_{j})}\frac{1}{d_{j}}.$
+    Define $B_{1}=\{e\mid x_{e}>1/\varepsilon\}$.
+    Let $H_{1}\subseteq H$ include a matching $M_{i}\in H$ iff no more than $2\varepsilon r$ edges of $M_{i}$ belong to $B_{1}$.
+For each matching $M_{i}\in H_{1}$, we then set $M_{i}\leftarrow M_{i}-B_{1}$.
+Observe that $|M_{i}|\geq(1-2\varepsilon)r$ for any $M_{i}\in H_{1}$.
+  • Step 2.
+    Let $H_{2}\subseteq H_{1}$ include every matching $M_{i}\in H_{1}$ independently with probability $\frac{\delta}{2d_{i}}$.
+  • Step 3. For any $i$, call an edge $e\in M_{i}\in H_{2}$ bad if
+    $e\in\bigcup_{j>i:M_{j}\in H_{2}}V^{2}(M_{j}).$
+    Let $B_{2}$ be the set of bad edges.
+Let $H_{3}\subseteq H_{2}$ include matching $M_{i}\in H_{2}$ iff there are at least $(1-\varepsilon)|M_{i}|$ edges in $M_{i}$ that do not belong to $B_{2}$.
+We then set $M_{i}\leftarrow M_{i}-B_{2}$ for any $M_{i}\in H_{3}$.
+Observe that $|M_{i}|\geq(1-\varepsilon)(1-2\varepsilon)r\geq(1-3\varepsilon)r$ for any $M_{i}\in H_{3}$.
+From the construction above, it can be verified that $H_{3}$ is an ORS graph on matchings of size at least $(1-3\varepsilon)r$.
+What remains, is to lower bound the number of matchings that survive in $H_{3}$.
+This is what the rest of the proof focuses on.
+First, we prove that $|B_{1}|\leq\varepsilon|E|$. Suppose for the sake of contradiction that $|B_{1}|>\varepsilon|E|$; then
+$\sum_{e\in B_{1}}x_{e}>|B_{1}|/\varepsilon>|E|.$
+On the other hand, we have
+$\sum_{e\in B_{1}}x_{e}\leq\sum_{e\in E}x_{e}=\sum_{i\in[t]}\frac{|E[G_{i}]\cap V^{2}(M_{i})|}{d}=\sum_{i\in[t]}\frac{d_{i}|M_{i}|}{d}\leq\sum_{i\in[t]}|M_{i}|=|E|,$
+which contradicts the inequality above.
+Here, the first equality follows by double counting: say matching $M_{i}$ “contributes” a value of $1/d$ to edge $e$ if $e\in E[G_{i}]\cap V^{2}(M_{i})$; then $x_{e}$ can be verified to be the total contribution of all matchings to $e$.
+The second equality follows from the definition of $d_{i}$.
+The inequality after that follows from $d_{i}\leq d$, which is by the assumption of the lemma.
+Now that we know $|B_{1}|\leq\varepsilon|E|$, we get that
+$|H_{1}|\geq t/2.$ (3)
+Otherwise, more than $t/2$ matchings have more than $2\varepsilon r$ edges in $B_{1}$, meaning that $|B_{1}|>\frac{t}{2}\cdot 2\varepsilon r=\varepsilon rt=\varepsilon|E|$, a contradiction.
+Let us now fix a matching $M_{i}\in H_{1}$ and fix an edge $e=(u,v)\in M_{i}$. We have
+$\Pr[e\in B_{2}\mid M_{i}\in H_{2}]\leq\sum_{j>i:u,v\in V^{2}(M_{j})}\frac{\delta}{2d_{j}}=\delta x_{e}/2\leq\frac{\delta}{2\varepsilon}=\varepsilon/2,$
+where the last inequality follows because $x_{e}\leq 1/\varepsilon$ for every edge in $M_{i}\in H_{1}$. By linearity of expectation, we have
+$\operatorname{\textbf{E}}[|M_{i}\cap B_{2}|\mid M_{i}\in H_{2}]=\sum_{e\in M_{i}}\Pr[e\in B_{2}\mid M_{i}\in H_{2}]\leq\frac{\varepsilon}{2}|M_{i}|.$
+Applying Markov’s inequality, we thus have
+$\Pr\big{[}|M_{i}\cap B_{2}|\geq\varepsilon|M_{i}|\,\,\,\big{|}\,\,\,M_{i}\in H_{2}\big{]}\leq 1/2.$
+Therefore,
+$\Pr[M_{i}\not\in H_{3}\mid M_{i}\in H_{2}]\leq 1/2.$
+This implies that
 […]
-First, we prove that this algorithm obtains a $\varepsilon$-approximation matching.
-Lemma 10.
-Given a contracted graph $G^{\prime}$ with $n^{\prime}$ vertices that the maximum matching size in $G^{\prime}$ is $O(n^{\prime})$ in $t$ updates, Section 4.2 maintains an $(1-\varepsilon$)-approximate maximum matching for $G^{\prime}$ in the dynamic model.
-Proof.
-Note that we proved in Observation 9 that if the maximal matching size is of order $n^{\prime}$, then we do not need to output a matching in $\varepsilon^{2}n^{\prime}$ updates.
-By Section 4.2, set $S_{1}$ contains edge insertions, and set $S_{2}$ contains all the matchings we found earlier in the algorithm.
-Using Proposition 3, we are given constant subsets of vertices that contain matchings of size at least $\varepsilon n$, we can run Section 4.1 on these subsets to output a $\varepsilon$-approximate matching in given subset $U$, if there are no large enough activated matchings in set $S_{2}\cup S_{1}$.
-Given that we aim to find edge-disjoint matchings, we need to find a maximal matching in $S_{2}\cup S_{1}$ and remove its endpoints from $U$.
-After finding a matching $M^{\prime}$ in $S_{2}\cup S_{1}$ of size $\varepsilon\mu$, we output this matching; otherwise, we deactivate all the endpoints in $M^{\prime}$.
-Note that we are certificated by Proposition 3 to have a matching of size $\mu$ in $U$.
-Since for maximal matching $M^{\prime}$ in $S_{1}\cup S_{2}$ we have $|M^{\prime}|<\varepsilon\mu$, then the maximum matching size in $S_{1}\cup S_{2}$ is at most of size $2\varepsilon\mu$.
-This certifies the existence of a matching of size $(1-2\varepsilon)\mu$ in the rest of $U$.
-Now, we run Section 4.1 on graph $U$ with endpoints of $M^{\prime}$ removed with parameter $\varepsilon/(1-2\varepsilon)$.
-This implies that we get a matching of size $\varepsilon\mu$.
-Since we deactivate vertices of $S_{2}$ we only add edge-disjoint matchings to set $S_{2}$.
-Note that any vertex in $G^{\prime}$ represents a bucket of vertices of $G$.
-By maintaining the number of vertices between two buckets in $G$, we can observe whether there is an edge between two vertices in $G^{\prime}$.
-This is how we can certify if a matching $M^{\prime}$ in $S_{2}$ is in $G^{\prime}$.
-We proved that Section 4.2 outputs a $\varepsilon$-approximate matching $M_{U}$ in each subset of $U$ presented by Proposition 3.
-Note that given $M_{U}$ for each subset $U$, Proposition 3 finds a $(1-\varepsilon)$-approximation matching in $G^{\prime}$.
-This is true by the assumption that the maximum matching size of $G^{\prime}$ is $O(n^{\prime})$.
-∎
-Now that we introduced an algorithm for each contraction, we explore how to build and maintain each contraction.
-To partition updates to the contraction graphs that we maintain we need to run Proposition 4 to approximate the matching size, and then use Proposition 6 to figure out the contracted graph preserving the maximum matching, which we will do after each $\varepsilon^{2}n^{\prime}$ updates, given that the last contracted graph had $n^{\prime}$ vertices.
-We introduce Section 4.2 that given updates calculates the matching size and finds a suitable contraction for the graph at each state.
-For each update, Section 4.2 calls Section 4.2 on all the contractions, however, it only outputs the matching build by the matching preserving contraction.
-We will run $O(\log^{2}n)$ parallel versions of Section 4.2 each for a maintained contracted graphs that we maintain. More formally we introduce Section 4.2:
-[…]
-Now, we prove that this algorithm outputs a $(1-\varepsilon)$-approximation matching.
-Lemma 11.
-Given graph $G$ with $n$ vertices, Section 4.2 maintains an $(1-\varepsilon$)-approximate maximum matching for $G$ in the dynamic model.
-Proof.
-In this algorithm, we maintain $O(\log^{2}n)$ contraction of graph $G$, where $G$ is the updated graph in the current batch of $t$ updates.
-If the maximum matching size of $G$ is $\mu$, by Proposition 6, there exists a contraction $c$ with $n_{c}$ vertices that preserves the maximum matching with $1-\delta$ factors and $\mu=O(n_{c})$.
-Denote $\delta$ a parameter sufficiently smaller than $\varepsilon$.
-We call $c$ a preserving contraction of $G$.
-This means that by Lemma 10, Section 4.2 outputs a $(1-\delta)$-approximation matching for $G^{\prime}$, that is $(1-\varepsilon)$-approximation matching for $G$.
-Note that, by running Proposition 4 on the maintained reduction, we find the preserving contraction of $G$.This concludes the $(1-\varepsilon)$-approximate.
-∎
-4.3 Running Time Analysis
-Now, we explore the time per update for Section 4.2.
-Lemma 12.
-If we set the threshold of Section 4.2 equal to $t=n^{1+\varepsilon/2}\sqrt{n\log n\cdot\mathsf{ORS}_{n}(\varepsilon\mu)}$, then the time spent on single contraction $G^{\prime}$ in both Section 4.2 is $O(\sqrt{n^{1+\varepsilon}\log n\cdot\mathsf{ORS}_{n}(\varepsilon\mu)})$ per each update, where $\mu$ is the maximum matching size in $G^{\prime}$.
-Proof.
-Fix a contracted graph $G^{\prime}=(V^{\prime},E^{\prime})$ with $n$ vertices. We explore the algorithm running time for this graph.
-First, let us determine the time the algorithm spends finding the size of the certificate matching, $\mu$, after $\varepsilon^{2}n$ updates.
-Note that we run Proposition 4 after each $\varepsilon^{2}n$ vertex update, this algorithm outputs $\mu$, which is a constant approximation of the maximum matching size of the graph.
-Moreover, this algorithm certifies the existence of a matching of size $\mu$ in the graph, the running time for which is $\widetilde{O}(n\operatorname{poly}(1/\varepsilon))$.
-By Proposition 3, after each $\varepsilon^{2}n$ update we spend $\widetilde{O}_{\varepsilon}(n)$ on finding the appropriate subsets of vertices such as $U$, including matching of size at least $\varepsilon n$.
-Now, we calculate the time Section 4.2 takes going through the set $S_{1}$ and $S_{2}$, after $\varepsilon^{2}n$ updates to find a matching $M^{\prime}$.
-Note that, after $\varepsilon^{2}n$ updates we go over the set $S_{1}\cup S_{2}$ to find an active matching.
-We add an active edge to the matching if none of the endpoints was added before it.
-The algorithm outputs a maximal matching this way.
-The total running time per $\varepsilon^{2}n$ updates is $|S_{1}|+|S_{2}|$.
-Finally, we explore the running time of Section 4.1.
-Let us assume that at the end of a phase of Section 4.2 we have $S_{2}=\{M_{1}\cup\dots\cup M_{l}\}$, We prove in Lemma 10 that these matchings are edge-disjoint.
-Note that the total edge-disjoint matchings in $G^{\prime}$ can be more than $l$ matching, $M_{1},\dots,M_{l}$ are all the matchings we find during the $t$ edge updates.
-We assume that the total edge-disjoint matchings in $G^{\prime}$ is $l_{t}$.
-We label these matchings as $M_{1},\dots,M_{l},M_{l+1},\dots,M_{l_{t}}$.
-Let us assume that for $i\in\{1,\dots,l\}$ we denote $d_{i}$ the average degree of matching $M_{i}$ in the edge set of $\{M_{i}\cup\dots\cup M_{l_{t}}\}$.
-Denoting by $M^{\prime}$ a maximal matching in $S_{1}\cup S_{2}$ found in Section 4.2, we prove that if Section 4.1 in graph $G^{\prime}\setminus V(M^{\prime})$, takes $O(n^{2+\varepsilon}/d)$ to output matching $M_{i}$ then we have $O(d)=d_{i}$.
-Note that by Lemma 8 we know that the average degree of matching $M_{i}$ in $G^{\prime}\setminus(S_{1}\cup S_{2})$ is $O(d)$.
-Note that, for matching $M_{1},\dots,M_{i-1}$, either their edges are removed from the graph or we deactivate the vertices that made a large matching.
-This implies that no edges from $M_{1},\dots,M_{i-1}$ remain in the graph.
-If we assumed that an edge $e$ was still in the graph with both endpoints not deactivated, we could add this edge to $M^{\prime}$, this is in contradiction with the fact that $M^{\prime}$ is a maximal matching.
-This implies that $d_{i}=O(d)$.
-This results in the conclusion that the running time of Section 4.1 is less than $n^{2+\varepsilon}/d_{i}$.
-This definition of matching average degree motivates us to define ORS graphs.
-Recall that the ORS graph is a set of matching $M_{1},\dots M_{l}$ that each matching is induced in the set of matching with higher indices.
-We then upper bound the running time of our algorithm by the size of ORS graphs according to Lemma 13.
-Now, let us compute the total running time per update.
-Note that since we add a matching of size $\varepsilon\mu$ to $S_{2}$ after $\varepsilon^{2}n$ updates, $|S_{2}|<|S_{1}|$.
-Let us assume that we add $M_{1},\dots M_{l}$ to $S_{2}$ by running Section 4.1.
-The total running time per update is:
-$\frac{\frac{|S_{1}|}{\varepsilon^{2}n}\cdot(|S_{1}|+|S_{2}|+\widetilde{O}(n))+n^{2+\varepsilon}\cdot\sum_{i=1}^{l}1/d_{i}}{t}.$
-Using Lemma 13 we have the running time as:
-$\frac{\frac{|S_{1}|}{\varepsilon^{2}n}\cdot(|S_{1}|+|S_{2}|+\widetilde{O}(n))+n^{2+\varepsilon}\cdot O(\mathsf{ORS}_{n}(\varepsilon\mu))\log n}{t}.$
-Now, if we set the threshold as $t=n^{1+\varepsilon/2}\sqrt{n\log n\cdot\mathsf{ORS}_{n}(\varepsilon\mu)}$ we have:
-$|S_{2}|,|S_{1}|<t=n^{1+\varepsilon/2}\sqrt{n\log n\cdot\mathsf{ORS}_{n}(\varepsilon\mu)},$
-This is because we could add a matching of size $\varepsilon\mu$ to $S_{2}$ after $\varepsilon^{2}n$ updates, and the total insertions are less than the total updates.
-The total running time per update is:
-$O(\sqrt{n^{1+\varepsilon}\log n\cdot\mathsf{ORS}_{n}(\varepsilon\mu)}).\qed$
-Lemma 13.
-Given a set of matching $M_{1},\dots,M_{t}$ with distinct edges, of size at least $r$ in graph $G$, for any index $i\in[1,t]$ let us define $d_{i}$ as the average degree of $V(M_{i})$ in $M_{1}\cup\dots\cup M_{i}$.
-For any $\alpha\in(0,1)$, we have $\sum_{i=1}^{t}1/d_{i}=O(\mathsf{ORS}_{n}(r(1-\alpha))\cdot\log n)$.
-Proof.
-Note that the maximum average degree of a matching $M$ is $n-1$, and the minimum is 1.
-Therefore, if we partition the input matchings into $\log n=k$ groups, the average degree of any matching in each partition is at most within 2 factors of the minimum average degree.
-We first prove that if we have $t^{\prime}$ matchings $M^{\prime}_{1},\dots,M^{\prime}_{t^{\prime}}$ a subset of $\{M_{1},\dots,M_{t}\}$, and their average degree as $d^{\prime}_{1}\star,\dots,d^{\prime}_{t^{\prime}}$ that for any $i\in[1,t^{\prime}]$, $1/d^{\prime}_{i}$ is in a $2$ factor of $\min_{i=1}^{t^{\prime}}1/d^{\prime}_{i}$.
-For any $\alpha\in(0,1)$, we have $\sum_{i=1}^{t^{\prime}}1/d^{\prime}_{i}=O(\mathsf{ORS}_{n}(r(1-\alpha)))$.
-To prove this, we first sample some of the matchings and the edges that are not induced in proper subsets of matchings.
-We prove there are plenty number of matchings remaining that are induced within their corresponding set of matchings.
-Let us fix a small variable $\delta<\varepsilon,\delta,\varepsilon\in(0,1)$. We set $H=M^{\prime}_{1}\cup...\cup M^{\prime}_{t^{\prime}}$.
-  • Step 1. Let us define $x_{e}$ for any edge $e=(u,v)$ appearing in matching $M^{\prime}_{\alpha_{j}}$ as follows
-    $\sum_{i\in\{j,j+1,\dots,t^{\prime}\};u,v\in V^{2}(M_{\alpha_{i}})}\frac{1}{d^{\prime}_{\alpha_{i}}}.$
-    We define $E_{1}=\{e|x_{e}>1/\varepsilon\}$ We remove any matching from $H$ with more than $2\varepsilon r$ edges in $E_{1}$ Let us call the remaining set of matchings $H_{1}$.
-  • Step 2.
-Let us pick a matching $M^{\prime}_{i}\in H$ with probability $\frac{\delta}{2d^{\prime}_{i}}$.
-Let us assume that at this step, we pick the following $k$ matchings $M^{\prime}_{\alpha_{1}},\dots,M^{\prime}_{\alpha_{k}}$.
-we define this subset of the graph as $H_{2}$.
-  • Step 3.
-Now, we prone any matching $M^{\prime}_{i}$ chosen in $H_{1}$ and has less than $(1-\delta)cn$ edges induced in $(M^{\prime}_{1}\cup\dots\cup M^{\prime}_{i})\cap(H_{2})$.
-Let us call the remaining graph $H_{3}$.
-To see that the claim is true, first, we need to prove that in Step 1 we remove at most $\varepsilon$ fraction of the edges in $H$. Second, we prove that $H_{2}$ is large enough.
-Let us assume that after Step 1 we remove more than $\varepsilon rt$ of the edges in $H$ from the total $rt^{\prime}$ edges. Therefore, we have
-$\sum_{e\in E_{1}}x_{e}>rt^{\prime}.$
-To compute $\sum_{e\in H}x_{e}$, we notice how much a matching $M^{\prime}_{\alpha_{i}}$ contributes to the total sum.
-Note that since the average degree of $M^{\prime}_{\alpha_{i}}$ in $\{M^{\prime}_{\alpha_{1}},\dots,M^{\prime}_{\alpha_{i}}\}$ is at most $d^{\prime}_{\alpha_{i}}$ therefore, the total contribution of matching $M^{\prime}_{\alpha_{i}}$ is $r\cdot d^{\prime}_{\alpha_{i}}\cdot\frac{1}{d^{\prime}_{\alpha_{i}}}$.
-In conclusion, the total sum of $x_{e}$ for any edge $e\in H$ is at most
-$\sum_{i\in\{1,\dots,t^{\prime}\}}d^{\prime}_{\alpha_{i}}\cdot r/d^{\prime}_{\alpha_{i}}=rt^{\prime},$
-which is a contradiction suggesting that we remove at most $\varepsilon rt^{\prime}$ edges.
-Now that we only removed at most $\varepsilon$ fraction of the edges, note that $H_{1}$ contains at least half of the matchings in $H$.
-This is because if more than $|H|/2$ matchings had more than $2\varepsilon r$ edges removed, then the total number of edges in $E_{1}$ would be more than $\varepsilon rt^{\prime}$ which is a contradiction.
-We aim to prove that $H_{3}$ contains $\delta|H_{1}|/2$ matchings.
-Let us fix a matching $M\in H_{1}$.
-Note that for any remaining edge $e$ in $M$, we have $x_{e}<1/\varepsilon$.
-Therefore, the probability that $e$ belongs to another matching is less than $\delta/2\varepsilon$ since we picked each matching $M^{\prime}_{i}$ with probability $\delta/2d^{\prime}_{i}$.
-This implies that the expected number of edges in $E(^{\prime})\cap E(H_{3})$ is more than $(1-2\varepsilon)(1-\delta/2\varepsilon)r$.
-Now, by applying Markov’s inequality to the random variable $r(1-2\varepsilon)-|E(M)\cap E(H_{3})|$, here we assume that the maximum number of edges is $r(1-2\varepsilon)$, and we get the following.
-$\Pr[r(1-2\varepsilon)-|E(M)\cap E(H_{3})|\geq(1-2\varepsilon)(\delta/\varepsilon)r|M\in H_{2}]\leq\frac{\operatorname{\textbf{E}}[r(1-2\varepsilon)-|E(M)\cap E(H_{3})|M\in H_{2}]}{(1-2\varepsilon)(\delta/\varepsilon)r}.$
-Since $\operatorname{\textbf{E}}[r(1-2\varepsilon)-|E(M)\cap E(H_{3})||M\in H_{2}]\leq((1-2\varepsilon)\delta/2\varepsilon)r$ we have
-$\Pr[r(1-2\varepsilon)-|E(M)\cap E(H_{3})|\geq(1-2\varepsilon)(\delta/\varepsilon)r|M\in H_{2}]\leq\frac{(\delta/2\varepsilon)r}{(\delta/\varepsilon)r}=1/2.$
-Therefore, we have
-$\Pr[r(1-2\varepsilon)-|E(M)\cap E(H_{3})|\leq(1-2\varepsilon)(\delta/\varepsilon)r|M\in H_{2}]>1/2,$
-which means
-$\Pr[|E(M)\cap E(H_{3})|\geq(1-2\varepsilon)(1-\delta/\varepsilon)r|M\in H_{2}]>1/2.$
-Multiplying with the probability that $M\in H_{2}$, we get that the probability of matching to have more than $(1-2\varepsilon)(1-\delta/\varepsilon)r$ edges and be selected is more than $\delta/(4\varepsilon d^{\prime}_{i})$.
-Now, summing over all matchings in $H_{1}$ we get
-$\operatorname{\textbf{E}}[|H_{3}|]\geq(\delta/(4\varepsilon))\sum_{i,M^{\prime}_{i}\in H_{1}}1/d^{\prime}_{i}.$
-Since $H_{3}$ is an ORS graph we get
-$(4\varepsilon)\mathsf{ORS}_{n}((1-2\varepsilon)(1-\delta/\varepsilon)r)/\delta\geq\sum_{i,M^{\prime}_{i}\in H_{1}}1/d^{\prime}_{i}.$
-Now, using that for any $j\in\{1,..,t^{\prime}\}$ we have $1/d^{\prime}_{j}\leq 2\min_{i=1}^{t^{\prime}}1/d^{\prime}_{i}$, for any $i,j\in\{1,...,t^{\prime}\}$ we get $1/d^{\prime}_{i}\leq 2/d^{\prime}_{j}$.Therefore,
-$\sum_{i,M^{\prime}_{i}\in H_{1}}1/d^{\prime}_{i}\geq\left(\sum_{j,M^{\prime}_{j}\in H\setminus H_{1}}1/d^{\prime}_{j}\right)/2.$
-Now, we give an upper bound for $\sum_{i,M_{i}\in H}1/d^{\prime}_{i}=\sum_{i,M^{\prime}_{i}\in H_{1}}1/d^{\prime}_{i}+\sum_{j,M^{\prime}_{j}\in H\setminus H_{1}}1/d^{\prime}_{j}$.
-We have the following.
-$(12\varepsilon)\mathsf{ORS}_{n}((1-2\varepsilon)(1-\delta/\varepsilon)r)/\delta\geq 3\sum_{i,M^{\prime}_{i}\in H_{1}}1/d^{\prime}_{i}\geq\sum_{i,M^{\prime}_{i}\in H_{1}}1/d^{\prime}_{i}+\sum_{j,M^{\prime}_{j}\in H\setminus H_{1}}1/d^{\prime}_{j}.$
-Setting $\delta=\varepsilon/2,\varepsilon=4\alpha/13$ we have:
-$\sum_{i,M^{\prime}_{i}\in H}1/d^{\prime}_{i}=O(\mathsf{ORS}_{n}((1-\alpha)r)).$
-Note that we showed an upper bound for the sum of the inverse average degree of each matching in one partition. Summing up the inverse average degree for all the partitions we get:
-$\sum_{i=1}^{t}1/d_{i}=O(\mathsf{ORS}_{n}((1-\alpha)r)\cdot\log n.\qed$
-Lemma 14.
-If we set the threshold of Section 4.2 equal to $t=n^{1+\varepsilon/2}\sqrt{n\log n\cdot\mathsf{ORS}_{n}(\Theta(\varepsilon^{2}n)}$, then the total time per update is $O(\operatorname{poly}\log n)\cdot\sqrt{n^{1+\varepsilon}\log n\cdot\mathsf{ORS}_{n}(\Theta(\varepsilon^{2}n))}$.
-Proof.
-We run the $O(\log^{2}n)$ parallel version of Section 4.2, and maintaining the contracted graphs takes $\operatorname{poly}\log n$ time.
-We will maintain the number of parallel edges in $G^{\prime}$ between any two contracted vertex in $\operatorname{poly}\log n$ time.
-Denote the set of contractions as $\mathcal{C}$ in one phase of the algorithm.
-Note that if the size of a contraction $c$ is $n_{c}$, and the maximum matching in $G$ is of order $n_{c}$ we run the algorithm for this contraction.
-Using Proposition 6 there is a reduction that preserves that matching in $G$ with a factor of $1-\varepsilon$.
-Now, integrating over the contractions in $\mathcal{C}$ we find the corresponding contraction by running Proposition 4.
-If the maximum matching in contraction $c$ is $M_{c}$, this algorithm certifies the existence of a matching of size $|M_{c}|/2<\mu\leq|M_{c}|$ in at least one contraction with $n_{c}$ vertices.
-On the other hand, Proposition 3 outputs subsets with a matching of size at least $\varepsilon n_{c}$.
-This means that the $\varepsilon$-approximation matching found on each subset determined by Proposition 3 is at least of size $\varepsilon^{2}n$.
-Now, since we run Section 4.2 on all contractions, nevertheless they contain the maximum matching or not, assuming $n_{c}$ is the number of vertices in $c$, we have the total running time as:
-$\sum_{c\in\mathcal{C}}\sqrt{n_{c}^{1+\varepsilon}\log n_{c}\cdot\mathsf{ORS}_{n_{c}}(\Theta(\varepsilon^{2}n_{c}))}\leq O(\operatorname{poly}\log n)\cdot\sqrt{n^{1+\varepsilon}\log n\cdot\mathsf{ORS}_{n}(\Theta(\varepsilon^{2}n))}.\qed$
-Now, using Lemma 11 and Lemma 14 the proof of Theorem 1 is complete.
+Note that for each matching $M_{i}\in H_{3}$, we have $|M_{i}|\geq(1-\varepsilon)(1-2\varepsilon)r\geq(1-3\varepsilon)r$ by construction.
+Additionally, $M_{i}$ is an induced matching in $\bigcup_{j<i;M_{j}\in H_{3}}M_{j}$, again, by construction.
+Therefore, $H_{3}$ is an ORS graph with matchings of size $(1-3\varepsilon)r$.
+This implies that
+$\mathsf{ORS}_{n}((1-3\varepsilon)r)\geq E[|H_{3}|]\geq\Omega(\varepsilon^{2}t/d).$
+Moving the terms and recalling that $\varepsilon=\alpha/3$, we get that
+$t=O\left(\frac{d}{\alpha^{2}}\cdot\mathsf{ORS}_{n}((1-\alpha)r)\right).\qed$
+Now, we are ready to prove Lemma 11.
+Proof of Lemma 11.
+Given $M_{1},\dots M_{t}$, we partition the matchings to $\log n$ groups $P_{1},\dots,P_{\log n}$ where $M_{j}$ is in the $i$-th group if $2^{i-1}\leq d_{j}<2^{i}$.
+Define $d^{\prime}_{j}:=|E[P_{i}]\cap V^{2}(M_{j})|/|M_{j}|$.
+Note that by definition for any matching $M_{j}\in P_{i}$ we have $d^{\prime}_{j}\leq d_{j}\leq 2^{i}$.
+Therefore, by Claim 12 we have
+$\sum_{i\in[\log n]}\sum_{j:M_{j}\in P_{i}}1/d_{j}$ $\leq\sum_{i\in[\log n]}|P_{i}|/2^{i-1}$
+$\leq O\left(\frac{1}{\alpha^{2}}\mathsf{ORS}_{n}((1-\alpha)r)\right)\cdot\sum_{i\in[\log n]}2^{i}/2^{i-1}$ (Applying Claim 12 with $d=2^{i}$, we get $|P_{i}|\leq O\left(\frac{1}{\alpha^{2}}\mathsf{ORS}_{n}((1-\alpha)r)2^{i}\right)$ which is plugged here.)
+$=O\left(\frac{1}{\alpha^{2}}\mathsf{ORS}_{n}((1-\alpha)r)\cdot\log n\right).$
+The proof is thus complete. ∎
 5 Bounding Density of (Ordered) Ruzsa-Szemerédi Graphs
@@
 Let us start with an observation that is extremely helpful in proving Lemma 17.
-Observation 16.
+Observation 14.
 Take a matching $M^{\prime}\in\mathcal{M^{\prime}}$ and take any vertex $v$ matched by $M^{\prime}$.
@@
 Let us define $Y_{M}:=V\setminus\cup_{v_{i}\in S}N_{i}$.
-From Observation 16, we get that all edges of $M$ must have both endpoints still in $Y_{M}$.
+From Observation 18, we get that all edges of $M$ must have both endpoints still in $Y_{M}$.
 Suppose for now that $M\in\mathcal{M}^{\prime}$ is chosen uniformly at random.
@@
 $\leq n-\sum_{i=1}^{k}\frac{4\delta c}{t}\cdot|N_{i}|^{2}.$ (Since $|N_{i}|\leq t/2$ as $\mathcal{M}$ is the union of $t/2$ matchings.)
-$\leq n-\frac{4\delta c}{t}\cdot(tcn/2-n^{2}/x)$ (By 2.)
+$\leq n-\frac{4\delta c}{t}\cdot(tcn/2-n^{2}/x)$ (By 12.)
 $=n-2\delta c^{2}n+\frac{4\delta cn^{2}}{tx}$
@@
 Instead of the expected value of $Y_{M}$, we need some (weak) concentration.
-Take $\varepsilon=\delta c^{2}/2$ and note from 3 that $(1+\varepsilon)\operatorname{\textbf{E}}[|Y_{M}|]<(1+\varepsilon)(1-\delta c^{2})n<(1-\delta c^{2}/2)n$.
+Take $\varepsilon=\delta c^{2}/2$ and note from 13 that $(1+\varepsilon)\operatorname{\textbf{E}}[|Y_{M}|]<(1+\varepsilon)(1-\delta c^{2})n<(1-\delta c^{2}/2)n$.
 We have
@@
 $\geq\varepsilon/2=\delta c^{2}/4.$ (7)
-We call a matching $M\in\mathcal{M}^{\prime}$ a good matching if $|Y_{M}|<(1-\delta c^{2}/2)n$. From 4 we get that at least $(\delta c^{2}/4)|\mathcal{M}^{\prime}|=\delta c^{2}t/8$ matchings are good.
+We call a matching $M\in\mathcal{M}^{\prime}$ a good matching if $|Y_{M}|<(1-\delta c^{2}/2)n$.
+From 14 we get that at least $(\delta c^{2}/4)|\mathcal{M}^{\prime}|=\delta c^{2}t/8$ matchings are good.
 Step 3: Grouping good matchings.
@@
 Note also that $Y=Y_{M^{\prime}}$ for any other matching $M^{\prime}\in C$ as well since they all match the same subset of pivots.
-Moreover, we have $|Y|<(1-\delta c^{2}/2)n$ since $M$ is good, and all matchings in $C$ only match vertices in $Y$ as discussed earlier (as a consequence of Observation 16).
+Moreover, we have $|Y|<(1-\delta c^{2}/2)n$ since $M$ is good, and all matchings in $C$ only match vertices in $Y$ as discussed earlier (as a consequence of Observation 18).
 This implies that the graph $H$ on vertex set $Y$ and edge-set $C$ (i.e., including all edges of all matchings in $C$) is an $\mathsf{ORS}_{n^{\prime}}(cn,t^{\prime})$ graph for $n^{\prime}<(1-\delta c^{2}/2)n$ and $t^{\prime}\geq\delta c^{2}t/(8\cdot 2^{x})$ as desired.
@@
 $c^{\prime}n^{\prime}$ $=c^{\prime\prime}n^{\prime\prime}$
-$\geq\frac{c^{\prime\prime}n^{\prime}}{(1-\delta c^{\prime\prime 2}/2)}$ (By 6.)
+$\geq\frac{c^{\prime\prime}n^{\prime}}{(1-\delta c^{\prime\prime 2}/2)}$ (By 16.)
 $\geq(1+\delta c^{\prime\prime 2}/2)c^{\prime\prime}n^{\prime}$
-$\geq\Big{(}1+\delta\Big{(}(1-\delta c^{2}/10)c\Big{)}^{2}/2\Big{)}\Big{(}(1-\delta c^{2}/10)c\Big{)}n^{\prime},$ (Since the RHS is minimized when $c^{\prime\prime}$ is minimized, thus we can replace $c^{\prime\prime}$ with its LB from 5.)
+$\geq\Big{(}1+\delta\Big{(}(1-\delta c^{2}/10)c\Big{)}^{2}/2\Big{)}\Big{(}(1-\delta c^{2}/10)c\Big{)}n^{\prime},$ (Since the RHS is minimized when $c^{\prime\prime}$ is minimized, thus we can replace $c^{\prime\prime}$ with its LB from 15.)
 $\geq\Big{(}1+0.4\delta c^{2}\Big{)}\Big{(}(1-\delta c^{2}/10)c\Big{)}n^{\prime}$ (Since $\delta\big{(}(1-\delta c^{2}/10)c\big{)}^{2}/2\geq\delta(0.9c)^{2}/2>0.4\delta c^{2}$)
@@
 $dist(v_{i},v_{j})=|V_{i}|+|V_{j}|-2|V_{i}\cap V_{j}|\geq 2r.$
-Note that in 7, $a_{i}b_{i}$ maximizes when we have $a_{i}=b_{i}=(t+1)/2$ for odd $t$ and $a_{i}=(t+2)/2,b_{i}=t/2$ for even $t$. This concludes that $t<1/\varepsilon+1$
+Note that in 10, $a_{i}b_{i}$ maximizes when we have $a_{i}=b_{i}=(t+1)/2$ for odd $t$ and $a_{i}=(t+2)/2,b_{i}=t/2$ for even $t$. This concludes that $t<1/\varepsilon+1$
+6 Bounding Density of (Ordered) Ruzsa-Szemerédi Graphs
+In this section we prove upper bounds on the density of ORS graphs.
+Our main result of this section is a proof of Theorem 2 which we present in Section 6.1.
+We then show that a much stronger upper bound can be proved if matchings cover more than half of vertices in Section 6.2.
+Before presenting our proof, let us first briefly discuss how the triangle removal lemma is useful for upper bounding density of RS graphs and why it does not seem to help for upper bounding ORS.
+The triangle removal lemma states that so long as the number of triangles in a graph are $o(n^{3})$, then we can make the graph triangle free by removing $o(n^{2})$ of its edges.
+Suppose we have a bipartite RS graph with induced matchings $M_{1},\ldots,M_{k}$.
+Add $k$ vertices $v_{1},\ldots,v_{k}$ to this graph and connect each $v_{i}$ to all the vertices matched by $M_{i}$.
+Now, crucially, because each $M_{i}$ is an induced matching, each $v_{i}$ is part of exactly $|M_{i}|$ triangles, one for each edge of $M_{i}$.
+As such, the total number of triangles can be upper bounded by $O(kn)=O(n^{2})$ which is well within the regime that one can apply triangle removal lemma.
+However, because the matchings in ORS are not induced matchings of the whole graph, the number of triangles cannot simply be upper bounded by $O(n^{2})$ after adding the auxiliary vertices.
+In fact, our upper bound completely deviates from this approach and does not rely on the triangle removal lemma.
+6.1 Linear Matchings
+The following lemma, which is one of our main tools in proving Theorem 2, shows that so long as vertex degrees are larger than a threshold, we can reduce the number of vertices rather significantly without hurting the size of a relatively large number of induced matchings.
+Our proof of Lemma 17 builds on the techniques developed for upper bounding RS graphs with matchings of size very close to $n/4$ in [29] that we extend to ORS graphs with balanced degrees.
+Lemma 17.
+Let $G$ be an $\mathsf{ORS}_{n}(r,t)$ graph with $r=cn$ with even $t$.
+Let $M_{1},\ldots,M_{t}$ be the $t$ matchings in $G$ as defined in Definition 2.
+Let us partition these matchings into two subsets
+=$\mathcal{M}=\{M_{1},\ldots,M_{t/2}\}\qquad\text{and}\qquad\mathcal{M}^{\prime}=\{M_{t/2+1},\ldots,M_{t}\}.$
+Suppose that for every vertex $v$, it holds that $\deg_{\mathcal{M}^{\prime}}(v)\geq\delta ct$ for some $\delta>0$.
+Then there exists an $\mathsf{ORS}_{n^{\prime}}(cn,t^{\prime})$ graph $H$ on $n^{\prime}<(1-\delta c^{2}/2)n$ vertices where $t^{\prime}\geq\delta c^{2}t/(8\cdot 2^{x})$ and $x=4n/ct$.
+Let us start with an observation that is extremely helpful in proving Lemma 17.
+Observation 18.
+Take a matching $M^{\prime}\in\mathcal{M^{\prime}}$ and take any vertex $v$ matched by $M^{\prime}$.
+Let $N_{\mathcal{M}}(v)$ be the set of neighbors of $v$ in $\mathcal{M}$.
+That is, $u\in N_{\mathcal{M}}(v)$ iff there is $M\in\mathcal{M}$ such that $(v,u)\in M$.
+Then no vertex in $N_{\mathcal{M}}(v)$ can be matched in $M^{\prime}$.
+Proof.
+Suppose there is $u\in N_{\mathcal{M}}(v)$ that is matched by $M^{\prime}$.
+Let $(u,w)\in M^{\prime}$ be the matching edge involving $u$.
+Also take the edge $(v,y)\in M^{\prime}$ that involves $v$ (which exists by definition of $v$).
+Note that since the matchings $M_{1},\ldots,M_{t}$ are edge-disjoint by Definition 2, $v$ and $u$ cannot be matched together in $M^{\prime}$ (as $v$ and $u$ must be matched in some other matching in $\mathcal{M}$ by definition of $N_{\mathcal{M}}(v)$).
+Now since $(u,w),(v,y)\in M^{\prime}$ but $(v,u)$ belongs to some matching in $\mathcal{M}$ that comes before $M^{\prime}$ in the ordering, matching $M^{\prime}$ cannot be an induced matching among its previous matchings, contradicting Definition 2 for ORS graphs.
+∎
+We are now ready to prove Lemma 17.
+Proof of Lemma 17.
+We explain the proof in a few steps.
+Step 1: The pivots.
+We iteratively take a vertex of highest remaining degree in $\mathcal{M}$, and remove its neighbors in $\mathcal{M}$ from the graph.
+More formally, take the graph $G_{0}=(V_{0},E_{0})$ where $V_{0}=V(G)$ and $E_{0}$ is the union of the matchings in $\mathcal{M}$.
+At step $i$, we choose $v_{i}$ in $V_{i-1}$ with the maximum degree in $G_{i-1}$.
+Let us define $N_{i}$ as the neighbors of $v_{i}$ in $G_{i-1}$.
+We then remove $v_{i}$ and its neighbors in $G_{i-1}$ to define $G_{i}=(V_{i},E_{i})$.
+Namely, $V_{i}=V_{i-1}\setminus(\{v_{i}\}\cup N_{i})$ and $E_{i}$ includes all edges in $E_{i-1}$ except those that have at least one endpoint that does not belong to $V_{i}$.
+Let $k$ be the maximum integer such that $|N_{k}|>n/x$.
+We define $P=\{v_{1},\ldots,v_{k}\}$ to be the set of pivots.
+Note that since $N_{1},\dots,N_{k}$ are disjoint sets, we get that $n\geq\sum_{i=1}^{k}|N_{i}|\geq kn/x$ which implies that
+$|P|\leq x.$ (11)
+Moreover, note that when removing $v_{i}$, we remove $|N_{i}|$ vertices from the graph and each of those vertices has remaining degree at most $|N_{i}|$ (as we choose $v_{i}$ to be the vertex of highest remaining degree).
+In other words, we remove at most $|N_{i}|^{2}$ from the graph after removing $v_{i}$ and its remaining neighbors.
+On the other hand, after removing $v_{1},\ldots,v_{k}$ and their neighbors, the maximum degree in the graph is at most $n/x$ (by definition of $k$), and so there are at most $n^{2}/x$ edges in the graph.
+This implies that:
+$|E_{0}|-\sum_{i=1}^{k}|N_{i}|^{2}\leq n^{2}/x.$
+Given that $E_{0}$ includes $t/2$ edge disjoint matchings of size $cn$, we get that $|E_{0}|\geq tcn/2$. Plugged into the inequality above, this implies that
+$\sum_{i=1}^{k}|N_{i}|^{2}\geq|E_{0}|-n^{2}/x\geq tcn/2-n^{2}/x.$ (12)
+Step 2: Vertex reduction.
+Take a matching $M\in\mathcal{M}^{\prime}$.
+Let $S_{M}$ be the set of pivots matched by $M$, i.e., $S_{M}=P\cap V(M)$.
+Let us define $Y_{M}:=V\setminus\cup_{v_{i}\in S}N_{i}$.
+From Observation 18, we get that all edges of $M$ must have both endpoints still in $Y_{M}$.
+Suppose for now that $M\in\mathcal{M}^{\prime}$ is chosen uniformly at random.
+We first argue that the expected size of $|Y_{M}|$ is relatively small (despite it preserving all edges of $M$ completely).
+We have
+$\operatorname{\textbf{E}}_{M\sim\mathcal{M}^{\prime}}[|Y_{M}|]$ $=n-\sum_{i=1}^{k}\Pr[v_{i}\in V(M)]\cdot|N_{i}|$
+$=n-\sum_{i=1}^{k}\frac{\deg_{\mathcal{M}^{\prime}}(v_{i})}{|\mathcal{M}^{\prime}|}\cdot|N_{i}|$ (As $M$ is chosen uniformly from $\mathcal{M}^{\prime}$ and $v_{i}$ is matched in $\deg_{\mathcal{M}^{\prime}}(v_{i})$ of matchings in $\mathcal{M}^{\prime}$.)
+$=n-\sum_{i=1}^{k}\frac{2\deg_{\mathcal{M}^{\prime}}(v_{i})}{t}\cdot|N_{i}|$ (Since $|\mathcal{M}^{\prime}|=t/2$.)
+$\leq n-\sum_{i=1}^{k}\frac{2\delta ct}{t}\cdot|N_{i}|$ (Since $\deg_{\mathcal{M}^{\prime}}(v_{i})\geq\delta ct$ as assumed in Lemma 17.)
+$\leq n-\sum_{i=1}^{k}\frac{4\delta c}{t}\cdot|N_{i}|^{2}.$ (Since $|N_{i}|\leq t/2$ as $\mathcal{M}$ is the union of $t/2$ matchings.)
+$\leq n-\frac{4\delta c}{t}\cdot(tcn/2-n^{2}/x)$ (By 12.)
+$=n-2\delta c^{2}n+\frac{4\delta cn^{2}}{tx}$
+$\leq n-2\delta c^{2}n+\frac{4\delta cn^{2}}{(4n/cx)x}$ (Since $t=4n/cx$ by definition of $x$ in Lemma 17.)
+$=n-2\delta c^{2}n+\delta c^{2}n$
+$=(1-\delta c^{2})n.$ (13)
+Instead of the expected value of $Y_{M}$, we need some (weak) concentration.
+Take $\varepsilon=\delta c^{2}/2$ and note from 13 that $(1+\varepsilon)\operatorname{\textbf{E}}[|Y_{M}|]<(1+\varepsilon)(1-\delta c^{2})n<(1-\delta c^{2}/2)n$.
+We have
+$\Pr_{M\sim\mathcal{M}^{\prime}}\Big{[}|Y_{M}|<(1+\varepsilon)\operatorname{\textbf{E}}[|Y_{M}|]<(1-\delta c^{2}/2)n\Big{]}$ $=1-\Pr_{M\sim\mathcal{M}^{\prime}}\Big{[}|Y_{M}|>(1+\varepsilon)\operatorname{\textbf{E}}[|Y_{M}|]\Big{]}$
+$\geq 1-\frac{1}{1+\varepsilon}$ (By Markov’s inequality.)
+$\geq\varepsilon/2=\delta c^{2}/4.$ (14)
+We call a matching $M\in\mathcal{M}^{\prime}$ a good matching if $|Y_{M}|<(1-\delta c^{2}/2)n$.
+From 14 we get that at least $(\delta c^{2}/4)|\mathcal{M}^{\prime}|=\delta c^{2}t/8$ matchings are good.
+Step 3: Grouping good matchings.
+Now, we partition the good matchings in $\mathcal{M}^{\prime}$ into $2^{|P|}\leq 2^{x}$ classes depending on which subset of the pivots they match.
+Since there are at least $\delta c^{2}t/8$ good matchings, at least $\frac{\delta c^{2}t}{8\cdot 2^{x}}$ of them must belong to the same class $C$.
+Let $Y=Y_{M}$ for some $M\in C$.
+Note also that $Y=Y_{M^{\prime}}$ for any other matching $M^{\prime}\in C$ as well since they all match the same subset of pivots.
+Moreover, we have $|Y|<(1-\delta c^{2}/2)n$ since $M$ is good, and all matchings in $C$ only match vertices in $Y$ as discussed earlier (as a consequence of Observation 18).
+This implies that the graph $H$ on vertex set $Y$ and edge-set $C$ (i.e., including all edges of all matchings in $C$) is an $\mathsf{ORS}_{n^{\prime}}(cn,t^{\prime})$ graph for $n^{\prime}<(1-\delta c^{2}/2)n$ and $t^{\prime}\geq\delta c^{2}t/(8\cdot 2^{x})$ as desired.
+∎
+The next lemma follows by applying Lemma 17 after carefully pruning vertices of low degree in $\mathcal{M}^{\prime}$.
+Intuitively, Lemma 19 significantly reduces the number of vertices and shows that a relatively large number of matchings will have a lot of edges in the remaining graph.
+Lemma 19.
+Let $G$ be an $\mathsf{ORS}_{n}(cn,t)$ graph for even $t$.
+Then for some $\kappa\geq c^{3}/160$, there exist $\mathsf{ORS}_{n^{\prime}}(c^{\prime}n^{\prime},t^{\prime})$ graphs such that $n^{\prime}\leq n$, $c^{\prime}\geq(1+\kappa)c$ and $t^{\prime}\geq t/2^{O(n/ct)}$.
+Proof.
+Let $\mathcal{M}$ and $\mathcal{M}^{\prime}$ be defined for $G$ as in Lemma 17.
+We would like to apply Lemma 17 iteratively to prove Lemma 19.
+However, Lemma 17 requires a lower bound on the degrees in $\mathcal{M}^{\prime}$ which might not necessarily hold for our graph $G$.
+Therefore to be able to use it, we first have to prune vertices of small degrees.
+Let $G^{\prime}_{0}$ be the subgraph of $G$ only including the edges of the matchings in $\mathcal{M}^{\prime}$.
+Let $\delta=1/4$.
+We iteratively remove vertices of low degree.
+Formally, in step $i$, if there is a vertex $v_{i}$ with degree less than $2\delta ct$ in $G^{\prime}_{i-1}$, we define the graph $G^{\prime}_{i}$ of the next step to be graph obtained after removing $v_{i}$ and its edges from $G^{\prime}_{i-1}$.
+Let $G^{\prime}_{k}$ be the final graph that does not have any vertex of degree smaller than $2\delta ct$.
+Suppose that $k=\alpha n$; that is, we remove $\alpha n$ vertices in the process above. We consider two cases depending on the value of $\alpha$.
+Case 1: $\alpha\geq\delta c^{3}/10$.
+Since we remove at most $\delta ct$ edges in every step, the total number of removed edges is at most $\delta\alpha ctn$ over the $k=\alpha n$ steps of constructing $G^{\prime}_{k}$.
+Say a matching $M\in\mathcal{M}^{\prime}$ is damaged if a total of at least $3\delta\alpha cn$ of its edges have been removed.
+In other words, $M$ is damaged if less than $|M|-3\delta\alpha cn=(1-3\delta\alpha)cn$ of its edges belong to $G^{\prime}_{k}$.
+Since the total number of edges removed is at most $\delta\alpha ctn$ and each damaged matching has at least $3\delta\alpha cn$ of its edges removed, the total number of damaged matchings can be upper bounded by $\delta\alpha ctn/3\delta\alpha cn=t/3$.
+Since $|\mathcal{M}^{\prime}|=t/2$, at least $t/6$ matchings in $\mathcal{M}^{\prime}$ are not damaged, and have size at least $(1-3\delta\alpha)cn$.
+These matchings themselves form an $\mathsf{ORS}_{n^{\prime}}(c^{\prime}n^{\prime},t/6)$ graph with parameters
+$n^{\prime}=(1-\alpha)n\leq(1-\delta c^{3}/10)n\stackrel{{\scriptstyle(\delta=1/4)}}{{=}}(1-c^{3}/40)n,$
+and
+$c^{\prime}n^{\prime}\geq(1-3\delta\alpha)cn=(1-3\delta\alpha)c\frac{n^{\prime}}{1-\alpha}\stackrel{{\scriptstyle(\delta=1/4)}}{{=}}\left(1-\frac{3}{4}\alpha\right)c\frac{n^{\prime}}{1-\alpha}\geq\left(1+\frac{1}{4}\alpha\right)cn^{\prime}\geq(1+c^{3}/160)cn^{\prime},$
+which implies that $c^{\prime}\geq(1+c^{3}/160)c$. Taking $\kappa=c^{3}/160$ proves the lemma in this case.
+Case 2 – $\alpha<\delta c^{3}/10$:
+Let $G_{k}$ be the graph $G$ after removing vertices $v_{1},\ldots,v_{k}$ (the difference between $G_{k}$ and $G^{\prime}_{k}$ is that $G^{\prime}_{k}$ only includes the edges of $\mathcal{M}^{\prime}$ but $G_{k}$ includes both edges of $\mathcal{M}^{\prime}$ and $\mathcal{M}$ that do not have any endpoint removed).
+Since we remove a total of $\alpha n$ vertices from $G$ to obtain $G_{k}$, each matching in $\mathcal{M}^{\prime}$ and $\mathcal{M}$ loses at most $\alpha n$ edges.
+Therefore, $G_{k}$ is an $\mathsf{ORS}_{n^{\prime\prime}}(c^{\prime\prime}n^{\prime\prime},t^{\prime\prime})$ graph with parameters
+$n^{\prime\prime}=(1-\alpha)n,\qquad c^{\prime\prime}\geq\frac{cn-\alpha n}{n^{\prime\prime}}=\frac{(c-\alpha)n}{(1-\alpha)n}\geq c-\alpha\geq(1-\delta c^{2}/10)c,\qquad t^{\prime\prime}=t.$ (15)
+Additionally, by the construction of $G^{\prime}_{k}$, all vertices $v$ in $G_{k}$ satisfy
+$\deg_{\mathcal{M}^{\prime}}(v)\geq 2\delta ct\geq\delta c^{\prime\prime}t^{\prime\prime}$ (Since $t^{\prime\prime}=t$ and $c^{\prime\prime}\leq\frac{cn}{n^{\prime\prime}}=\frac{cn}{(1-\alpha)n}\leq 2c$ where the last inequality follows from $\alpha<1/2$.)
+This now satisfies the requirements of Lemma 17. Applying it on graph $G_{k}$, we obtain an $\mathsf{ORS}_{n^{\prime}}(c^{\prime}n^{\prime},t^{\prime})$ graph where the number of vertices satisfies
+$n^{\prime}$ $<(1-\delta c^{\prime\prime 2}/2)n^{\prime\prime},$ (16)
+the number of matchings satisfies
+$t^{\prime}\geq\delta c^{\prime\prime 2}t^{\prime\prime}/(8\cdot 2^{4n^{\prime\prime}/c^{\prime\prime}t^{\prime\prime}})=\delta c\cdot t/2^{O(n/ct)}=t/2^{O(n/ct)},$
+and finally since Lemma 17 does not change the size of matchings, we get
+$c^{\prime}n^{\prime}$ $=c^{\prime\prime}n^{\prime\prime}$
+$\geq\frac{c^{\prime\prime}n^{\prime}}{(1-\delta c^{\prime\prime 2}/2)}$ (By 16.)
+$\geq(1+\delta c^{\prime\prime 2}/2)c^{\prime\prime}n^{\prime}$
+$\geq\Big{(}1+\delta\Big{(}(1-\delta c^{2}/10)c\Big{)}^{2}/2\Big{)}\Big{(}(1-\delta c^{2}/10)c\Big{)}n^{\prime},$ (Since the RHS is minimized when $c^{\prime\prime}$ is minimized, thus we can replace $c^{\prime\prime}$ with its LB from 15.)
+$\geq\Big{(}1+0.4\delta c^{2}\Big{)}\Big{(}(1-\delta c^{2}/10)c\Big{)}n^{\prime}$ (Since $\delta\big{(}(1-\delta c^{2}/10)c\big{)}^{2}/2\geq\delta(0.9c)^{2}/2>0.4\delta c^{2}$)
+$>(1+0.07c^{2})cn^{\prime}.$
+Dividing both sides of the inequality by $n^{\prime}$ implies $c^{\prime}\geq(1+0.07c^{2})c$. Taking $\kappa=0.07c^{2}\geq c^{3}/160$ implies the lemma. ∎
+We are now ready to prove Theorem 2.
+Proof of Theorem 2.
+We prove that for some $d=\operatorname{poly}(1/c)$, there does not exist any $\mathsf{ORS}_{n}(cn,t)$ graph $G$ with $t\geq n/\log_{b}^{(d)}n$ where $b=2^{\beta/c}$ for some large enough constant $\beta\geq 1$.
+Note that the base of the iterated log is different from the statement of the theorem, but since $\log^{(x)}_{b}z=\log^{(\Theta(x\log^{*}b))}z$, this implies the theorem as well by letting $\ell=\Theta(d\cdot\log^{*}b)=\operatorname{poly}(1/c)$.
+Suppose for the sake of contradiction that that there exists an $\mathsf{ORS}_{n}(cn,t)$ graph $G$ with $t\geq n/\log_{b}^{(d)}n$.
+Let $\kappa$ be as in Lemma 19.
+We iteratively apply Lemma 19 for $k=\log_{2}(1/c)/\kappa=\operatorname{poly}(c)$ steps to obtain a sequence of graphs $G_{0},G_{1},G_{2},G_{3},\ldots,G_{k}$ where $G_{0}=G$ is the original graph, $G_{1}$ is obtained by applying Lemma 19 on $G_{0}$, $G_{2}$ is obtained by applying Lemma 19 on $G_{1}$, and so on so forth.
+Let us now analyze the ORS properties of graph $G_{k}$, starting with the parameter $c_{k}$.
+Since every application of Lemma 19 multiplies the parameter $c_{i}$ by a factor of at least $(1+\kappa)$, we have
+$c_{k}$ $\geq(1+\kappa)^{k}c$
+$=(1+\kappa)^{\log_{2}(1/c)/k}c$
+$\geq 2^{\log_{2}(1/c)}c$ (Since $(1+\kappa)^{1/\kappa}\geq 2$ for all $0<\kappa\leq 1$)
+$\geq 1.$
+This implies that in graph $G_{k}$, there must be $t_{k}$ edge-disjoint matchings of size $c_{k}n_{k}\geq n_{k}$, but each matching in a graph on $n_{k}$ vertices can have size at most $n_{k}/2$.
+In other words, we must have $t_{k}=0$.
+We will obtain a contradiction by proving that $t_{k}\geq 1$.
+We prove by induction that for every $i\in[d]$,
+$t_{i}\geq\frac{n}{\log^{(d-i)}_{b}n}.$
+For the base case $i=0$ this holds as assumed at the beginning of the proof. By choosing appropriately large $\beta$, we know by Lemma 19 that,
+$t_{i}$ $\geq t_{i-1}/2^{0.5\beta(n_{i-1}/c_{i-1}t_{i-1})}$
+$\geq t_{i-1}/2^{0.5\beta(n/ct_{i-1})}$ (Since $n_{i-1}\leq n$ and $c_{i-1}\geq c$ as we iteratively apply Lemma 19.)
+$=t_{i-1}/b^{0.5(n/t_{i-1})}$ (Since we defined $b=2^{\beta/c}$.)
+$\geq\frac{n}{\log^{(d-i+1)}_{b}n}\cdot b^{-0.5\left(n/\frac{n}{\log^{(d-i+1)}_{b}n}\right)}$ (By the induction hypothesis $t_{i-1}\geq\frac{n}{\log^{(d-i+1)}_{b}n}$.)
+$=\frac{n}{\log^{(d-i+1)}_{b}n}\cdot b^{-0.5\log^{(d-i+1)}_{b}n}$
+$=\frac{n}{\log^{(d-i+1)}_{b}n\cdot\sqrt{\log^{(d-i)}_{b}n}}$
+$\geq\frac{n}{\log^{(d-i)}_{b}n},$
+concluding the proof.
+Now let $d>k=\operatorname{poly}(1/c)$. From the above, we get that $t_{k}\geq n/(\log_{b}^{(d-k)})\geq 1$, which as discussed is a contradiction. ∎
+6.2 When Matchings are (Very) Large
+In this section, we prove that the number of matching of an ORS graph with $r>n/4$ is bounded by a constant.
+More precisely we prove Theorem 4.
+The same upper bound was already known for RS graphs [29], and we show that nearly the same proof carries over to ORS graphs as well.
+Theorem 4.
+For any $c>1/4+\varepsilon$, $\mathsf{ORS}_{n}(cn,t)\leq 1/\varepsilon+1$.
+Proof.
+Suppose $G$ is an $\mathsf{ORS}(r,t)$ graph on $n$ vertices and let $M_{1},\ldots,M_{t}$ be the corresponding ordered list of matchings as in Definition 2.
+We show that for any $i\not=j$ in $[t]$, we have $|V(M_{i})\cap V(M_{j})|\leq r$.
+The rest of the proof follows exactly as in the upper bound of [29, Section 2] for RS, which we omit here.
+Let us assume w.l.o.g.
+that $i<j$ and suppose that $|V(M_{i})\cap V(M_{j})|\geq r+1$.
+By the pigeonhole principle, this means that there must be an edge in $M_{i}$ whose both endpoints are matched in $M_{j}$, contradicting the fact from Definition 2 that $M_{j}$ is an induced matching in $M_{1}\cup\ldots\cup M_{j}$.
+Hence, $|V(M_{i})\cap V(M_{j})|\leq r$ and the proof is complete.
+∎
 Acknowledgements
-The first author thanks Sepehr Assadi, Sanjeev Khanna, Huan Li, Ray Li, Mohammad Roghani, and Aviad Rubinstein for helpful discussions about RS graphs and their applications in dynamic graphs over the years.
+We thank Sepehr Assadi, Sanjeev Khanna, Huan Li, Ray Li, Mohammad Roghani, and Aviad Rubinstein for helpful discussions about RS graphs and their applications in dynamic graphs over the years.
+Additionally, we thank Sepehr Assadi, Jiale Chen, and anonymous FOCS’24 reviewers for helpful suggestions on improving the exposition of the paper.

### v4

@@
 Let us define $Y_{M}:=V\setminus\cup_{v_{i}\in S}N_{i}$.
-From Observation 18, we get that all edges of $M$ must have both endpoints still in $Y_{M}$.
+From Observation 14, we get that all edges of $M$ must have both endpoints still in $Y_{M}$.
 Suppose for now that $M\in\mathcal{M}^{\prime}$ is chosen uniformly at random.
@@
 $\leq n-\sum_{i=1}^{k}\frac{4\delta c}{t}\cdot|N_{i}|^{2}.$ (Since $|N_{i}|\leq t/2$ as $\mathcal{M}$ is the union of $t/2$ matchings.)
-$\leq n-\frac{4\delta c}{t}\cdot(tcn/2-n^{2}/x)$ (By 12.)
+$\leq n-\frac{4\delta c}{t}\cdot(tcn/2-n^{2}/x)$ (By 5.)
 $=n-2\delta c^{2}n+\frac{4\delta cn^{2}}{tx}$
@@
 Instead of the expected value of $Y_{M}$, we need some (weak) concentration.
-Take $\varepsilon=\delta c^{2}/2$ and note from 13 that $(1+\varepsilon)\operatorname{\textbf{E}}[|Y_{M}|]<(1+\varepsilon)(1-\delta c^{2})n<(1-\delta c^{2}/2)n$.
+Take $\varepsilon=\delta c^{2}/2$ and note from 6 that $(1+\varepsilon)\operatorname{\textbf{E}}[|Y_{M}|]<(1+\varepsilon)(1-\delta c^{2})n<(1-\delta c^{2}/2)n$.
 We have
@@
 $\geq\varepsilon/2=\delta c^{2}/4.$ (7)
-We call a matching $M\in\mathcal{M}^{\prime}$ a good matching if $|Y_{M}|<(1-\delta c^{2}/2)n$.
-From 14 we get that at least $(\delta c^{2}/4)|\mathcal{M}^{\prime}|=\delta c^{2}t/8$ matchings are good.
+We call a matching $M\in\mathcal{M}^{\prime}$ a good matching if $|Y_{M}|<(1-\delta c^{2}/2)n$. From 7 we get that at least $(\delta c^{2}/4)|\mathcal{M}^{\prime}|=\delta c^{2}t/8$ matchings are good.
 Step 3: Grouping good matchings.
@@
 Note also that $Y=Y_{M^{\prime}}$ for any other matching $M^{\prime}\in C$ as well since they all match the same subset of pivots.
-Moreover, we have $|Y|<(1-\delta c^{2}/2)n$ since $M$ is good, and all matchings in $C$ only match vertices in $Y$ as discussed earlier (as a consequence of Observation 18).
+Moreover, we have $|Y|<(1-\delta c^{2}/2)n$ since $M$ is good, and all matchings in $C$ only match vertices in $Y$ as discussed earlier (as a consequence of Observation 14).
 This implies that the graph $H$ on vertex set $Y$ and edge-set $C$ (i.e., including all edges of all matchings in $C$) is an $\mathsf{ORS}_{n^{\prime}}(cn,t^{\prime})$ graph for $n^{\prime}<(1-\delta c^{2}/2)n$ and $t^{\prime}\geq\delta c^{2}t/(8\cdot 2^{x})$ as desired.
@@
 $c^{\prime}n^{\prime}$ $=c^{\prime\prime}n^{\prime\prime}$
-$\geq\frac{c^{\prime\prime}n^{\prime}}{(1-\delta c^{\prime\prime 2}/2)}$ (By 16.)
+$\geq\frac{c^{\prime\prime}n^{\prime}}{(1-\delta c^{\prime\prime 2}/2)}$ (By 9.)
 $\geq(1+\delta c^{\prime\prime 2}/2)c^{\prime\prime}n^{\prime}$
-$\geq\Big{(}1+\delta\Big{(}(1-\delta c^{2}/10)c\Big{)}^{2}/2\Big{)}\Big{(}(1-\delta c^{2}/10)c\Big{)}n^{\prime},$ (Since the RHS is minimized when $c^{\prime\prime}$ is minimized, thus we can replace $c^{\prime\prime}$ with its LB from 15.)
-$\geq\Big{(}1+0.4\delta c^{2}\Big{)}\Big{(}(1-\delta c^{2}/10)c\Big{)}n^{\prime}$ (Since $\delta\big{(}(1-\delta c^{2}/10)c\big{)}^{2}/2\geq\delta(0.9c)^{2}/2>0.4\delta c^{2}$)
-$>(1+0.07c^{2})cn^{\prime}.$
-Dividing both sides of the inequality by $n^{\prime}$ implies $c^{\prime}\geq(1+0.07c^{2})c$. Taking $\kappa=0.07c^{2}\geq c^{3}/160$ implies the lemma. ∎
-We are now ready to prove Theorem 2.
-Proof of Theorem 2.
-We prove that for some $d=\operatorname{poly}(1/c)$, there does not exist any $\mathsf{ORS}_{n}(cn,t)$ graph $G$ with $t\geq n/\log_{b}^{(d)}n$ where $b=2^{\beta/c}$ for some large enough constant $\beta\geq 1$.
-Note that the base of the iterated log is different from the statement of the theorem, but since $\log^{(x)}_{b}z=\log^{(\Theta(x\log^{*}b))}z$, this implies the theorem as well by letting $\ell=\Theta(d\cdot\log^{*}b)=\operatorname{poly}(1/c)$.
-Suppose for the sake of contradiction that that there exists an $\mathsf{ORS}_{n}(cn,t)$ graph $G$ with $t\geq n/\log_{b}^{(d)}n$.
-Let $\kappa$ be as in Lemma 19.
-We iteratively apply Lemma 19 for $k=\log_{2}(1/c)/\kappa=\operatorname{poly}(c)$ steps to obtain a sequence of graphs $G_{0},G_{1},G_{2},G_{3},\ldots,G_{k}$ where $G_{0}=G$ is the original graph, $G_{1}$ is obtained by applying Lemma 19 on $G_{0}$, $G_{2}$ is obtained by applying Lemma 19 on $G_{1}$, and so on so forth.
-Let us now analyze the ORS properties of graph $G_{k}$, starting with the parameter $c_{k}$.
-Since every application of Lemma 19 multiplies the parameter $c_{i}$ by a factor of at least $(1+\kappa)$, we have
-$c_{k}$ $\geq(1+\kappa)^{k}c$
-$=(1+\kappa)^{\log_{2}(1/c)/k}c$
-$\geq 2^{\log_{2}(1/c)}c$ (Since $(1+\kappa)^{1/\kappa}\geq 2$ for all $0<\kappa\leq 1$)
-$\geq 1.$
-This implies that in graph $G_{k}$, there must be $t_{k}$ edge-disjoint matchings of size $c_{k}n_{k}\geq n_{k}$, but each matching in a graph on $n_{k}$ vertices can have size at most $n_{k}/2$.
-In other words, we must have $t_{k}=0$.
-We will obtain a contradiction by proving that $t_{k}\geq 1$.
-We prove by induction that for every $i\in[d]$,
-$t_{i}\geq\frac{n}{\log^{(d-i)}_{b}n}.$
-For the base case $i=0$ this holds as assumed at the beginning of the proof. By choosing appropriately large $\beta$, we know by Lemma 19 that,
-$t_{i}$ $\geq t_{i-1}/2^{0.5\beta(n_{i-1}/c_{i-1}t_{i-1})}$
-$\geq t_{i-1}/2^{0.5\beta(n/ct_{i-1})}$ (Since $n_{i-1}\leq n$ and $c_{i-1}\geq c$ as we iteratively apply Lemma 19.)
-$=t_{i-1}/b^{0.5(n/t_{i-1})}$ (Since we defined $b=2^{\beta/c}$.)
-$\geq\frac{n}{\log^{(d-i+1)}_{b}n}\cdot b^{-0.5\left(n/\frac{n}{\log^{(d-i+1)}_{b}n}\right)}$ (By the induction hypothesis $t_{i-1}\geq\frac{n}{\log^{(d-i+1)}_{b}n}$.)
-$=\frac{n}{\log^{(d-i+1)}_{b}n}\cdot b^{-0.5\log^{(d-i+1)}_{b}n}$
-$=\frac{n}{\log^{(d-i+1)}_{b}n\cdot\sqrt{\log^{(d-i)}_{b}n}}$
-$\geq\frac{n}{\log^{(d-i)}_{b}n},$
-concluding the proof.
-Now let $d>k=\operatorname{poly}(1/c)$. From the above, we get that $t_{k}\geq n/(\log_{b}^{(d-k)})\geq 1$, which as discussed is a contradiction. ∎
-5.2 When Matchings are (Very) Large
-The goal of this section is to prove that the number of matching of an $\mathsf{ORS}$ graph with $r>n/4$ is bounded by a constant value.
-More precisely we prove Theorem 4.
-An equivalent of this theorem was already known for RS graphs (see [29]), and we show that almost the same proof carries over to ORS graphs as well.
-Theorem 3.
-For any $c>1/4+\varepsilon$, $\mathsf{ORS}_{n}(cn,t)\leq 1/\varepsilon+1$.
-First, we argue that the number of common vertices of any two matching in ORS is at most $r$, using this we can prove that a constant value bounds the number of matching since the number of subsets of length $2r$ with at most $r$ common vertices between any two sets.
-Lemma 16.
-Suppose $G$ is an $\mathsf{ORS}(r,t)$ graph on $n$ vertices and let $M_{1},\ldots,M_{t}$ be the corresponding ordered list of matchings as in Definition 2.
-Then for any $i\not=j$ in $[t]$, we have $|V(M_{i})\cap V(M_{j})|\leq r$.
-Proof.
-Let us assume w.l.o.g.
-that $i<j$ and suppose that $|V(M_{i})\cap V(M_{j})|\geq r+1$.
-By the pigeonhole principle, this means that there must be an edge in $M_{i}$ whose both endpoints are matched in $M_{j}$, contradicting the fact from Definition 2 that $M_{j}$ is an induced matching in $M_{1}\cup\ldots\cup M_{j}$.
-∎
-Now, we use a similar approach as [29] to prove that $t<1/\varepsilon$. This theorem only uses the fact that the number of common vertices of two matchings is at most $r$.
-Proof of Theorem 4.
-Using Lemma 16 we have that the distance between sets of vertices of any two matchings is at most $r$, this implies a constant upper bound for the size of $\mathsf{ORS}$ using bounds from coding theory.
-To improve the upper bound we set $v_{i}\in\{0,1\}^{n}$, a vector that represents vertices in $V(M_{i})$.
-Let us use $a_{i}$ to denote the number of vectors $v_{j}$ with 00 in their $i$-th coordinate and let $b_{i}$ be the number of vectors with $1$ in their $i$-th coordinate.
-If we set $v_{0}$ a vector of all zeros.
-Note that we have $a_{i}+b_{i}=t+1$.
-Now, by double counting we have:
-$2r{t+1\choose 2}\leq\sum_{0\leq i<j\leq t}dist(v_{i},v_{j})=\sum_{i=1}^{n}a_{i}b_{i}.$ (10)
-By $dist(v_{i},v_{j})$ we mean the Hamming distance which satisfies the following inequality:
-$dist(v_{i},v_{j})=|V_{i}|+|V_{j}|-2|V_{i}\cap V_{j}|\geq 2r.$
-Note that in 10, $a_{i}b_{i}$ maximizes when we have $a_{i}=b_{i}=(t+1)/2$ for odd $t$ and $a_{i}=(t+2)/2,b_{i}=t/2$ for even $t$. This concludes that $t<1/\varepsilon+1$
-6 Bounding Density of (Ordered) Ruzsa-Szemerédi Graphs
-In this section we prove upper bounds on the density of ORS graphs.
-Our main result of this section is a proof of Theorem 2 which we present in Section 6.1.
-We then show that a much stronger upper bound can be proved if matchings cover more than half of vertices in Section 6.2.
-Before presenting our proof, let us first briefly discuss how the triangle removal lemma is useful for upper bounding density of RS graphs and why it does not seem to help for upper bounding ORS.
-The triangle removal lemma states that so long as the number of triangles in a graph are $o(n^{3})$, then we can make the graph triangle free by removing $o(n^{2})$ of its edges.
-Suppose we have a bipartite RS graph with induced matchings $M_{1},\ldots,M_{k}$.
-Add $k$ vertices $v_{1},\ldots,v_{k}$ to this graph and connect each $v_{i}$ to all the vertices matched by $M_{i}$.
-Now, crucially, because each $M_{i}$ is an induced matching, each $v_{i}$ is part of exactly $|M_{i}|$ triangles, one for each edge of $M_{i}$.
-As such, the total number of triangles can be upper bounded by $O(kn)=O(n^{2})$ which is well within the regime that one can apply triangle removal lemma.
-However, because the matchings in ORS are not induced matchings of the whole graph, the number of triangles cannot simply be upper bounded by $O(n^{2})$ after adding the auxiliary vertices.
-In fact, our upper bound completely deviates from this approach and does not rely on the triangle removal lemma.
-6.1 Linear Matchings
-The following lemma, which is one of our main tools in proving Theorem 2, shows that so long as vertex degrees are larger than a threshold, we can reduce the number of vertices rather significantly without hurting the size of a relatively large number of induced matchings.
-Our proof of Lemma 17 builds on the techniques developed for upper bounding RS graphs with matchings of size very close to $n/4$ in [29] that we extend to ORS graphs with balanced degrees.
-Lemma 17.
-Let $G$ be an $\mathsf{ORS}_{n}(r,t)$ graph with $r=cn$ with even $t$.
-Let $M_{1},\ldots,M_{t}$ be the $t$ matchings in $G$ as defined in Definition 2.
-Let us partition these matchings into two subsets
-=$\mathcal{M}=\{M_{1},\ldots,M_{t/2}\}\qquad\text{and}\qquad\mathcal{M}^{\prime}=\{M_{t/2+1},\ldots,M_{t}\}.$
-Suppose that for every vertex $v$, it holds that $\deg_{\mathcal{M}^{\prime}}(v)\geq\delta ct$ for some $\delta>0$.
-Then there exists an $\mathsf{ORS}_{n^{\prime}}(cn,t^{\prime})$ graph $H$ on $n^{\prime}<(1-\delta c^{2}/2)n$ vertices where $t^{\prime}\geq\delta c^{2}t/(8\cdot 2^{x})$ and $x=4n/ct$.
-Let us start with an observation that is extremely helpful in proving Lemma 17.
-Observation 18.
-Take a matching $M^{\prime}\in\mathcal{M^{\prime}}$ and take any vertex $v$ matched by $M^{\prime}$.
-Let $N_{\mathcal{M}}(v)$ be the set of neighbors of $v$ in $\mathcal{M}$.
-That is, $u\in N_{\mathcal{M}}(v)$ iff there is $M\in\mathcal{M}$ such that $(v,u)\in M$.
-Then no vertex in $N_{\mathcal{M}}(v)$ can be matched in $M^{\prime}$.
-Proof.
-Suppose there is $u\in N_{\mathcal{M}}(v)$ that is matched by $M^{\prime}$.
-Let $(u,w)\in M^{\prime}$ be the matching edge involving $u$.
-Also take the edge $(v,y)\in M^{\prime}$ that involves $v$ (which exists by definition of $v$).
-Note that since the matchings $M_{1},\ldots,M_{t}$ are edge-disjoint by Definition 2, $v$ and $u$ cannot be matched together in $M^{\prime}$ (as $v$ and $u$ must be matched in some other matching in $\mathcal{M}$ by definition of $N_{\mathcal{M}}(v)$).
-Now since $(u,w),(v,y)\in M^{\prime}$ but $(v,u)$ belongs to some matching in $\mathcal{M}$ that comes before $M^{\prime}$ in the ordering, matching $M^{\prime}$ cannot be an induced matching among its previous matchings, contradicting Definition 2 for ORS graphs.
-∎
-We are now ready to prove Lemma 17.
-Proof of Lemma 17.
-We explain the proof in a few steps.
-Step 1: The pivots.
-We iteratively take a vertex of highest remaining degree in $\mathcal{M}$, and remove its neighbors in $\mathcal{M}$ from the graph.
-More formally, take the graph $G_{0}=(V_{0},E_{0})$ where $V_{0}=V(G)$ and $E_{0}$ is the union of the matchings in $\mathcal{M}$.
-At step $i$, we choose $v_{i}$ in $V_{i-1}$ with the maximum degree in $G_{i-1}$.
-Let us define $N_{i}$ as the neighbors of $v_{i}$ in $G_{i-1}$.
-We then remove $v_{i}$ and its neighbors in $G_{i-1}$ to define $G_{i}=(V_{i},E_{i})$.
-Namely, $V_{i}=V_{i-1}\setminus(\{v_{i}\}\cup N_{i})$ and $E_{i}$ includes all edges in $E_{i-1}$ except those that have at least one endpoint that does not belong to $V_{i}$.
-Let $k$ be the maximum integer such that $|N_{k}|>n/x$.
-We define $P=\{v_{1},\ldots,v_{k}\}$ to be the set of pivots.
-Note that since $N_{1},\dots,N_{k}$ are disjoint sets, we get that $n\geq\sum_{i=1}^{k}|N_{i}|\geq kn/x$ which implies that
-$|P|\leq x.$ (11)
-Moreover, note that when removing $v_{i}$, we remove $|N_{i}|$ vertices from the graph and each of those vertices has remaining degree at most $|N_{i}|$ (as we choose $v_{i}$ to be the vertex of highest remaining degree).
-In other words, we remove at most $|N_{i}|^{2}$ from the graph after removing $v_{i}$ and its remaining neighbors.
-On the other hand, after removing $v_{1},\ldots,v_{k}$ and their neighbors, the maximum degree in the graph is at most $n/x$ (by definition of $k$), and so there are at most $n^{2}/x$ edges in the graph.
-This implies that:
-$|E_{0}|-\sum_{i=1}^{k}|N_{i}|^{2}\leq n^{2}/x.$
-Given that $E_{0}$ includes $t/2$ edge disjoint matchings of size $cn$, we get that $|E_{0}|\geq tcn/2$. Plugged into the inequality above, this implies that
-$\sum_{i=1}^{k}|N_{i}|^{2}\geq|E_{0}|-n^{2}/x\geq tcn/2-n^{2}/x.$ (12)
-Step 2: Vertex reduction.
-Take a matching $M\in\mathcal{M}^{\prime}$.
-Let $S_{M}$ be the set of pivots matched by $M$, i.e., $S_{M}=P\cap V(M)$.
-Let us define $Y_{M}:=V\setminus\cup_{v_{i}\in S}N_{i}$.
-From Observation 18, we get that all edges of $M$ must have both endpoints still in $Y_{M}$.
-Suppose for now that $M\in\mathcal{M}^{\prime}$ is chosen uniformly at random.
-We first argue that the expected size of $|Y_{M}|$ is relatively small (despite it preserving all edges of $M$ completely).
-We have
-$\operatorname{\textbf{E}}_{M\sim\mathcal{M}^{\prime}}[|Y_{M}|]$ $=n-\sum_{i=1}^{k}\Pr[v_{i}\in V(M)]\cdot|N_{i}|$
-$=n-\sum_{i=1}^{k}\frac{\deg_{\mathcal{M}^{\prime}}(v_{i})}{|\mathcal{M}^{\prime}|}\cdot|N_{i}|$ (As $M$ is chosen uniformly from $\mathcal{M}^{\prime}$ and $v_{i}$ is matched in $\deg_{\mathcal{M}^{\prime}}(v_{i})$ of matchings in $\mathcal{M}^{\prime}$.)
-$=n-\sum_{i=1}^{k}\frac{2\deg_{\mathcal{M}^{\prime}}(v_{i})}{t}\cdot|N_{i}|$ (Since $|\mathcal{M}^{\prime}|=t/2$.)
-$\leq n-\sum_{i=1}^{k}\frac{2\delta ct}{t}\cdot|N_{i}|$ (Since $\deg_{\mathcal{M}^{\prime}}(v_{i})\geq\delta ct$ as assumed in Lemma 17.)
-$\leq n-\sum_{i=1}^{k}\frac{4\delta c}{t}\cdot|N_{i}|^{2}.$ (Since $|N_{i}|\leq t/2$ as $\mathcal{M}$ is the union of $t/2$ matchings.)
-$\leq n-\frac{4\delta c}{t}\cdot(tcn/2-n^{2}/x)$ (By 12.)
-$=n-2\delta c^{2}n+\frac{4\delta cn^{2}}{tx}$
-$\leq n-2\delta c^{2}n+\frac{4\delta cn^{2}}{(4n/cx)x}$ (Since $t=4n/cx$ by definition of $x$ in Lemma 17.)
-$=n-2\delta c^{2}n+\delta c^{2}n$
-$=(1-\delta c^{2})n.$ (13)
-Instead of the expected value of $Y_{M}$, we need some (weak) concentration.
-Take $\varepsilon=\delta c^{2}/2$ and note from 13 that $(1+\varepsilon)\operatorname{\textbf{E}}[|Y_{M}|]<(1+\varepsilon)(1-\delta c^{2})n<(1-\delta c^{2}/2)n$.
-We have
-$\Pr_{M\sim\mathcal{M}^{\prime}}\Big{[}|Y_{M}|<(1+\varepsilon)\operatorname{\textbf{E}}[|Y_{M}|]<(1-\delta c^{2}/2)n\Big{]}$ $=1-\Pr_{M\sim\mathcal{M}^{\prime}}\Big{[}|Y_{M}|>(1+\varepsilon)\operatorname{\textbf{E}}[|Y_{M}|]\Big{]}$
-$\geq 1-\frac{1}{1+\varepsilon}$ (By Markov’s inequality.)
-$\geq\varepsilon/2=\delta c^{2}/4.$ (14)
-We call a matching $M\in\mathcal{M}^{\prime}$ a good matching if $|Y_{M}|<(1-\delta c^{2}/2)n$.
-From 14 we get that at least $(\delta c^{2}/4)|\mathcal{M}^{\prime}|=\delta c^{2}t/8$ matchings are good.
-Step 3: Grouping good matchings.
-Now, we partition the good matchings in $\mathcal{M}^{\prime}$ into $2^{|P|}\leq 2^{x}$ classes depending on which subset of the pivots they match.
-Since there are at least $\delta c^{2}t/8$ good matchings, at least $\frac{\delta c^{2}t}{8\cdot 2^{x}}$ of them must belong to the same class $C$.
-Let $Y=Y_{M}$ for some $M\in C$.
-Note also that $Y=Y_{M^{\prime}}$ for any other matching $M^{\prime}\in C$ as well since they all match the same subset of pivots.
-Moreover, we have $|Y|<(1-\delta c^{2}/2)n$ since $M$ is good, and all matchings in $C$ only match vertices in $Y$ as discussed earlier (as a consequence of Observation 18).
-This implies that the graph $H$ on vertex set $Y$ and edge-set $C$ (i.e., including all edges of all matchings in $C$) is an $\mathsf{ORS}_{n^{\prime}}(cn,t^{\prime})$ graph for $n^{\prime}<(1-\delta c^{2}/2)n$ and $t^{\prime}\geq\delta c^{2}t/(8\cdot 2^{x})$ as desired.
-∎
-The next lemma follows by applying Lemma 17 after carefully pruning vertices of low degree in $\mathcal{M}^{\prime}$.
-Intuitively, Lemma 19 significantly reduces the number of vertices and shows that a relatively large number of matchings will have a lot of edges in the remaining graph.
-Lemma 19.
-Let $G$ be an $\mathsf{ORS}_{n}(cn,t)$ graph for even $t$.
-Then for some $\kappa\geq c^{3}/160$, there exist $\mathsf{ORS}_{n^{\prime}}(c^{\prime}n^{\prime},t^{\prime})$ graphs such that $n^{\prime}\leq n$, $c^{\prime}\geq(1+\kappa)c$ and $t^{\prime}\geq t/2^{O(n/ct)}$.
-Proof.
-Let $\mathcal{M}$ and $\mathcal{M}^{\prime}$ be defined for $G$ as in Lemma 17.
-We would like to apply Lemma 17 iteratively to prove Lemma 19.
-However, Lemma 17 requires a lower bound on the degrees in $\mathcal{M}^{\prime}$ which might not necessarily hold for our graph $G$.
-Therefore to be able to use it, we first have to prune vertices of small degrees.
-Let $G^{\prime}_{0}$ be the subgraph of $G$ only including the edges of the matchings in $\mathcal{M}^{\prime}$.
-Let $\delta=1/4$.
-We iteratively remove vertices of low degree.
-Formally, in step $i$, if there is a vertex $v_{i}$ with degree less than $2\delta ct$ in $G^{\prime}_{i-1}$, we define the graph $G^{\prime}_{i}$ of the next step to be graph obtained after removing $v_{i}$ and its edges from $G^{\prime}_{i-1}$.
-Let $G^{\prime}_{k}$ be the final graph that does not have any vertex of degree smaller than $2\delta ct$.
-Suppose that $k=\alpha n$; that is, we remove $\alpha n$ vertices in the process above. We consider two cases depending on the value of $\alpha$.
-Case 1: $\alpha\geq\delta c^{3}/10$.
-Since we remove at most $\delta ct$ edges in every step, the total number of removed edges is at most $\delta\alpha ctn$ over the $k=\alpha n$ steps of constructing $G^{\prime}_{k}$.
-Say a matching $M\in\mathcal{M}^{\prime}$ is damaged if a total of at least $3\delta\alpha cn$ of its edges have been removed.
-In other words, $M$ is damaged if less than $|M|-3\delta\alpha cn=(1-3\delta\alpha)cn$ of its edges belong to $G^{\prime}_{k}$.
-Since the total number of edges removed is at most $\delta\alpha ctn$ and each damaged matching has at least $3\delta\alpha cn$ of its edges removed, the total number of damaged matchings can be upper bounded by $\delta\alpha ctn/3\delta\alpha cn=t/3$.
-Since $|\mathcal{M}^{\prime}|=t/2$, at least $t/6$ matchings in $\mathcal{M}^{\prime}$ are not damaged, and have size at least $(1-3\delta\alpha)cn$.
-These matchings themselves form an $\mathsf{ORS}_{n^{\prime}}(c^{\prime}n^{\prime},t/6)$ graph with parameters
-$n^{\prime}=(1-\alpha)n\leq(1-\delta c^{3}/10)n\stackrel{{\scriptstyle(\delta=1/4)}}{{=}}(1-c^{3}/40)n,$
-and
-$c^{\prime}n^{\prime}\geq(1-3\delta\alpha)cn=(1-3\delta\alpha)c\frac{n^{\prime}}{1-\alpha}\stackrel{{\scriptstyle(\delta=1/4)}}{{=}}\left(1-\frac{3}{4}\alpha\right)c\frac{n^{\prime}}{1-\alpha}\geq\left(1+\frac{1}{4}\alpha\right)cn^{\prime}\geq(1+c^{3}/160)cn^{\prime},$
-which implies that $c^{\prime}\geq(1+c^{3}/160)c$. Taking $\kappa=c^{3}/160$ proves the lemma in this case.
-Case 2 – $\alpha<\delta c^{3}/10$:
-Let $G_{k}$ be the graph $G$ after removing vertices $v_{1},\ldots,v_{k}$ (the difference between $G_{k}$ and $G^{\prime}_{k}$ is that $G^{\prime}_{k}$ only includes the edges of $\mathcal{M}^{\prime}$ but $G_{k}$ includes both edges of $\mathcal{M}^{\prime}$ and $\mathcal{M}$ that do not have any endpoint removed).
-Since we remove a total of $\alpha n$ vertices from $G$ to obtain $G_{k}$, each matching in $\mathcal{M}^{\prime}$ and $\mathcal{M}$ loses at most $\alpha n$ edges.
-Therefore, $G_{k}$ is an $\mathsf{ORS}_{n^{\prime\prime}}(c^{\prime\prime}n^{\prime\prime},t^{\prime\prime})$ graph with parameters
-$n^{\prime\prime}=(1-\alpha)n,\qquad c^{\prime\prime}\geq\frac{cn-\alpha n}{n^{\prime\prime}}=\frac{(c-\alpha)n}{(1-\alpha)n}\geq c-\alpha\geq(1-\delta c^{2}/10)c,\qquad t^{\prime\prime}=t.$ (15)
-Additionally, by the construction of $G^{\prime}_{k}$, all vertices $v$ in $G_{k}$ satisfy
-$\deg_{\mathcal{M}^{\prime}}(v)\geq 2\delta ct\geq\delta c^{\prime\prime}t^{\prime\prime}$ (Since $t^{\prime\prime}=t$ and $c^{\prime\prime}\leq\frac{cn}{n^{\prime\prime}}=\frac{cn}{(1-\alpha)n}\leq 2c$ where the last inequality follows from $\alpha<1/2$.)
-This now satisfies the requirements of Lemma 17. Applying it on graph $G_{k}$, we obtain an $\mathsf{ORS}_{n^{\prime}}(c^{\prime}n^{\prime},t^{\prime})$ graph where the number of vertices satisfies
-$n^{\prime}$ $<(1-\delta c^{\prime\prime 2}/2)n^{\prime\prime},$ (16)
-the number of matchings satisfies
-$t^{\prime}\geq\delta c^{\prime\prime 2}t^{\prime\prime}/(8\cdot 2^{4n^{\prime\prime}/c^{\prime\prime}t^{\prime\prime}})=\delta c\cdot t/2^{O(n/ct)}=t/2^{O(n/ct)},$
-and finally since Lemma 17 does not change the size of matchings, we get
-$c^{\prime}n^{\prime}$ $=c^{\prime\prime}n^{\prime\prime}$
-$\geq\frac{c^{\prime\prime}n^{\prime}}{(1-\delta c^{\prime\prime 2}/2)}$ (By 16.)
-$\geq(1+\delta c^{\prime\prime 2}/2)c^{\prime\prime}n^{\prime}$
-$\geq\Big{(}1+\delta\Big{(}(1-\delta c^{2}/10)c\Big{)}^{2}/2\Big{)}\Big{(}(1-\delta c^{2}/10)c\Big{)}n^{\prime},$ (Since the RHS is minimized when $c^{\prime\prime}$ is minimized, thus we can replace $c^{\prime\prime}$ with its LB from 15.)
+$\geq\Big{(}1+\delta\Big{(}(1-\delta c^{2}/10)c\Big{)}^{2}/2\Big{)}\Big{(}(1-\delta c^{2}/10)c\Big{)}n^{\prime},$ (Since the RHS is minimized when $c^{\prime\prime}$ is minimized, thus we can replace $c^{\prime\prime}$ with its LB from 8.)
 $\geq\Big{(}1+0.4\delta c^{2}\Big{)}\Big{(}(1-\delta c^{2}/10)c\Big{)}n^{\prime}$ (Since $\delta\big{(}(1-\delta c^{2}/10)c\big{)}^{2}/2\geq\delta(0.9c)^{2}/2>0.4\delta c^{2}$)
@@
 Theorem 3.
-For any $c>1/4+\varepsilon$, $\mathsf{ORS}_{n}(cn,t)\leq 1/\varepsilon+1$.
+For any $c>1/4+\varepsilon$, $\mathsf{ORS}_{n}(cn)\leq 1/\varepsilon+1$.
 Proof.